### v1

# Redefining Machine Unlearning: A Conformal Prediction-Motivated Approach

Abstract
Machine unlearning seeks to systematically remove specified data from a trained model, effectively achieving a state as though the data had never been encountered during training. While metrics such as Unlearning Accuracy (UA) and Membership Inference Attack (MIA) provide a baseline for assessing unlearning performance, they fall short of evaluating the completeness and reliability of forgetting. This is because the ground truth labels remain potential candidates within the scope of uncertainty quantification, leaving gaps in the evaluation of true forgetting. In this paper, we identify critical limitations in existing unlearning metrics and propose enhanced evaluation metrics inspired by conformal prediction. Our metrics can effectively capture the extent to which ground truth labels are excluded from the prediction set. Furthermore, we observe that many existing machine unlearning methods do not achieve satisfactory forgetting performance when evaluated with our new metrics. To address this, we propose an unlearning framework that integrates conformal prediction insights into Carlini & Wagner adversarial attack loss. Extensive experiments on the image classification task demonstrate that our enhanced metrics offer deeper insights into unlearning effectiveness, and that our unlearning framework significantly improves the forgetting quality of unlearning methods.
1 Introduction
Machine unlearning has become crucial in advancing data privacy, especially under legal requirements like the General Data Protection Regulation (GDPR) (Bourtoule et al., 2021). These regulations emphasize the right for individuals to have their data removed or forgotten, creating a demand for machine unlearning methods that can enable machine learning models to behave as if specific forget data were never used in training stage. The existing post hoc machine unlearning methods can be categorized into training-based (Graves et al., 2021; Warnecke et al., 2021; Thudi et al., 2022; Tarun et al., 2023) and training-free (Guo et al., 2019; Sekhari et al., 2021; Nguyen et al., 2020; Golatkar et al., 2020b, 2021; Foster et al., 2024) approaches, depending on whether they require any traditional model training steps during the unlearning process (Foster et al., 2024).
Table 1: Grad-CAM maps of one original model in CIFAR-10 with ResNet18 and two corresponding unlearning models. The Prediction row indicates whether the model correctly predicts the image’s true label, while the In Set row represents whether the true label is included in the prediction set. Although the Finetune unlearning method, can misclassify the forget data, Grad-CAM can still highlight key features of the object under this model since the true label is included in the prediction set. In contrast, our unlearning method removes the true label from the set, with activation regions shifting significantly away from the object’s key features. This confirms that the forgetting quality is better if the true label can be excluded from the prediction set.
Class Name Forget Data Original Model Finetune Method Our Method
Wok [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Swimming Trunks [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Prediction
In Set
To measure the forgetting quality and predictive performance of the model after the unlearning process (i.e., unlearning model), several unlearning metrics have been proposed (Kashef, 2021; Shokri et al., 2017; Chen et al., 2021; Brophy & Lowd, 2021; Cao & Yang, 2015). However, existing unlearning metrics, such as unlearning accuracy (UA) (Brophy & Lowd, 2021; Foster et al., 2024) and membership inference attack (MIA) (Shokri et al., 2017; Chen et al., 2021), fall short in fully evaluating forgetting quality and reliability. These metrics primarily focus on whether models can predict forget data accurately without sufficiently addressing how well the model forgets. In a nutshell, misclassifying forget data does not mean that the model has completely forgotten it to some extent.
To verify this view, we apply conformal prediction (Papadopoulos et al., 2002; Lei & Wasserman, 2014) to the unlearning models. Through experiments, we find that although the model misclassifies some forget data, a significant portion of these misclassified instances still appear in the conformal prediction set. Pictures in Table 1, which visualize the important feature of models’ prediction of models by using Grad-CAM (Selvaraju et al., 2017), further explain this phenomenon. Despite the Finetune method incorrectly predicting the forget data, the Grad-CAM maps still focus heavily on the important features of the object itself.
Based on the above findings, we design two novel metrics that capture the uncertainty and robustness of unlearning performance more effectively inspired by conformal prediction. Additionally, motivated by Carlini & Wagner (C&W) attack (Carlini & Wagner, 2017) and conformal prediction, we propose a general unlearning framework to improve training-based unlearning methods and promote reliable forgetting. Grad-CAM maps of our method in Table 1 reveal that once the true label no longer falls within the conformal prediction set, the activation regions shift significantly. To sum up, the contributions of our work are as follows:
  • We identify pivotal limitations in current unlearning metrics, as they overlook misclassified data where ground truth labels remain potential candidates under uncertainty quantification.
  • We design two novel metrics to address the limitations motivated by conformal prediction.
  • We propose a general unlearning framework for training-based machine unlearning methods motivated by conformal prediction and C&W loss.
  • Extensive experiments demonstrate the effectiveness of novel metrics and our unlearning framework.
2 Enhancing Metrics for Machine Unlearning Based on Conformal Prediction
2.1 Preliminaries and Notations
Machine Unlearning.
Machine unlearning is the targeted removal of certain training data effects from a machine learning model. In our work, two different forgetting scenarios are considered: (i) random data forgetting focuses on randomly forgetting specific instances within the training data, and (ii) class-wise forgetting aims to remove all information associated with an entire class. Let $\mathcal{D}_{train}$ denote the original training data used to obtain an original model ${\bm{\theta}}_{o}$. We split the whole training data $\mathcal{D}_{train}$ into two subsets, forget data $\mathcal{D}_{f}$ and retain data $\mathcal{D}_{r}=\mathcal{D}_{train}\setminus\mathcal{D}_{f}$. In random data forgetting, $\mathcal{D}_{test}$ represents test data. In class-wise forgetting, $\mathcal{D}_{tf}$ corresponds to the test-forget data, exclusively containing the forget class, while $\mathcal{D}_{tr}$ represents the test-retain data within the test data $\mathcal{D}_{test}$. Let $\theta_{u}$ denote the model after the unlearning process, where the influence of forget data $\mathcal{D}_{f}$ has been removed. The purpose of machine unlearning is to enable $\theta_{u}$ to forget $\mathcal{D}_{f}$ (thus no longer recognizing $\mathcal{D}_{tf}$ in the class-wise forgetting scenario), while maintaining accuracy on the retain data $\mathcal{D}_{r}$ and avoiding a significant drop in accuracy on $\mathcal{D}_{test}$ in the random data forgetting scenario or $\mathcal{D}_{tr}$ in the class-wise setting.
Conformal Prediction.
Conformal prediction is proposed to quantify uncertainty in machine learning models, providing prediction sets that contain the true label with a guaranteed probability (Angelopoulos & Bates, 2021). Among the various types of conformal prediction, this work specifically focuses on split conformal prediction (SCP) [^1] Note that while the goal is to remove the influence of the forget data so that it behaves similarly to the calibration data, the exchangeability property may not always hold in machine unlearning settings. Here, we are directly leveraging the concept of conformal prediction to evaluate machine unlearning performance.. To construct the conformal prediction set, it involves four steps:
  1. 1.
    Calibration Data. SCP first chooses unseen data as calibration data. The number of calibration data points should be enough to evaluate the model’s uncertainty.
  2. 2.
    Non-conformity Score. In our work, we follow the conventional choice and set the non-conformity score as
    $S({\bm{x}},y_{i})=1-p_{i}({\bm{x}}),$ (1)
    where $p_{i}({\bm{x}})$ represents the probability output (through a softmax function) for the class $y_{i}$.
  3. 3.
    Quantile Computation. Given a target miscoverage rate $\alpha\in[0,1]$, SCP obtains threshold $\hat{q}$ by taking the $1-\alpha$ quantile of the non-conformity score on the calibration data $\mathcal{D}_{c}$.
  4. 4.
    Prediction Set. For the data point ${\bm{x}}$ that needs to be tested, labels with non-conformity scores lower than the threshold $\hat{q}$ are selected for the final prediction set:
    ${\mathbb{C}}({\bm{x}})=\{y_{i}:S({\bm{x}},y_{i})\leq\hat{q}\},$ (2)
    where $y_{i}$ represents different classes.
2.2 Identifying Limitations in Current Unlearning Metrics
Evaluating the effectiveness of machine unlearning involves a trade-off between how well the model forgets unwanted data (forgetting quality) and how much it retains useful knowledge (predictive performance). The first key question we pose is as follows:
(Q1) Are existing commonly used metrics sufficient to evaluate unlearning performance?
[…]
$10\%$ Forgetting $50\%$ Forgetting
Methods UA $\uparrow$ RA $\uparrow$ TA $\uparrow$ MIA $\downarrow$ UA $\uparrow$ RA $\uparrow$ TA $\uparrow$ MIA $\downarrow$
RT 8.62 99.69 91.83 86.92 10.98 99.80 89.16 82.79
FT 3.84 98.14 91.57 92.00 2.59 99.08 91.77 92.92
RL 7.55 97.41 90.60 74.21 10.48 93.91 85.78 61.15
Table 3: Mis-label (mis-classification) count and in-set ratio of UA and MIA metrics for RT, FT and RL on CIFAT-10 with ResNet-18 under $10\%$ and $50\%$ random data forgetting scenarios. In all settings, over 30% of mis-label data remains within the conformal prediction set in both UA and MIA. More results of other unlearning methods can be found in Appendix B.
$10\%$ Forgetting $50\%$ Forgetting
Methods Mis-label $\uparrow$ In-set $\downarrow$ Ratio $\downarrow$ Mis-label $\uparrow$ In-set $\downarrow$ Ratio $\downarrow$
Mis-label and In-set Ratio of UA
RT 431 132 30.6% 2,745 1,573 57.3%
FT 192 112 58.3% 647 431 66.6%
RL 380 173 45.5% 2,625 1,795 68.4%
Mis-label and In-set Ratio of MIA
RT 654 209 32.0% 4,303 1,391 32.3%
FT 400 216 54.0% 1,769 813 46.0%
RL 1,289 1,011 78.4% 9,713 8,295 85.4%
We apply the most commonly used metrics, unlearning accuracy (UA), retain accuracy (RA), test accuracy (TA) and MIA, to evaluate three classic machine unlearning methods: Retrain (RT), Finetune (FT) (Warnecke et al., 2021), and Random Label (RL) (Graves et al., 2021). These methods are implemented on a ResNet-18 model trained with the CIFAR-10 dataset in a random data forgetting scenario. In Table 2, compared with RA on $\mathcal{D}_{r}$, the UA of the three methods proves that the model indeed forgets some forget data. The MIA column also indicates that many data points seem to be forgotten. These metrics appear to provide valuable insights into the extent of forgetting achieved. However, higher unlearning accuracy cannot ensure that these forget data points are no longer represented in any form within the model’s predictions.
To validate our hypothesis, we employ the process of conformal prediction to examine whether the true labels of the forget data appear within the prediction set. We set the confidence level to $95\%$, i.e., $\alpha=0.05$, and set the calibration set size to $2,000$. In Table 3, we count both UA metric and MIA metric’s misclassified data points (mis-label) and, after applying conformal prediction, determined how many of these points fall within the conformal prediction set (in-set). Table 3 demonstrates that even though the model misclassifies part of the forget data, on average 54.6% of these misclassified instances are still present in the conformal prediction set. This indicates that a high UA does not mean the model has truly forgotten the data. Relying solely on UA to evaluate the forgetting quality is insufficient.
In MIA, the prediction of ‘0’ indicates a data point is predicted as “truly forget”, while ‘1’ signifies “still retain”. And “mis-label” refers to the number of the model predicts the forget data points as ‘0’. In conformal prediction of MIA, there are three potential prediction set results, i.e., $\{0\},\{1\},\{0,1\}$. The “in-set” here refers to the number of mislabeled data points whose conformal prediction set includes $1$. Thus, the in-set ratio indicates that, although the MIA fails to identify an average of 18.33% of the forget data points, conformal prediction can still recognize more than half of these forget data points within its prediction set.
Overall, high in-set ratios in conformal prediction highlight that a considerable portion of the forget data remains identifiable through conformal prediction. This implies these unlearning methods may not be fully effective in eliminating traces of the forget data from another perspective.
2.3 Designing Metrics Motivated by Conformal Prediction
Based on the limitations of UA and MIA metrics shown in Section 2.2, it raises a question as follows:
(Q2) Can we develop metrics to address the shortcomings of UA and MIA?
Thus, we propose enhanced UA and MIA metrics that draw intuition from conformal prediction.
2.3.1 Definition of New Metrics
Conformal Ratio (CR).
To overcome the limitations of UA, we introduce a novel metric, Conformal Ratio (CR), which incorporates both coverage and set size in conformal prediction to provide a more comprehensive evaluation. Before defining CR, we introduce Coverage and Set Size.
The definition of Coverage, given a dataset $\mathcal{D}$, is as follows:
$\text{Coverage}:=\frac{1}{|\mathcal{D}|}\sum_{({\bm{x}},y_{t})\in\mathcal{D}}\mathbb{I}(y_{t}\in{\mathbb{C}}({\bm{x}})),$ (3)
where $y_{t}$ is the true label of data point ${\bm{x}}$. Coverage reflects the probability that the true label falls within the prediction set ${\mathbb{C}}({\bm{x}})$. For $\mathcal{D}=\mathcal{D}_{f}$, high coverage indicates that the model retains significant information about forget data, suggesting incomplete unlearning. Conversely, higher coverage is better for test data $\mathcal{D}_{test}$.
Given a dataset $\mathcal{D}$, Set Size can be defined as follows:
$\text{Set Size}:=\frac{1}{|\mathcal{D}|}\sum_{({\bm{x}},y_{t})\in\mathcal{D}}|{\mathbb{C}}({\bm{x}})|,$ (4)
where $|{\mathbb{C}}({\bm{x}})|$ represents the prediction set size of a specific instance. When $\mathcal{D}=\mathcal{D}_{f}$, a smaller set size implies that the model cannot forget specific information about $\mathcal{D}_{f}$ well. In contrast, when $\mathcal{D}=\mathcal{D}_{test}$, a small set size may indicate that the model maintains a strong performance after unlearning.
Based on Coverage and Set Size, we introduce the definition of CR for a dataset $\mathcal{D}$ as follows:
$\text{CR}:=\frac{\text{Coverage}}{\text{Set Size}}=\frac{\sum_{({\bm{x}},y_{t})\in\mathcal{D}}\mathbb{I}(y_{t}\in{\mathbb{C}}({\bm{x}}))}{\sum_{({\bm{x}},y_{t})\in\mathcal{D}}|{\mathbb{C}}({\bm{x}})|}.$ (5)
CR balances the information captured by coverage and set size. A lower CR value implies successful unlearning on forget data $\mathcal{D}_{f}$, but a significant drop in the model’s utility for $\mathcal{D}_{test}$. CR is inspired by conformal prediction which is proposed to assess the model’s behavior on new, unseen data (e.g., validation or test data), not on the training data. Thus, we emphasize that CR only measures forget data $\mathcal{D}_{f}$ and test data $\mathcal{D}_{test}$.
MIA Conformal Ratio (MIACR).
To address the limitations of the current MIA metric, we propose a new metric called MIACR. For three potential prediction sets {0}, {1} and {0, 1}, set {0} is the only ideal case for MIA prediction since ‘1’ represents the data point is part of the training data. Therefore, we introduce a new metric to measure the probability of the set {0} occurring, referred to as MIACR:
$\text{MIACR}:=\frac{1}{|\mathcal{D}_{f}|}\sum_{({\bm{x}},y_{t})\in\mathcal{D}_{f}}\mathbb{I}({\mathbb{C}}({\bm{x}})=\{0\}),$ (6)
where ${\mathbb{C}}({\bm{x}})=\{0\}$ denotes prediction set is $\{0\}$. MIACR only calculates the average of ${\mathbb{C}}({\bm{x}})=\{0\}$, and a higher MIACR value represents better forgetting for $\mathcal{D}_{f}$.
Superiority of Our Metrics.
Unlike existing metrics, both the CR and MIACR metrics incorporate the fact that the true labels of some forgotten data points may remain within the prediction set, even when the predictions deviate from the true labels. This allows for a more robust assessment of the model’s forgetting quality and predictive performance.
Evaluation Criteria.
Similar to existing unlearning metrics, we consider two different criteria to measure unlearning performance with our metrics. ➊ Gap to RT Criterion: A lower gap to RT method is better for both CR and MIACR metrics. The performance gap relative to the RT method is represented in (). ➋ Limit-Based Criterion: For the CR metric, a lower CR value of forget data $\mathcal{D}_{f}$ indicates better unlearning performance while a higher CR value of $\mathcal{D}_{test}$ is desirable. For the MIACR metric, a higher MIACR value for forget data ($\mathcal{D}_{f}$) reflects improved unlearning effectiveness.
2.3.2 Confidence Level and Calibration Data
Figure 1: The stability of $\hat{q}$ in different calibration set sizes (results of RT method on CIFAR-10 with ResNet-18). When the calibration set size is greater than $2,000$, the fluctuation of $\hat{q}$ can be kept within a certain range.
In conformal prediction, miscoverage rate $\alpha$ and calibration set size are two crucial parameters for determining prediction set coverage and reliability. We next discuss the suitable settings for these two parameters and the rationale behind our chosen configuration.
In terms of $\alpha$, a higher confidence level (which is $1-\alpha$) provides more reliable coverage but typically results in larger prediction sets. Common settings of confidence level are $0.95$ and $0.9$, corresponding to $\alpha$ values of $0.05$ and $0.1$, respectively. To make our work more broadly applicable and provide a comprehensive reference, we expand our experiments by setting $\alpha$ to a wider range of values: $0.05$, $0.1$, $0.15$, and $0.2$. Notably, our default and recommended value for $\bm{\alpha}$ is $\bm{0.1}$ and subsequent experimental analysis is based on $\bm{\alpha=0.1}$. Higher values of $\alpha$ would result in smaller, potentially more precise prediction sets, but this reduction comes at the cost of decreased confidence, which may not be acceptable in many applications.
Table 4: Unlearning performance of 9 unlearning methods on CIFAR-10 with ResNet-18 in $10\%$ random data forgetting scenario. The sign $\uparrow$ represents the greater is better, while $\downarrow$ denotes ideally small. The results are average values from $3$ independent trials and the standard deviation values are reported in Appendix B. The performance gap relative to the RT method is represented in () in results tables. It shows the unlearning methods that excel under traditional metrics (UA, RA and TA) do not necessarily perform well under the CR metric.
Coverage Set Size CR
Methods $\alpha$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{test}\uparrow$ $\mathcal{D}_{f}\uparrow$ $\mathcal{D}_{test}\downarrow$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{test}\uparrow$ $\hat{q}$
RT 0.1 $0.881({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.895({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.934({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.947({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.943({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.945({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.192$
[…]
FT 0.1 $0.968({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.087})$ $0.899({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $0.969({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.035})$ $0.924({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.023})$ $0.998({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.972({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$ $0.079$
[…]
RL 0.1 $0.913({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.897({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.975({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.041})$ $0.980({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.033})$ $0.936({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.916({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.572$
[…]
GA 0.1 $0.990({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.109})$ $0.905({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.990({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.056})$ $0.928({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $0.998({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.973({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.028})$ $0.062$
[…]
Teacher 0.1 $0.967({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$ $0.898({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.963({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.929({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.018})$ $0.998({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.969({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $0.591$
[…]
FF 0.1 $0.933({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.899({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $7.129({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.195})$ $6.566({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.619})$ $0.131({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.812})$ $0.137({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.808})$ $0.998$
[…]
SSD 0.1 $0.987({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.106})$ $0.902({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.990({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.056})$ $0.926({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $0.998({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.973({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.028})$ $0.063$
[…]
NegGrad+ 0.1 $0.895({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.014})$ $0.898({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.964({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.030})$ $0.950({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.928({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.946({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.044$
[…]
Salun 0.1 $0.936({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.896({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.956({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.022})$ $0.954({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.979({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.939({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.489$
[…]
As for calibration data, it should be independent of train and test data. One requirement for the choice of the calibration set size is that the distribution of the calibration set should match the distribution of the test data as closely as possible. That means a small sample size may lead to distributional dissimilarity and unstable coverage estimation. Therefore, a sufficient sample size of calibration data must be ensured to obtain stable estimates. Figure 1 illustrates the stability of $\hat{q}$ across varying calibration set sizes. We implement them on CIFAR-10 with ResNet-18 in $10\%$ and $50\%$ random data forgetting scenarios. The results shown in Figure 1 are smoothed using B-spline. It reveals that $2,000$ is a sufficient calibration set size for different settings on CIFAR-10 with ResNet-18 to get a stable threshold $\hat{q}$. We also analyze the calibration set size choice of the class-wise forgetting scenario and find that $50$ data points per class as a calibration set is enough. The reason is that, the model’s prediction distribution becomes more complex in a dataset with more classes, requiring a larger calibration set to stabilize the estimation.
2.4 Evaluating Unlearning Methods with New Metrics
Having established the new metrics CR and MIACR, a natural question to ask is:
(Q3) Do existing machine unlearning methods truly achieve forgetting, and how do they perform under our new evaluation metrics?
In this section, we assess the performance of various unlearning methods using the newly introduced metrics, specifically CR, together with coverage and set size. We employ 9 different unlearning methods, RT, FT (Warnecke et al., 2021), RL (Graves et al., 2021), Gradient Ascent (GA) (Thudi et al., 2022), Teacher (Tarun et al., 2023), FisherForgetting (FF) (Golatkar et al., 2020a), SSD (Foster et al., 2024), NegGrad+ (Kurmanji et al., 2024) and Salun (Fan et al., 2023).
The experimental results are presented in Table 4, which summarizes the unlearning performance under $10\%$ random data forgetting scenario on CIFAR-10 with ResNet-18. For more experimental results on CR and MIACR, please check Section 4 and Appendix B. The table 4 includes coverage percentages, set size, CR for both $\mathcal{D}_{f}$ and $\mathcal{D}_{test}$, as well as $\hat{q}$ in different settings of $\alpha$. As observed, it shows a decreasing coverage and set size as the alpha increases in all methods since the threshold $\hat{q}$ is relaxed due to the reduced confidence level.
Based on evaluation criterion ➊, while NegGrad+ outperforms RL under the traditional UA metric, RL exhibits superior performance under the CR metric. This observation suggests that methods excelling in traditional UA metric may not perform well under the CR metric. The underlying rationale behind this is that the CR metric takes into account the possibility that the true labels of some misclassified forget data points may still remain within the prediction set. This observation aligns with the insights we discussed in Section 2.2 regarding the limitation of UA. Based on evaluation criterion ➋, NegGrad+ and RL exhibit the best trade-off between forgetting quality and model performance. FF achieves the lowest CR, indicating the highest forgetting quality, but it exhibits the worst prediction performance on $\mathcal{D}_{test}$. On the other hand, GA, Teacher, SSD, FT and Salun maintain relatively high CR of the forget data $\mathcal{D}_{f}$ across different alpha settings, reflecting their poor forgetting quality.
3 Enhancing Machine Unlearning via Conformal Prediction
Our analysis indicates that machine unlearning methods generally do not exhibit strong performance under our newly developed evaluation metrics. Even methods with relatively low CR values demonstrate significant compromises in other metrics. This observation raises an important question:
(Q4) Can we explore advanced learning techniques via conformal prediction to optimize unlearning model’s performance under our evaluation metrics?
A training-based machine unlearning method is typically optimized for prediction loss but cannot efficiently support the improvement of forgetting quality in our new metrics. Therefore, we propose a novel and general unlearning framework for training-based machine unlearning methods to enhance machine unlearning. The key innovation of this framework lies in leveraging conformal prediction principles combined with a loss function inspired by the C&W attack (Carlini & Wagner, 2017) to improve forgetting quality.
The original C&W loss function is designed to find adversarial examples that cause misclassification. Here we extend it to an unlearning loss. For forget data $\mathcal{D}_{f}$, the objective of the unlearning loss function is to decrease the model’s confidence in the true labels of $\mathcal{D}_{f}$. Based on this, the C&W-inspired unlearning loss can be defined as:
${\mathcal{L}}_{\text{cw}}({\bm{x}},y_{t})=\max\left(p_{t}({\bm{x}})-\max_{i\neq t}\{p_{i}({\bm{x}})\},\ -\Delta\right),$ (7)
where $({\bm{x}},y_{t})\in\mathcal{D}_{f}$. $p_{i}({\bm{x}})$ is the possibility (after softmax function) of class $y_{i}$, while $z_{t}({\bm{x}})$ is the true label $y_{t}$’s possibility output. We denote $\max_{i\neq t}\{p_{i}({\bm{x}})\}$ as the highest possibility value of the incorrect class. $\Delta$ is the margin parameter (also known as the confidence parameter) that controls the attack strength.
This loss is minimized when $p_{t}({\bm{x}})-\max_{i\neq t}\{p_{i}({\bm{x}})\}\leq-\Delta$ to make $\max_{i\neq t}\{p_{i}({\bm{x}})\}-p_{t}({\bm{x}})\geq\Delta$. It means that the loss maximizes the difference between the highest possibility value for class $y_{i}$ ($i\neq t$) and the possibility value for the true label $y_{t}$, which makes it more difficult for the model to include the true label in the prediction set during conformal prediction process. Additionally, increasing the value of $\Delta$ further reduces the confidence of the true label $y_{t}$ being included in the prediction set. In our work, we set $\Delta=0.01$.
Based on the insight discussed in Section 2, we further improve the C&W-inspired unlearning loss function by combining conformal prediction. In conformal prediction, calibration data helps in estimating non-conformity scores and determining a threshold to ensure valid statistical guarantees about the model’s uncertainty estimates. Therefore, we reserve a portion of calibration data $\mathcal{D}_{c}^{\prime}$ for the unlearning phase, which is kept separate from the calibration data $\mathcal{D}_{c}$ used in the evaluation phase.
According to Eq. 1, calibration data $\mathcal{D}_{c}^{\prime}$ is used to calculate non-conformity scores and generate a threshold $\bar{q}$ based on an $\alpha$. Then, we calculate the non-conformity scores of $\mathcal{D}_{f}$ to obtain the corresponding prediction set. Then, by revising C&W-inspired unlearning loss with this calibration step, a general unlearning loss function is defined as follows:
${\mathcal{L}}_{\text{unlearn}}({\bm{x}},y_{t})=\max\left(S({\bm{x}},y_{t})-\bar{q},\ -\Delta\right),$ (8)
where $({\bm{x}},y_{t})\in\mathcal{D}_{f}$. We replace $p_{t}({\bm{x}})$ in Eq. 7 with the non-conformity score $S({\bm{x}},y_{t})$ of true label $y_{t}$, and replace $\max_{i\neq t}\{p_{i}({\bm{x}})\}$ with $\bar{q}$ which will be updated before each epoch. It ensures that ${\mathcal{L}}_{\text{unlearn}}$ adheres to the principle of ${\mathcal{L}}_{\text{cw}}$, which encourages $\bar{q}-S({\bm{x}},y_{t})\geq\Delta$.
It is a general framework applicable across various training-based machine unlearning methods. To preserve the effectiveness of specific machine unlearning methods, we retain their original loss ${\mathcal{L}}_{\text{original}}$. Consequently, we sum these terms to form the final objective loss function as follows:
${\mathcal{L}}_{\text{total}}={\mathcal{L}}_{\text{original}}+\lambda\cdot{\mathcal{L}}_{\text{unlearn}},$ (9)
where $\lambda$ is a hyperparameter that controls the forgetting degree.
4 Experiment
4.1 Experimental setting
Datasets and Models.
We report experiments on CIFAR-10 (Krizhevsky, 2009) with ResNet-18 (He et al., 2016) and Tiny ImageNet (Le & Yang, 2015) with ViT (Dosovitskiy et al., 2021).
Table 5: Performance of 8 unlearning methods on Tiny ImageNet with ViT in $10\%$ random data forgetting scenario.
Coverage Set Size CR
Methods $\alpha$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{test}\uparrow$ $\mathcal{D}_{f}\uparrow$ $\mathcal{D}_{test}\downarrow$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{test}\uparrow$ $\hat{q}$
RT 0.1 $0.892({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.900({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.151({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.144({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.775({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.786({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.853$
[…]
FT 0.1 $0.978({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$ $0.903({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $1.234({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.083})$ $1.317({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.173})$ $0.792({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.017})$ $0.685({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.101})$ $0.935$
[…]
RL 0.1 $0.892({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.902({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $2.639({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.488})$ $1.843({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.699})$ $0.338({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.437})$ $0.489({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.297})$ $0.971$
[…]
GA 0.1 $0.986({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.094})$ $0.900({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.104({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$ $1.224({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.080})$ $0.894({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.119})$ $0.736({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.050})$ $0.899$
[…]
Teacher 0.1 $0.930({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.038})$ $0.902({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $1.991({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.840})$ $1.959({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.815})$ $0.467({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.308})$ $0.460({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.326})$ $0.971$
[…]
SSD 0.1 $0.993({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.101})$ $0.897({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $1.039({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.112})$ $1.134({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.956({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.181})$ $0.791({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.852$
[…]
NegGrad+ 0.1 $0.995({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.103})$ $0.848({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.898({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.253})$ $1.093({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.051})$ $1.225({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.450})$ $1.287({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.501})$ $0.933$
[…]
Salun 0.1 $0.977({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.085})$ $0.924({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $1.229({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.078})$ $1.281({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.137})$ $0.918({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.143})$ $0.884({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.097})$ $0.939$
[…]
Table 6: MIACR performance on CIFAR-10 with ResNet-18. We show the MIA value in $10\%$ random data forgetting scenario, while the MIA results in $50\%$ scenario are presented in Appendix B.
$10\%$ Forgetting $50\%$ Forgetting
Methods $\alpha$ MIACR $\uparrow$ $\hat{q}$ MIACR $\uparrow$ $\hat{q}$
RT 0.1 $0.147({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.589$ $0.201({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.570$
[…]
FT 0.1 $0.077({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.070})$ $0.627$ $0.103({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.098})$ $0.558$
[…]
RL 0.1 $0.178({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$ $0.572$ $0.137({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.064})$ $0.547$
[…]
GA 0.1 $0.032({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.115})$ $0.502$ $0.055({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.146})$ $0.486$
[…]
Teacher 0.1 $0.038({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.109})$ $0.672$ $0.065({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.136})$ $0.582$
[…]
FF 0.1 $0.051({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.096})$ $0.486$ $0.089({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.112})$ $0.509$
[…]
SSD 0.1 $0.031({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.116})$ $0.511$ $0.051({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.150})$ $0.488$
[…]
NegGrad+ 0.1 $0.128({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $0.481$ $0.109({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.092})$ $0.511$
[…]
Salun 0.1 $0.113({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.034})$ $0.681$ $0.115({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$ $0.630$
[…]
Implementation Details.
For CIFAR-10/Tiny ImageNet, we randomly select $200$/$50$ data points per class ($2,000$/$10,000$ data points in total) as calibration data $\mathcal{D}_{c}$ and $\mathcal{D}_{c}^{\prime}$, respectively. The calibration data $\mathcal{D}_{c}$ does not participate in the model training or unlearning processes and is only used for calibrating the threshold $\hat{q}$, while $\mathcal{D}_{c}^{\prime}$ is used in the process of our unlearning framework to generate $\bar{q}$.
For the hyperparameter in our work, miscoverage rate $\alpha\in[0.05,0.1,0.15,0.2]$, margin parameter $\Delta=0.01$, unlearning loss weight $\lambda\in[0,0.2,0.5,1]$. Due to space limitations, more training details and experimental results under settings of $\alpha\in[0.05,0.15]$, $\lambda=0.2$ and class-wise forgetting scenario can be found in Appendix A-C.
4.2 Measure Machine Unlearning Method via New Metrics
CR Metric.
For the CR metric, we already show the results of CIFAR-10 with ResNet-18 in Tabel 4. Here, we further evaluate the unlearning performance in Tiny ImageNet with ViT under $10\%$ random data forgetting scenario by CR metric. The results of Tiny ImageNet with ViT and CIFAR-10 with ReNet18 in $50\%$ random data forgetting and class-wise forgetting scenarios can be found in Appendix B. We remark that it is computationally intractable for the FF method in Tiny ImageNet with ViT, so we omit its results. The results of Tiny ImageNet with ViT in $10\%$ random data forgetting scenario are summarized in Table 5. Our findings reveal the following insights:
For all machine unlearning methods, as $\alpha$ level increases, it results in reduced Coverage and smaller Set Size. This happens because a higher $\alpha$ loosens the conformal threshold $\hat{q}$, allowing fewer predictions to be included within the prediction set for each data point. On the contrary, the CR tends to increase with increasing $\alpha$. Although both Coverage and Set Size may decrease, Set Size often decreases more significantly. Consequently, the CR value generally becomes larger as $\alpha$ increases.
When $\alpha$ is set to $0.2$, most methods show a value of Set Size less than $1$ in both Table 4 and 5. The intuition behind it is that conformal prediction, as a static predictor, is intrinsically tied to the model’s base prediction performance and accuracy. When the model’s accuracy is significantly higher than the confidence level, conformal prediction can achieve the required coverage with ease. In fact, it can generate empty prediction sets for some data points while still meeting the target coverage. Thus, the choice of $\alpha$ is crucial. Overly high $\alpha$ values may skew evaluation results by failing to let CR accurately reflect model performance.
Notably, when $\alpha=0.2$, RL’s coverage on the forget data $\mathcal{D}_{f}$ decreases significantly compared to other methods. This results from the use of label corruption during RL’s unlearning process, which enlarges the distribution difference between calibration data and forget data. We can mitigate this gap if we know the details of specific unlearning methods, and a potential solution to this issue is discussed in Appendix B.
Regarding the forgetting quality of specific machine unlearning methods, the Teacher method achieves the strongest performance on the traditional UA metric under both evaluation criteria ➊ and ➋. However, its forgetting quality is significantly weaker when measured by the CR metric. This highlights that the CR captures critical scenarios overlooked by UA, specifically the potential retention of true labels within prediction sets for the forget data points. By explicitly addressing this phenomenon, CR ensures a more robust and reliable evaluation for unlearning quality.
Table 7: Performance of our unlearning framework. We show the performance on CIFAR-10 with ResNet-18 and Tiny ImageNet with ViT in $10\%$ random data forgetting scenario. $\lambda=0$ represents the baseline without our framework applied. It shows our framework significantly improves the forgetting quality, while preserving stable predictive performance.
$\lambda=0$ $\lambda=0.2$ $\lambda=0.5$
[…]
CIFAR-10 with ResNet-18
RT $0.1$ $8.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $99.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $91.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $0.943({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.945({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $10.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.2})$ $98.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.4})$ $91.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.8})$ $0.914({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.924({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $14.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.4})$ $97.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.9})$ $90.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.4})$ $0.879({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.064})$ $0.912({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.033})$
$0.2$ $0.988({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.981({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.977({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.976({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.963({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.025})$ $0.966({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$
FT $0.1$ $3.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.8})$ $98.1\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.6})$ $91.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.2})$ $0.998({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.972({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$ $6.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.8})$ $97.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.7})$ $90.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.0})$ $0.948({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.924({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $7.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.7})$ $96.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.8})$ $90.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $0.940({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.927({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.018})$
$0.2$ $1.000({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.993({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.989({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.974({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.983({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.975({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$
RL $0.1$ $7.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.0})$ $97.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.3})$ $90.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.2})$ $0.936({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.916({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $9.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.1})$ $96.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.1})$ $89.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.896({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$ $0.887({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.058})$ $9.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.3})$ $96.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.8})$ $89.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.1})$ $0.902({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.041})$ $0.896({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.049})$
$0.2$ $0.976({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.959({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.022})$ $0.964({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $0.949({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.959({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.950({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$
Tiny ImageNet with ViT
RT $0.1$ $14.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $98.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $86.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $0.775({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.786({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $19.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.6})$ $98.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $86.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $0.729({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.163})$ $0.786({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.114})$ $26.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}11.7})$ $98.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.1})$ $85.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.2})$ $0.649({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.243})$ $0.765({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.135})$
$0.2$ $0.934({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.935({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.898({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.108})$ $0.932({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.133})$ $0.839({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.049})$ $0.929({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.130})$
FT $0.1$ $6.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}7.8})$ $97.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $84.1\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.9})$ $0.791({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.087})$ $0.685({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.201})$ $9.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.9})$ $97.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.4})$ $83.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.753({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.139})$ $0.683({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.217})$ $13.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $97.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.6})$ $83.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.718({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.174})$ $0.683({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.217})$
$0.2$ $0.969({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.905({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.942({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.152})$ $0.893({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.094})$ $0.914({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.124})$ $0.890({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.091})$
RL $0.1$ $26.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}12.2})$ $96.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.8})$ $81.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.6})$ $0.338({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.540})$ $0.488({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.398})$ $31.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}17.1})$ $95.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}17.9})$ $80.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.1})$ $0.278({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.614})$ $0.451({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.449})$ $36.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}21.5})$ $95.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.5})$ $80.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.6})$ $0.254({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.638})$ $0.449({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.451})$
$0.2$ $0.821({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.143})$ $0.848({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.110})$ $0.752({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.038})$ $0.825({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.026})$ $0.718({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.072})$ $0.827({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.028})$
MIACR Metric.
In Table 6, we show the MIACR results on CIFAR-10 under both $10\%$ and $50\%$ random data forgetting scenarios. Compared to MIACR, traditional MIA remains a privacy concern. Albeit MIA cannot predict some forget data points as training data, it still has high confidence in appearing in the prediction set. For example, under evaluation criterion ➊, the teacher’s MIA is the method closest to the RT method, but it performs poorly in the CR metric. Similarly, FF and Salun are the top two unlearning methods in the traditional MIA metric (ideally small) under evaluation criterion ➋, but FF has low MIACR (ideally greater) values in both 10% and 50% forgetting scenarios.
4.3 Performance of Our Unlearning Framework
In this experiment, we apply the RT, FT, and RL methods to our framework. Table 7 presents the results for CIFAR-10 with ResNet-18 and Tiny ImageNet with ViT in $10\%$ random data forgetting scenario.
[…]
From the perspective of evaluation criterion ➊, the FT method achieves an average gap of $1.47$ across UA, RA, and TA on ResNet-18 when $\lambda=0.5$, compared to an average gap of $2.2$ when $\lambda=0$. Similarly, on the ViT model, FT reduces the average gap from $3.53$ to $1.63$ when $\lambda=0.5$. If UA of an unlearning method is already lower than that of RT, e.g., RL, our framework cannot further improve their performance under evaluation criterion ➊.
[…]
5 Related Work
Machine unlearning has emerged as a vital research topic due to several privacy, regulatory, and ethical concerns associated with machine learning models. It refers to the process of selectively removing specific data points from a trained machine-learning model. Generally, post-hoc machine unlearning can be divided into training-based (Graves et al., 2021; Warnecke et al., 2021; Thudi et al., 2022; Tarun et al., 2023) and training-free approaches (Guo et al., 2019; Sekhari et al., 2021; Nguyen et al., 2020; Golatkar et al., 2020b, 2021; Foster et al., 2024).
To evaluate these methods, several unlearning metrics have been proposed, including UA (Brophy & Lowd, 2021; Foster et al., 2024) and MIA (Shokri et al., 2017; Chen et al., 2021). However, these metrics often fail to account for the confidence of the forgetting quality. To address this limitation, we improve it in our work motivated by conformal prediction (Angelopoulos & Bates, 2021), which stands out among uncertainty quantification techniques for its ability to provide well-calibrated, reliable confidence measures. As a generic methodology, conformal prediction can transform the outputs of any black box prediction algorithm into a prediction set. Due to its versatility, many works have specifically designed numerous conformal prediction methods tailored to particular prediction problems (Papadopoulos et al., 2002; Lei & Wasserman, 2014; Romano et al., 2020; Lei et al., 2018).
One work (Becker & Liebig, 2022) has primarily focused on parameter-level uncertainty without fully addressing the broader implications of unlearning on prediction confidence. It assesses the sensitivity of model parameters to the target data through Fisher Information Matrix, but they often rely on computationally intensive operations and may struggle to scale to large models or datasets.
6 Conclusion
Motivated by conformal prediction, we introduce new metrics, CR and MIACR, to enhance the evaluation and reliability of machine unlearning. In addiction, our unlearning framework, which incorporates the adapted C&W loss with conformal prediction, improves unlearning effectiveness. Together, we provide a more rigorous foundation for privacy-preserving machine learning.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Setting Details
For CIFAR-10 with a ResNet-18 architecture, we train the original model from scratch for 200 epochs using SGD with a Cosine Annealing learning rate schedule, starting from an initial learning rate of 0.1. We set momentum to 0.9 and a batch size of 64. For the ViT architecture, we initialize the original model by training a pretrained ViT model for 15 epochs on Tiny ImageNet. We start with a learning rate of 0.001, while other training parameters match those used for ResNet-18. In class-wise forgetting, in each dataset, we designate the same class as the forget data $\mathcal{D}_{f}$ across all unlearning methods, with the remaining classes used as the retain data $\mathcal{D}_{r}$.
Appendix B Evaluating MU methods
B.1 Mis-label Number and In-set Ratios
Table 8: Mis-label number and in-set ratios of UA and MIA metrics for 9 unlearning methods.
$10\%$ Forgetting $50\%$ Forgetting
Methods Mis-label $\uparrow$ In-set $\downarrow$ Ratio $\downarrow$ Mis-label $\uparrow$ In-set $\downarrow$ Ratio $\downarrow$
Mis-label and In-set Ratio of UA
RT 431 132 30.6% 2,745 1,573 57.3%
FT 192 112 58.3% 647 431 66.6%
RL 380 173 45.5% 2,625 1,795 68.4%
GA 30 2 6.7% 150 9 6.0%
Teacher 40 4 10% 400 37 9.3%
FF 2,995 2,751 91.9% 15,083 14,061 93.2%
SSD 25 2 8.0% 116 9 7.8%
NegGrad+ 435 115 26.4% 711 249 35.5%
Salun 185 117 63.2% 1065 695 65.3%
Mis-label and In-set Ratio of MIA
RT 654 209 32.0% 4,303 1,391 32.3%
FT 400 216 54.0% 1,769 813 46.0%
RL 1,289 1,011 78.4% 9,713 8,295 85.4%
GA 60 10 16.7% 284 31 10.9%
Teacher 638 586 91.8% 1,689 895 53.0%
FF 1,424 1,424 100% 5,996 4,850 80.9%
SSD 61 11 18.0% 282 24 8.5%
NegGrad+ 486 106 21.8% 1,545 415 26.9%
Salun 2,121 1,848 87.1% 10,221 9,121 89.2%
Conformal prediction is applied to UA and MIA predictions to determine the number of misclassified data points (mis-label) and the number of these points that fall within the conformal prediction set (in-set) across 9 unlearning methods. We evaluate both the UA and MIA metrics by counting the misclassified data points and calculating how many of them are included in the conformal prediction set. The detailed results are presented in Table 8.
B.2 CR Metric
Tables 9 and 10 show the unlearning performance of seven unlearning methods on CIFAR-10 with ResNet-18 in $10\%$ and $50\%$ random data forgetting scenarios, while Table 11 is the results in class-wise forgetting scenario. Tables 12 and 13 present the unlearning performance on Tiny ImageNet with ResNet-18 in random data forgetting scenario, while Table 11 details the unlearning performance in class-wise forgetting scenario.
For all machine unlearning methods, increasing the $\alpha$ level leads to lower Coverage and smaller Set Size. Conversely, the CR tends to grow as $\alpha$ increases. Most methods yield a Set Size of less than $1$ when $\alpha$ is set too low, since conformal prediction is fundamentally linked to the model’s baseline prediction performance (accuracy). If the model’s accuracy significantly exceeds the confidence level, conformal prediction can easily achieve the required coverage. In fact, it may produce empty prediction sets for certain data points while still satisfying the coverage target. Excessively high $\alpha$ values can distort evaluation results, preventing the CR metric from accurately reflecting the model’s performance. Therefore, the choice of $\alpha$ should be carefully considered in relation to the model’s inherent performance. The more experimental results for random data forgetting scenarios have already been analyzed in detail in Section 4 and will not be reiterated here.
Notably, the insights gained from the random data forgetting scenario can also be extended to the class-wise forgetting scenario. Additionally, in the class-wise scenario, some unlearning methods like RT and RL with UA = $100$ and CR approaching 00 indicate they are truly effective at forgetting the specified class.
B.3 MIACR Metric
Table 15 presents the performance of seven machine unlearning methods on CIFAR-10 in ResNet-18, evaluated with the MIACR metric. In addition to the settings discussed in Section 4, we include results for $\alpha\in[0.05,0.15]$ in Table 15.
(a) Without adjusting.
(b) Semi-shadow with 1 epoch.
(c) Semi-shadow with 3 epoch.
(d) Semi-shadow with 4 epoch.
(e) Semi-shadow with 5 epoch.
(f) Shadow with 5 epoch.
Figure 2: Distribution shifting processing with different strategies. The distribution of calibration data gradually converges with that of forget data.
Figure 3: Non-conformity density of calibration data $\mathcal{D}_{c}$ and forget data $\mathcal{D}_{f}$ without our unlearning framework in CIFAR-10 with ResNet-18 under $10\%$ random data forgetting scenario.
Figure 4: Non-conformity score density of calibration data $\mathcal{D}_{c}$ and forget data $\mathcal{D}_{f}$ with our unlearning framework in CIFAR-10 with ResNet-18 under $10\%$ random data forgetting scenario. Our unlearning framework shifts the distribution of the forget data to the right, demonstrating improved forgetting quality.
B.4 How can we better measure forgetting under distribution shifts
Recall that the coverage value of RL has a relatively large deviation in Table 5 since it employs label corruption in its unlearning strategy which can cause distribution shifts. Here, we introduce how to better measure forgetting under these circumstances.
Figure 2(a) shows the non-conformity score distribution of calibration data $\mathcal{D}_{c}$ and forget data $\mathcal{D}_{f}$ in unlearning model $\theta_{u}$ obtained by RL method in Tiny ImageNet with ViT. It looks like there is a significant discrepancy between the distribution of the forget data and the calibration data.
To align the distribution of $\mathcal{D}_{c}$ with that of $\mathcal{D}_{f}$ and minimize the differences between them, we design a shadow model. To make the explanation clearer and more intuitive, we take RL as an example. In the RL unlearning method, the forget data is assigned random labels. Therefore, we apply the same random labeling process to the calibration data and train a shadow model accordingly. We designed two methods:
  1. 1.
    Shadow model. A shadow model replicates the behavior of forget data $\mathcal{D}_{f}$ throughout the unlearning process. A shadow model is a two-step approach: (1) it firstly trains a shadow original model $\theta_{o}^{\prime}$ using train data $\mathcal{D}_{train}$ and clean calibration data $\mathcal{D}_{c}$ with the same epoch number as the original model $\theta_{o}$; (2) subsequently, we finetune the $\theta_{o}^{\prime}$ using the random labeled calibration data.
  2. 2.
    Semi-shadow model. The semi-shadow model only adopts the second step in the shadow model. It finetunes the original model $\theta_{o}$ with random-labeled calibration data.
The results are presented in Figure 2, where (b)-(e) present the results of the semi-shadow model with different epochs and (f) illustrates the shadow model’s result. Under the semi-shadow model, as the number of epochs increases, the distribution of calibration data gradually moves to the right until it becomes consistent with the distribution of forget data. It also shows that the shadow model demonstrates the best ability to handle distribution shifts compared to the semi-shadow model. However, this comes at the cost of higher computational overhead. Overall, the semi-shadow model offers a balanced trade-off between handling distribution shifts effectively and maintaining lower computational costs.
Appendix C Performance of Our Unlearning Framework
[…]
Furthermore, we conduct an ablation study and analyze the impact of using our unlearning framework. As illustrated in Figures 3 and 4, we compare the density distributions of non-conformity scores for calibration data $\mathcal{D}_{c}$ and forget data $\mathcal{D}_{f}$ under the RT, FT, and RL unlearning methods. We set $\lambda$ to 1. Clearly, a higher non-conformity score for $\mathcal{D}_{f}$ indicates that it is less likely to be included in the conformal prediction set, reflecting more effective forgetting. Comparing Figures 3 and 4, it is evident that after applying our unlearning framework, the distribution of non-conformity scores for forget data shifts noticeably to the right. This demonstrates the effectiveness of our framework in enhancing forgetting quality.
Table 9: Unlearning performance of 9 unlearning methods on CIFAR-10 with ResNet-18 in $10\%$ random data forgetting scenario. The results are reported in the format a±$\pm$±b, where a is the mean and b is the standard deviation from $3$ independent trials. The performance gap relative to Retrain method is represented in ().
Coverage Set Size CR
Methods $\alpha$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{test}\uparrow$ $\mathcal{D}_{f}\uparrow$ $\mathcal{D}_{test}\downarrow$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{test}\uparrow$ $\hat{q}$
RT 0.05 $0.941_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.944_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.089_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.074_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.864_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.879_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.883_{\pm 0.007}$
0.1 $0.881_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.895_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.934_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.947_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.943_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.945_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.192_{\pm 0.001}$
UA8.6%, RA99.7%, TA91.8% 0.15 $0.820_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.839_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.841_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.867_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.975_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.968_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.015_{\pm 0.011}$
0.2 $0.780_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.808_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.789_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.824_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.988_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.981_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.003_{\pm 0.002}$
FT 0.05 $0.994_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.951_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $1.008_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.081})$ $1.026_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.048})$ $0.986_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.122})$ $0.927_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.048})$ $0.721_{\pm 0.045}$
0.1 $0.968_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.087})$ $0.899_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $0.969_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.035})$ $0.924_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.023})$ $0.998_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.972_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$ $0.079_{\pm 0.013}$
UA3.8%, RA98.1%, TA91.6% 0.15 $0.915_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.095})$ $0.848_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.916_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.075})$ $0.860_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.025})$ $0.986_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.018})$ $0.008_{\pm 0.000}$
0.2 $0.861_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.081})$ $0.806_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.861_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.072})$ $0.811_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.013})$ $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.993_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.002_{\pm 0.000}$
RL 0.05 $0.970_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.949_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $1.242_{\pm 0.151}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.153})$ $1.197_{\pm 0.098}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.123})$ $0.788_{\pm 0.089}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.076})$ $0.796_{\pm 0.061}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.083})$ $0.877_{\pm 0.057}$
0.1 $0.913_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.897_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.975_{\pm 0.028}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.041})$ $0.980_{\pm 0.025}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.033})$ $0.936_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.916_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.572_{\pm 0.059}$
UA7.6%, RA97.4%, TA90.6% 0.15 $0.825_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.843_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $0.854_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.013})$ $0.888_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $0.966_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.949_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $0.329_{\pm 0.021}$
0.2 $0.755_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.025})$ $0.798_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.774_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.832_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.976_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.959_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.022})$ $0.234_{\pm 0.028}$
GA 0.05 $0.994_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.945_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $1.002_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.087})$ $1.009_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.065})$ $0.994_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.130})$ $0.936_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.057})$ $0.621_{\pm 0.015}$
0.1 $0.990_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.109})$ $0.905_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.990_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.056})$ $0.928_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $0.998_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.973_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.028})$ $0.062_{\pm 0.016}$
UA0.6%, RA99.5%, TA94.1% 0.15 $0.969_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.149})$ $0.848_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.969_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.128})$ $0.858_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $1.000_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.025})$ $0.986_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.018})$ $0.006_{\pm 0.009}$
0.2 $0.925_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.145})$ $0.805_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.924_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.135})$ $0.811_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.013})$ $0.998_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.992_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.003_{\pm 0.005}$
Teacher 0.05 $0.991_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.050})$ $0.941_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $1.003_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$ $1.021_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.993_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.129})$ $0.922_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.043})$ $0.744_{\pm 0.015}$
0.1 $0.967_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$ $0.898_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.963_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.929_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.018})$ $0.998_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.969_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $0.591_{\pm 0.005}$
UA0.8%, RA99.4%, TA93.5% 0.15 $0.913_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.093})$ $0.845_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.912_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.071})$ $0.859_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.996_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $0.983_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.481_{\pm 0.009}$
0.2 $0.865_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.085})$ $0.806_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.866_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.077})$ $0.816_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.998_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.988_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.426_{\pm 0.007}$
FF 0.05 $0.973_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.949_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $7.966_{\pm 0.212}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.877})$ $7.408_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.334})$ $0.122_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.742})$ $0.128_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.751})$ $0.999_{\pm 0.000}$
0.1 $0.933_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.899_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $7.129_{\pm 0.148}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.195})$ $6.566_{\pm 0.166}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.619})$ $0.131_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.812})$ $0.137_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.808})$ $0.998_{\pm 0.001}$
UA59.9%, RA40.1%, TA41.1% 0.15 $0.888_{\pm 0.029}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.068})$ $0.852_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.013})$ $6.431_{\pm 0.078}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.590})$ $5.903_{\pm 0.262}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.036})$ $0.138_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.837})$ $0.144_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.824})$ $0.996_{\pm 0.002}$
0.2 $0.835_{\pm 0.048}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.794_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.014})$ $5.750_{\pm 0.034}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.961})$ $5.219_{\pm 0.368}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.395})$ $0.145_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.843})$ $0.153_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.828})$ $0.993_{\pm 0.003}$
SSD 0.05 $0.996_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.945_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.999_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.090})$ $1.008_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.066})$ $0.994_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.130})$ $0.936_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.057})$ $0.622_{\pm 0.019}$
0.1 $0.987_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.106})$ $0.902_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.990_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.056})$ $0.926_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $0.998_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.973_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.028})$ $0.063_{\pm 0.022}$
UA0.5%, RA99.5%, TA94.2% 0.15 $0.967_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.147})$ $0.849_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.965_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.124})$ $0.862_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $1.002_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$ $0.990_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.022})$ $0.007_{\pm 0.007}$
0.2 $0.922_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.142})$ $0.803_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.923_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.134})$ $0.811_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.013})$ $1.002_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.014})$ $0.992_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.001_{\pm 0.005}$
NegGrad+ 0.05 $0.934_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.948_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $1.068_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $1.086_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.875_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.873_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.989_{\pm 0.013}$
0.1 $0.895_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.014})$ $0.898_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.964_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.030})$ $0.950_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.928_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.946_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.044_{\pm 0.041}$
UA8.7%, RA98.8%, TA92.2% 0.15 $0.851_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$ $0.851_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.896_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.876_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.950_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.025})$ $0.971_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.000_{\pm 0.000}$
0.2 $0.800_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.020})$ $0.799_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.832_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.043})$ $0.813_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.961_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$ $0.983_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.000_{\pm 0.000}$
Salun 0.05 $0.987_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.046})$ $0.950_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $1.132_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.043})$ $1.143_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.069})$ $0.872_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.832_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$ $0.867_{\pm 0.001}$
0.1 $0.936_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.896_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.956_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.022})$ $0.954_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.979_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.939_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.489_{\pm 0.029}$
UA3.7%, RA98.9%, TA91.8% 0.15 $0.871_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.051})$ $0.849_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.881_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.040})$ $0.886_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $0.989_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.014})$ $0.958_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.314_{\pm 0.020}$
0.2 $0.788_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.794_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.014})$ $0.794_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.821_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.992_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $0.966_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.221_{\pm 0.005}$
Table 10: Unlearning performance of 9 unlearning methods on CIFAR-10 with ResNet18 in $50\%$ random data forgetting scenario.
Coverage Set Size CR
Methods $\alpha$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{test}\uparrow$ $\mathcal{D}_{f}\uparrow$ $\mathcal{D}_{test}\downarrow$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{test}\uparrow$ $\hat{q}$
RT $0.05$ $0.955_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.947_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.287_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.214_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.742_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.780_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.984_{\pm 0.002}$
$0.1$ $0.898_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.904_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.023_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.021_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.878_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.886_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.650_{\pm 0.004}$
UA11.0%, RA99.8%, TA89.2% $0.15$ $0.833_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.847_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.883_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.906_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.943_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.934_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.090_{\pm 0.004}$
$0.2$ $0.782_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.814_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.812_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.850_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.964_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.958_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.018_{\pm 0.006}$
FT $0.05$ $0.996_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.041})$ $0.952_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $1.007_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.280})$ $1.029_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.185})$ $0.989_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.247})$ $0.925_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.145})$ $0.738_{\pm 0.014}$
$0.1$ $0.975_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.077})$ $0.896_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.976_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$ $0.921_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.100})$ $0.999_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.121})$ $0.972_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$ $0.081_{\pm 0.033}$
UA2.6%, RA99.1%, TA91.8% $0.15$ $0.936_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.103})$ $0.854_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.936_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.867_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.039})$ $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.057})$ $0.985_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.051})$ $0.011_{\pm 0.002}$
$0.2$ $0.859_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.077})$ $0.790_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $0.859_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$ $0.795_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.993_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.035})$ $0.001_{\pm 0.000}$
RL $0.05$ $0.976_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.022})$ $0.949_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $1.973_{\pm 0.396}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.686})$ $1.971_{\pm 0.406}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.757})$ $0.508_{\pm 0.100}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.234})$ $0.495_{\pm 0.098}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.285})$ $0.899_{\pm 0.012}$
$0.1$ $0.942_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.043})$ $0.907_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $1.227_{\pm 0.103}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.204})$ $1.235_{\pm 0.107}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.214})$ $0.771_{\pm 0.064}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.107})$ $0.738_{\pm 0.064}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.147})$ $0.837_{\pm 0.016}$
UA10.5%, RA93.9%, TA85.8% $0.15$ $0.891_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.058})$ $0.856_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $1.009_{\pm 0.047}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.125})$ $1.011_{\pm 0.045}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.105})$ $0.884_{\pm 0.039}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.059})$ $0.847_{\pm 0.037}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.087})$ $0.770_{\pm 0.022}$
$0.2$ $0.834_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.051})$ $0.799_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.016})$ $0.897_{\pm 0.026}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$ $0.893_{\pm 0.025}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.043})$ $0.929_{\pm 0.024}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.034})$ $0.895_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.063})$ $0.713_{\pm 0.028}$
GA $0.05$ $0.996_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.041})$ $0.945_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $1.003_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.284})$ $1.005_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.209})$ $1.050_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.308})$ $0.945_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.165})$ $0.616_{\pm 0.008}$
$0.1$ $0.985_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.087})$ $0.902_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.989_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.034})$ $0.926_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.095})$ $1.095_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.217})$ $0.916_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.030})$ $0.057_{\pm 0.005}$
UA0.6%, RA99.5%, TA94.3% $0.15$ $0.966_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.133})$ $0.848_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.966_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.083})$ $0.857_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.049})$ $1.141_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.198})$ $0.879_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.005_{\pm 0.007}$
$0.2$ $0.929_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.147})$ $0.809_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.932_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.120})$ $0.817_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.033})$ $1.150_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.186})$ $0.871_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.087})$ $0.001_{\pm 0.007}$
Teacher $0.05$ $0.985_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.030})$ $0.944_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $1.066_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.221})$ $1.143_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.071})$ $0.923_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.181})$ $0.823_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.043})$ $0.857_{\pm 0.013}$
$0.1$ $0.949_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.051})$ $0.909_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.970_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.986_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.035})$ $0.980_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.102})$ $0.918_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.834_{\pm 0.005}$
UA1.6%, RA98.3%, TA91.7% $0.15$ $0.885_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.849_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.894_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.893_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.013})$ $0.992_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.049})$ $0.950_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.016})$ $0.813_{\pm 0.013}$
$0.2$ $0.818_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.798_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.016})$ $0.823_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.826_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $0.997_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.033})$ $0.971_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.013})$ $0.793_{\pm 0.012}$
FF $0.05$ $0.972_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.017})$ $0.954_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $8.023_{\pm 0.189}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.736})$ $7.461_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.247})$ $0.121_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.621})$ $0.128_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.652})$ $0.999_{\pm 0.000}$
$0.1$ $0.930_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.897_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $7.091_{\pm 0.151}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.068})$ $6.521_{\pm 0.171}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.501})$ $0.131_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.747})$ $0.138_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.748})$ $0.998_{\pm 0.001}$
UA60.0%, RA40.1%, TA40.6% $0.15$ $0.887_{\pm 0.024}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.054})$ $0.852_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $6.402_{\pm 0.034}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.519})$ $5.837_{\pm 0.438}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.931})$ $0.139_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.804})$ $0.146_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.788})$ $0.996_{\pm 0.001}$
$0.2$ $0.840_{\pm 0.046}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.058})$ $0.803_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $5.805_{\pm 0.042}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.994})$ $5.253_{\pm 0.376}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.403})$ $0.145_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.819})$ $0.153_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.804})$ $0.993_{\pm 0.003}$
SSD $0.05$ $0.993_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.038})$ $0.944_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.999_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.288})$ $1.001_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.213})$ $0.995_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.253})$ $0.941_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.161})$ $0.585_{\pm 0.014}$
$0.1$ $0.991_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.093})$ $0.904_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.991_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.929_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.092})$ $1.000_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.122})$ $0.975_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.089})$ $0.060_{\pm 0.011}$
UA0.5%, RA99.5%, TA94.3% $0.15$ $0.964_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.131})$ $0.850_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.967_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.084})$ $0.860_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.046})$ $1.000_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.057})$ $0.988_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.054})$ $0.005_{\pm 0.010}$
$0.2$ $0.930_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.148})$ $0.807_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.929_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.117})$ $0.814_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $1.000_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.992_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.034})$ $0.002_{\pm 0.005}$
NegGrad+ $0.05$ $0.986_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$ $0.949_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $1.039_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.248})$ $1.062_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.152})$ $0.949_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.207})$ $0.893_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.113})$ $0.855_{\pm 0.028}$
$0.1$ $0.951_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.903_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.964_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.059})$ $0.944_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.076})$ $0.987_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.109})$ $0.956_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.070})$ $0.177_{\pm 0.055}$
UA2.8%, RA99.6%, TA92.9% $0.15$ $0.889_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.056})$ $0.845_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.892_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.861_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.045})$ $0.996_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.981_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$ $0.012_{\pm 0.002}$
$0.2$ $0.825_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.043})$ $0.796_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.018})$ $0.827_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.805_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.045})$ $0.999_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.035})$ $0.989_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.002_{\pm 0.000}$
Salun $0.05$ $0.988_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.034})$ $0.951_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $1.314_{\pm 0.113}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$ $1.381_{\pm 0.121}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.167})$ $0.756_{\pm 0.064}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.014})$ $0.692_{\pm 0.058}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.088})$ $0.871_{\pm 0.013}$
$0.1$ $0.956_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.058})$ $0.897_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $1.015_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $1.021_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.941_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.064})$ $0.878_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.776_{\pm 0.002}$
UA4.3%, RA97.7%, TA89.4% $0.15$ $0.910_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.078})$ $0.847_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.937_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.054})$ $0.916_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.972_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.924_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.714_{\pm 0.010}$
$0.2$ $0.856_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.074})$ $0.796_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $0.872_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.060})$ $0.844_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.982_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $0.943_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.669_{\pm 0.008}$
Table 11: Unlearning performance of 9 unlearning methods on CIFAR-10 with ResNet18 in class-wise forgetting scenario.
Coverage Set Size CR
Methods $\alpha$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{tf}\downarrow$ $\mathcal{D}_{tr}\uparrow$ $\mathcal{D}_{f}\uparrow$ $\mathcal{D}_{tf}\uparrow$ $\mathcal{D}_{tr}\downarrow$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{tf}\downarrow$ $\mathcal{D}_{tr}\uparrow$ $\hat{q}_{f}$ $\hat{q}_{test}$
RT 0.05 $1.000_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.964_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $10.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $10.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.148_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.100_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.100_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.840_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.000}$ $0.982_{\pm 0.003}$
0.1 $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.882_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $10.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $10.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.922_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.100_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.100_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.956_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.001}$ $0.080_{\pm 0.003}$
UA100%, UAtf100%, 0.15 $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.856_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $10.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $10.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.882_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.100_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.100_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.970_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.000}$ $0.018_{\pm 0.010}$
RA99.9%, TA92.4% 0.2 $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.814_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $10.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $10.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.830_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.100_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.100_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.981_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.000}$ $0.003_{\pm 0.001}$
FT 0.05 $0.994_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.962_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.038})$ $0.944_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.020})$ $9.854_{\pm 0.127}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.146})$ $9.403_{\pm 0.501}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.597})$ $1.045_{\pm 0.040}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.103})$ $0.101_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.102_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.904_{\pm 0.028}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.065})$ $1.000_{\pm 0.000}$ $0.731_{\pm 0.166}$
0.1 $0.969_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$ $0.882_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.118})$ $0.908_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.026})$ $9.495_{\pm 0.255}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.505})$ $8.528_{\pm 0.571}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.472})$ $0.956_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.034})$ $0.102_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.104_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $0.950_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $1.000_{\pm 0.000}$ $0.314_{\pm 0.010}$
UA100%, UAtf100%, 0.15 $0.951_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.049})$ $0.840_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.160})$ $0.851_{\pm 0.031}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $9.265_{\pm 0.279}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.735})$ $8.131_{\pm 0.523}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.869})$ $0.872_{\pm 0.039}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.103_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.103_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.976_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $1.000_{\pm 0.000}$ $0.073_{\pm 0.054}$
RA96.7%, TA90.8% 0.2 $0.942_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.058})$ $0.818_{\pm 0.072}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.182})$ $0.838_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.023})$ $9.163_{\pm 0.245}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.837})$ $7.934_{\pm 0.533}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.066})$ $0.854_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $0.103_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.103_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.981_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.000}$ $0.039_{\pm 0.017}$
RL 0.05 $0.995_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.954_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.046})$ $0.959_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $9.993_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $9.900_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.100})$ $1.170_{\pm 0.155}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.022})$ $0.100_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.096_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $0.828_{\pm 0.097}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $1.000_{\pm 0.000}$ $0.870_{\pm 0.145}$
0.1 $0.984_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.016})$ $0.907_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.093})$ $0.918_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $9.978_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.022})$ $9.800_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.200})$ $0.982_{\pm 0.036}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.059})$ $0.099_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.093_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.936_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $1.000_{\pm 0.000}$ $0.469_{\pm 0.250}$
UA100%, UAtf100%, 0.15 $0.961_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.039})$ $0.859_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.141})$ $0.870_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.014})$ $9.950_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.050})$ $9.700_{\pm 0.066}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.300})$ $0.904_{\pm 0.045}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $0.097_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.089_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.964_{\pm 0.027}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $1.000_{\pm 0.000}$ $0.144_{\pm 0.163}$
RA98.0%, TA92.7% 0.2 $0.935_{\pm 0.027}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.065})$ $0.815_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.185})$ $0.804_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $9.919_{\pm 0.035}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.081})$ $9.637_{\pm 0.076}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.363})$ $0.820_{\pm 0.026}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.094_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.085_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.981_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.999_{\pm 0.001}$ $0.014_{\pm 0.013}$
GA 0.05 $1.000_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.948_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.016})$ $10.000_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $10.000_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.204_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.056})$ $0.100_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.100_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.787_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $1.000_{\pm 0.010}$ $0.988_{\pm 0.000}$
0.1 $1.000_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.899_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.017})$ $10.000_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $10.000_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.005_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.083})$ $0.100_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.100_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.894_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.062})$ $1.000_{\pm 0.000}$ $0.562_{\pm 0.003}$
UA84.6%, UAtf82.5%, 0.15 $1.000_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.843_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.013})$ $10.000_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $10.000_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.893_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.100_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.100_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.944_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.026})$ $1.000_{\pm 0.001}$ $0.051_{\pm 0.002}$
RA96.4%, TA89.6% 0.2 $0.828_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.172})$ $0.782_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.218})$ $0.838_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $9.550_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.450})$ $9.366_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.634})$ $0.884_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.054})$ $0.087_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.013})$ $0.084_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.016})$ $0.948_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.033})$ $1.000_{\pm 0.002}$ $0.038_{\pm 0.003}$
Teacher 0.05 $0.994_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.959_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.041})$ $0.939_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.025})$ $9.877_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.123})$ $9.502_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.498})$ $1.000_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.148})$ $0.101_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.101_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.939_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.099})$ $0.955_{\pm 0.005}$ $0.588_{\pm 0.004}$
0.1 $0.931_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.069})$ $0.904_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.096})$ $0.890_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $9.199_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.801})$ $8.604_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.396})$ $0.914_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.101_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.105_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.974_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.018})$ $0.926_{\pm 0.004}$ $0.116_{\pm 0.005}$
UA90.1%, UAtf86.5%, 0.15 $0.879_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.121})$ $0.881_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.119})$ $0.834_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.022})$ $8.730_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.270})$ $8.081_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.919})$ $0.845_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.037})$ $0.101_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.109_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.986_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.016})$ $0.921_{\pm 0.001}$ $0.017_{\pm 0.002}$
RA99.5%, TA94.0% 0.2 $0.809_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.191})$ $0.841_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.159})$ $0.816_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $8.141_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.859})$ $7.525_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.475})$ $0.824_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.099_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.112_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.990_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.916_{\pm 0.005}$ $0.010_{\pm 0.003}$
FF 0.05 $0.984_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.016})$ $0.962_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.038})$ $0.950_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $7.760_{\pm 1.386}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.240})$ $7.479_{\pm 1.223}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.521})$ $7.627_{\pm 0.058}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.479})$ $0.129_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.130_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.030})$ $0.125_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.715})$ $0.999_{\pm 0.000}$ $0.997_{\pm 0.003}$
0.1 $0.964_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.917_{\pm 0.037}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.083})$ $0.909_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$ $7.360_{\pm 0.999}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.640})$ $6.988_{\pm 0.749}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.012})$ $6.853_{\pm 0.390}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.930})$ $0.132_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.132_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.133_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.823})$ $0.998_{\pm 0.000}$ $0.996_{\pm 0.003}$
UA100%, UAtf100%, 0.15 $0.922_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.078})$ $0.852_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.148})$ $0.863_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $6.941_{\pm 1.088}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.059})$ $6.562_{\pm 0.856}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.438})$ $6.278_{\pm 0.232}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.395})$ $0.134_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.034})$ $0.131_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$ $0.138_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.832})$ $0.997_{\pm 0.001}$ $0.994_{\pm 0.006}$
RA34.3%, TA36.1% 0.2 $0.897_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.103})$ $0.818_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.182})$ $0.789_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.025})$ $6.757_{\pm 0.978}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.243})$ $6.352_{\pm 0.757}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.648})$ $5.307_{\pm 0.527}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.477})$ $0.134_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.034})$ $0.130_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.030})$ $0.150_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.832})$ $0.993_{\pm 0.002}$ $0.993_{\pm 0.006}$
SSD 0.05 $0.995_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.935_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.065})$ $0.940_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $1.030_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}8.970})$ $1.067_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}8.933})$ $0.991_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.157})$ $0.966_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.866})$ $0.876_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.776})$ $0.949_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.109})$ $0.804_{\pm 0.015}$ $0.447_{\pm 0.007}$
0.1 $0.984_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.016})$ $0.910_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.090})$ $0.880_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.992_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}9.008})$ $0.982_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}9.018})$ $0.896_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.026})$ $0.992_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.892})$ $0.926_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.826})$ $0.981_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.025})$ $0.434_{\pm 0.007}$ $0.022_{\pm 0.005}$
UA1.16%, UAtf7.75%, 0.15 $0.960_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.040})$ $0.876_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.124})$ $0.847_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.962_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}9.038})$ $0.931_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}9.069})$ $0.857_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.025})$ $0.998_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.898})$ $0.941_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.841})$ $0.989_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $0.215_{\pm 0.007}$ $0.005_{\pm 0.017}$
RA99.5%, TA94.3% 0.2 $0.895_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.105})$ $0.816_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.184})$ $0.823_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.895_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}9.105})$ $0.850_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}9.150})$ $0.831_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.999_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.899})$ $0.960_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.860})$ $0.991_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.078_{\pm 0.003}$ $0.002_{\pm 0.009}$
Salun 0.05 $0.996_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $0.941_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.059})$ $0.952_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $9.996_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $9.892_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.108})$ $1.028_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.121})$ $0.100_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.095_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.926_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.087})$ $1.000_{\pm 0.000}$ $0.785_{\pm 0.049}$
0.1 $0.988_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.906_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.094})$ $0.901_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.020})$ $9.985_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $9.817_{\pm 0.045}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.183})$ $0.928_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.099_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.092_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.971_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $1.000_{\pm 0.000}$ $0.042_{\pm 0.011}$
UA100%, UAtf100%, 0.15 $0.960_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.040})$ $0.851_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.149})$ $0.878_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.022})$ $9.952_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.048})$ $9.677_{\pm 0.088}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.323})$ $0.896_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.013})$ $0.096_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $0.088_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.980_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $1.000_{\pm 0.000}$ $0.009_{\pm 0.001}$
RA99.6%, TA94.3% 0.2 $0.915_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.085})$ $0.807_{\pm 0.038}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.193})$ $0.820_{\pm 0.035}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $9.893_{\pm 0.024}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.107})$ $9.511_{\pm 0.192}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.489})$ $0.828_{\pm 0.039}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.092_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.085_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.990_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $1.000_{\pm 0.000}$ $0.001_{\pm 0.001}$
Table 12: Unlearning performance of 8 unlearning methods on Tiny ImageNet with ViT in $10\%$ random data forgetting scenario.
Coverage Set Size CR
Methods $\alpha$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{test}\uparrow$ $\mathcal{D}_{f}\uparrow$ $\mathcal{D}_{test}\downarrow$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{test}\uparrow$ $\hat{q}$
RT 0.05 $0.944_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.949_{\pm 0.026}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.876_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.840_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.503_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.516_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.984_{\pm 0.002}$
0.1 $0.892_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.900_{\pm 0.025}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.151_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.144_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.775_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.786_{\pm 0.026}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.853_{\pm 0.003}$
UA14.7%, RA98.8%, TA86.0% 0.15 $0.841_{\pm 0.024}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.850_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.956_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.956_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.880_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.889_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.539_{\pm 0.001}$
0.2 $0.790_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.799_{\pm 0.023}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.846_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.854_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.934_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.935_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.238_{\pm 0.012}$
FT 0.05 $0.994_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.050})$ $0.950_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $2.133_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.257})$ $2.440_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.600})$ $0.466_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.037})$ $0.389_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.127})$ $0.994_{\pm 0.020}$
0.1 $0.978_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$ $0.903_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $1.234_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.083})$ $1.317_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.173})$ $0.792_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.017})$ $0.685_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.101})$ $0.935_{\pm 0.012}$
UA6.9%, RA97.9%, TA84.1% 0.15 $0.938_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.097})$ $0.851_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $1.014_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.058})$ $1.017_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.061})$ $0.925_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.045})$ $0.836_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.681_{\pm 0.003}$
0.2 $0.888_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.098})$ $0.801_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.915_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.069})$ $0.885_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$ $0.970_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.905_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.030})$ $0.326_{\pm 0.011}$
RL 0.05 $0.969_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.025})$ $0.952_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $17.890_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}16.014})$ $8.572_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.732})$ $0.054_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.449})$ $0.111_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.405})$ $0.996_{\pm 0.019}$
0.1 $0.892_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.902_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $2.639_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.488})$ $1.843_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.699})$ $0.338_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.437})$ $0.489_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.297})$ $0.971_{\pm 0.014}$
UA26.9%, RA96.0%, TA81.4% 0.15 $0.793_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.048})$ $0.855_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $1.225_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.269})$ $1.164_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.208})$ $0.648_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.232})$ $0.734_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.155})$ $0.894_{\pm 0.022}$
0.2 $0.681_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.109})$ $0.803_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $0.831_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.946_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.092})$ $0.820_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.114})$ $0.849_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$ $0.715_{\pm 0.013}$
GA 0.05 $0.996_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.947_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $1.539_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.337})$ $2.018_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.178})$ $0.647_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.144})$ $0.469_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$ $0.988_{\pm 0.004}$
0.1 $0.986_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.094})$ $0.900_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.104_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$ $1.224_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.080})$ $0.894_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.119})$ $0.736_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.050})$ $0.899_{\pm 0.001}$
UA3.2%, RA97.4%, TA84.9% 0.15 $0.967_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.126})$ $0.852_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $1.003_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$ $0.993_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.037})$ $0.964_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.084})$ $0.859_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.030})$ $0.632_{\pm 0.009}$
0.2 $0.934_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.144})$ $0.800_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.946_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.100})$ $0.871_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.017})$ $0.987_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.919_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.016})$ $0.296_{\pm 0.009}$
Teacher 0.05 $0.977_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.033})$ $0.956_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $5.473_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.597})$ $5.080_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.240})$ $0.179_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.324})$ $0.188_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.328})$ $0.987_{\pm 0.008}$
0.1 $0.930_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.038})$ $0.902_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $1.991_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.840})$ $1.959_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.815})$ $0.467_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.308})$ $0.460_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.326})$ $0.971_{\pm 0.007}$
UA17.3%, RA86.7%, TA79.0% 0.15 $0.873_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.850_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.295_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.339})$ $1.319_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.363})$ $0.674_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.206})$ $0.645_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.244})$ $0.944_{\pm 0.006}$
0.2 $0.816_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.026})$ $0.803_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $1.020_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.174})$ $1.058_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.204})$ $0.800_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.134})$ $0.758_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.177})$ $0.910_{\pm 0.006}$
SSD 0.05 $0.998_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.054})$ $0.950_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $1.354_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.522})$ $1.827_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.013})$ $0.737_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.234})$ $0.520_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $0.985_{\pm 0.005}$
0.1 $0.993_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.101})$ $0.897_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $1.039_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.112})$ $1.134_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.956_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.181})$ $0.791_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.852_{\pm 0.001}$
UA1.5%, RA98.5%, TA86.1% 0.15 $0.981_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.140})$ $0.853_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.993_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.037})$ $0.962_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.988_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.108})$ $0.887_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.542_{\pm 0.007}$
0.2 $0.956_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.166})$ $0.805_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.960_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.114})$ $0.864_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.996_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.062})$ $0.932_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.249_{\pm 0.006}$
NegGrad+ 0.05 $0.999_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.890_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.059})$ $0.949_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.927})$ $1.614_{\pm 0.023}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.227})$ $2.184_{\pm 0.052}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.681})$ $2.499_{\pm 0.059}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.984})$ $0.995_{\pm 0.000}$
0.1 $0.995_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.103})$ $0.848_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.898_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.253})$ $1.093_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.051})$ $1.225_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.450})$ $1.287_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.501})$ $0.933_{\pm 0.002}$
UA19.4%, RA98.3%, TA84.0% 0.15 $0.987_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.146})$ $0.814_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.850_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.106})$ $1.009_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $1.017_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.137})$ $1.023_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.133})$ $0.685_{\pm 0.002}$
0.2 $0.966_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.176})$ $0.783_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.016})$ $0.802_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.044})$ $0.972_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.118})$ $0.922_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.891_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.043})$ $0.320_{\pm 0.001}$
Salun 0.05 $0.995_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.051})$ $0.964_{\pm 0.026}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $2.803_{\pm 1.607}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.927})$ $2.726_{\pm 0.727}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.886})$ $1.311_{\pm 1.810}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.808})$ $1.157_{\pm 1.481}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.641})$ $0.988_{\pm 0.001}$
0.1 $0.977_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.085})$ $0.924_{\pm 0.040}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $1.229_{\pm 0.286}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.078})$ $1.281_{\pm 0.120}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.137})$ $0.918_{\pm 0.387}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.143})$ $0.884_{\pm 0.374}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.097})$ $0.939_{\pm 0.005}$
UA9.2%, RA97.7%, TA83.6% 0.15 $0.936_{\pm 0.041}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.095})$ $0.874_{\pm 0.041}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $0.972_{\pm 0.103}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.016})$ $1.032_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.076})$ $0.935_{\pm 0.087}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.893_{\pm 0.124}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $0.819_{\pm 0.003}$
0.2 $0.870_{\pm 0.081}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.080})$ $0.810_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.845_{\pm 0.036}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.925_{\pm 0.046}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.071})$ $0.924_{\pm 0.047}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.894_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.041})$ $0.630_{\pm 0.003}$
Table 13: Unlearning performance of 8 unlearning methods on Tiny ImageNet with ViT in $50\%$ random data forgetting scenario.
Coverage Set Size CR
Methods $\alpha$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{test}\uparrow$ $\mathcal{D}_{f}\uparrow$ $\mathcal{D}_{test}\downarrow$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{test}\uparrow$ $\hat{q}$
RT 0.05 $0.946_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.948_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $2.146_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $2.106_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.441_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.450_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.987_{\pm 0.004}$
0.1 $0.892_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.899_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.222_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.211_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.730_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.742_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.889_{\pm 0.009}$
UA16.0%, RA98.8%, TA84.9% 0.15 $0.838_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.847_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.977_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.977_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.858_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.868_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.607_{\pm 0.001}$
0.2 $0.786_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.796_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.856_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.863_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.918_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.922_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.304_{\pm 0.008}$
FT 0.05 $0.995_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.051})$ $0.949_{\pm 0.024}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.879_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $2.216_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.376})$ $0.527_{\pm 0.028}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $0.428_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.088})$ $0.992_{\pm 0.019}$
0.1 $0.979_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.087})$ $0.901_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $1.183_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $1.281_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.137})$ $0.828_{\pm 0.029}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.701_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.085})$ $0.926_{\pm 0.025}$
UA5.4%, RA97.1%, TA84.4% 0.15 $0.953_{\pm 0.024}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.112})$ $0.850_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.014_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.058})$ $1.017_{\pm 0.026}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.061})$ $0.940_{\pm 0.027}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.060})$ $0.839_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.050})$ $0.681_{\pm 0.020}$
0.2 $0.910_{\pm 0.029}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.120})$ $0.806_{\pm 0.024}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.937_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.091})$ $0.895_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.041})$ $0.977_{\pm 0.029}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.043})$ $0.902_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.033})$ $0.345_{\pm 0.016}$
RL 0.05 $0.974_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.028})$ $0.953_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $26.032_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}23.886})$ $23.369_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}21.263})$ $0.038_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.403})$ $0.038_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.412})$ $0.994_{\pm 0.010}$
0.1 $0.930_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.038})$ $0.902_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $5.277_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.055})$ $4.621_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.410})$ $0.178_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.552})$ $0.197_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.545})$ $0.987_{\pm 0.008}$
UA22.5%, RA93.5%, TA77.1% 0.15 $0.875_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.037})$ $0.856_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $1.758_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.781})$ $1.657_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.680})$ $0.496_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.362})$ $0.516_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.352})$ $0.970_{\pm 0.017}$
0.2 $0.810_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $0.805_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $1.147_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.291})$ $1.144_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.281})$ $0.707_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.211})$ $0.707_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.215})$ $0.945_{\pm 0.005}$
GA 0.05 $0.998_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.949_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $1.807_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.339})$ $2.338_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.232})$ $0.552_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.111})$ $0.407_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.043})$ $0.992_{\pm 0.006}$
0.1 $0.986_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.094})$ $0.896_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $1.147_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.075})$ $1.278_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.067})$ $0.863_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.133})$ $0.703_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.039})$ $0.918_{\pm 0.010}$
UA3.9%, RA96.1%, TA84.2% 0.15 $0.968_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.130})$ $0.850_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $1.015_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.038})$ $1.020_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.043})$ $0.954_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.096})$ $0.835_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.033})$ $0.696_{\pm 0.009}$
0.2 $0.931_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.145})$ $0.804_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.948_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.092})$ $0.893_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.030})$ $0.983_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.065})$ $0.900_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.022})$ $0.363_{\pm 0.002}$
Teacher 0.05 $0.967_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $0.950_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $6.465_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.319})$ $6.233_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.127})$ $0.151_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.290})$ $0.151_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.299})$ $0.990_{\pm 0.014}$
0.1 $0.922_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.030})$ $0.899_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $2.202_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.980})$ $2.167_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.956})$ $0.418_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.312})$ $0.419_{\pm 0.024}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.323})$ $0.977_{\pm 0.001}$
UA22.1%, RA85.7%, TA76.2% 0.15 $0.869_{\pm 0.025}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$ $0.852_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $1.467_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.490})$ $1.459_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.482})$ $0.591_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.267})$ $0.581_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.287})$ $0.958_{\pm 0.021}$
0.2 $0.814_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.028})$ $0.801_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $1.125_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.269})$ $1.138_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.275})$ $0.718_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.200})$ $0.704_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.218})$ $0.927_{\pm 0.017}$
SSD 0.05 $0.999_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.952_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $1.346_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.800})$ $1.824_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.282})$ $0.742_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.301})$ $0.522_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.072})$ $0.986_{\pm 0.001}$
0.1 $0.995_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.103})$ $0.897_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $1.033_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.189})$ $1.135_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.076})$ $0.959_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.229})$ $0.790_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.048})$ $0.847_{\pm 0.001}$
UA1.3%, RA98.4%, TA86.1% 0.15 $0.982_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.144})$ $0.847_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.987_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.956_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $0.989_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.131})$ $0.890_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.022})$ $0.517_{\pm 0.001}$
0.2 $0.959_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.173})$ $0.804_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.961_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.105})$ $0.862_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.995_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.077})$ $0.932_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.243_{\pm 0.001}$
NegGrad+ 0.05 $0.999_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.979_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$ $0.946_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.200})$ $1.443_{\pm 0.028}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.663})$ $2.248_{\pm 0.063}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.807})$ $2.358_{\pm 0.095}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.908})$ $0.992_{\pm 0.001}$
0.1 $0.996_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.104})$ $0.946_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$ $0.900_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.322})$ $1.078_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.134})$ $1.295_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.565})$ $1.332_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.590})$ $0.933_{\pm 0.003}$
UA11.5%, RA98.7%, TA83.8% 0.15 $0.990_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.152})$ $0.900_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.853_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.124})$ $1.008_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$ $1.032_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.174})$ $1.033_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.165})$ $0.712_{\pm 0.015}$
0.2 $0.977_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.191})$ $0.848_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.805_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.982_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.119})$ $0.909_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.898_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $0.381_{\pm 0.009}$
Salun 0.05 $0.993_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$ $0.962_{\pm 0.026}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.014})$ $3.284_{\pm 2.048}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.138})$ $4.112_{\pm 0.813}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.007})$ $1.546_{\pm 2.290}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.105})$ $1.558_{\pm 2.336}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.108})$ $0.989_{\pm 0.001}$
0.1 $0.976_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.084})$ $0.924_{\pm 0.039}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.026})$ $1.386_{\pm 0.423}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.164})$ $1.579_{\pm 0.130}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.368})$ $0.922_{\pm 0.566}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.192})$ $0.896_{\pm 0.607}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.154})$ $0.973_{\pm 0.002}$
UA9.2%, RA95.7%, TA81.9% 0.15 $0.944_{\pm 0.024}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.106})$ $0.876_{\pm 0.046}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $1.051_{\pm 0.175}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.074})$ $1.139_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.162})$ $0.919_{\pm 0.194}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.061})$ $0.871_{\pm 0.226}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.942_{\pm 0.002}$
0.2 $0.900_{\pm 0.044}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.114})$ $0.825_{\pm 0.049}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.910_{\pm 0.097}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.054})$ $0.969_{\pm 0.037}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.105})$ $0.928_{\pm 0.040}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.876_{\pm 0.063}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.045})$ $0.893_{\pm 0.002}$
Table 14: Unlearning performance of 8 unlearning methods on Tiny ImageNet with ViT in class-wise forgetting scenario.
Coverage Set Size CR
Methods $\alpha$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{tf}\downarrow$ $\mathcal{D}_{tr}\uparrow$ $\mathcal{D}_{f}\uparrow$ $\mathcal{D}_{tf}\uparrow$ $\mathcal{D}_{tr}\downarrow$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{tf}\downarrow$ $\mathcal{D}_{tr}\uparrow$ $\hat{q}_{f}$ $\hat{q}_{test}$
RT 0.05 $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.950_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $200.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $200.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.785_{\pm 0.056}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.532_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.000}$ $0.984_{\pm 0.002}$
0.1 $0.936_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.960_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.903_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $192.882_{\pm 0.912}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $193.340_{\pm 2.620}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.146_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.788_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.000}$ $0.859_{\pm 0.004}$
UA100%, UAtf100%, 0.15 $0.904_{\pm 0.039}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.960_{\pm 0.046}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.853_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $186.791_{\pm 2.173}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $188.880_{\pm 1.802}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.957_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.892_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.000}$ $0.535_{\pm 0.002}$
RA98.7%, TA86.4% 0.2 $0.787_{\pm 0.061}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.860_{\pm 0.024}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.805_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $171.051_{\pm 3.183}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $174.480_{\pm 2.311}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.860_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.936_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.000}$ $0.232_{\pm 0.001}$
FT 0.05 $0.993_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.960_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.040})$ $0.952_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $8.360_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}191.640})$ $8.280_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}191.720})$ $2.442_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.657})$ $0.119_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.114})$ $0.116_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.111})$ $0.390_{\pm 0.023}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.142})$ $0.999_{\pm 0.006}$ $0.993_{\pm 0.005}$
0.1 $0.984_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.048})$ $0.860_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.100})$ $0.898_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $1.802_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}191.080})$ $1.660_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}191.680})$ $1.287_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.141})$ $0.546_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.541})$ $0.518_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.513})$ $0.698_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.090})$ $0.971_{\pm 0.019}$ $0.924_{\pm 0.016}$
UA13.8%, UAtf22.0%, 0.15 $0.902_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.800_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.160})$ $0.852_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $1.120_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}185.671})$ $1.040_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}187.840})$ $1.021_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.064})$ $0.806_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.801})$ $0.769_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.764})$ $0.835_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.057})$ $0.809_{\pm 0.010}$ $0.686_{\pm 0.004}$
RA97.5%, TA84.1% 0.2 $0.860_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.073})$ $0.760_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.100})$ $0.800_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.969_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}170.082})$ $0.960_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}173.520})$ $0.882_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.022})$ $0.888_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.883})$ $0.792_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.787})$ $0.907_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.595_{\pm 0.002}$ $0.338_{\pm 0.019}$
RL 0.05 $0.998_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.980_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.020})$ $0.952_{\pm 0.049}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $199.489_{\pm 0.512}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.511})$ $195.220_{\pm 1.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.780})$ $2.317_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.532})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.411_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.121})$ $1.000_{\pm 0.000}$ $0.995_{\pm 0.032}$
0.1 $0.971_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.035})$ $0.900_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.060})$ $0.900_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $180.442_{\pm 0.710}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}12.440})$ $170.960_{\pm 0.948}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}22.380})$ $1.237_{\pm 0.050}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.991})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.727_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.061})$ $1.000_{\pm 0.000}$ $0.925_{\pm 0.024}$
UA100%, UAtf100%, 0.15 $0.922_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.018})$ $0.900_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.060})$ $0.852_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $165.884_{\pm 2.037}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}20.907})$ $159.980_{\pm 1.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}28.900})$ $1.001_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.044})$ $0.006_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.006_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.851_{\pm 0.023}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.041})$ $1.000_{\pm 0.000}$ $0.641_{\pm 0.035}$
RA98.2%, TA84.6% 0.2 $0.882_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.095})$ $0.860_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.807_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $154.896_{\pm 2.028}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}16.155})$ $149.280_{\pm 3.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}25.200})$ $0.886_{\pm 0.032}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.026})$ $0.006_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.006_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.912_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $1.000_{\pm 0.000}$ $0.262_{\pm 0.022}$
GA 0.05 $1.000_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.980_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.020})$ $0.948_{\pm 0.026}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $22.836_{\pm 0.045}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}177.164})$ $20.600_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}179.400})$ $1.781_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $0.044_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $0.048_{\pm 0.028}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.043})$ $0.532_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.000}$ $0.984_{\pm 0.033}$
0.1 $0.991_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.900_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.060})$ $0.897_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $1.631_{\pm 0.031}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}191.251})$ $1.720_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}191.620})$ $1.133_{\pm 0.044}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.013})$ $0.608_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.603})$ $0.523_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.518})$ $0.792_{\pm 0.037}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $0.972_{\pm 0.033}$ $0.849_{\pm 0.039}$
UA9.1%, UAtf20.0%, 0.15 $0.958_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.054})$ $0.820_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.140})$ $0.850_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $1.151_{\pm 0.039}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}185.640})$ $1.140_{\pm 0.042}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}187.740})$ $0.958_{\pm 0.026}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.832_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.827})$ $0.719_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.714})$ $0.887_{\pm 0.044}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.868_{\pm 0.023}$ $0.535_{\pm 0.011}$
RA98.6%, TA86.1% 0.2 $0.880_{\pm 0.047}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.093})$ $0.800_{\pm 0.051}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.060})$ $0.803_{\pm 0.025}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.929_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}170.122})$ $0.900_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}173.580})$ $0.861_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.947_{\pm 0.036}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.942})$ $0.889_{\pm 0.029}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.884})$ $0.933_{\pm 0.027}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.473_{\pm 0.016}$ $0.238_{\pm 0.000}$
Teacher 0.05 $0.982_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.018})$ $1.000_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.952_{\pm 0.025}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $199.971_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $200.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $5.095_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.310})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.187_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.345})$ $1.000_{\pm 0.000}$ $0.989_{\pm 0.001}$
0.1 $0.909_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$ $0.940_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.020})$ $0.903_{\pm 0.032}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $199.813_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.931})$ $199.900_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.560})$ $2.033_{\pm 0.031}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.887})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.444_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.344})$ $1.000_{\pm 0.000}$ $0.965_{\pm 0.003}$
UA100%, UAtf100%, 0.15 $0.887_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.017})$ $0.880_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.080})$ $0.854_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $199.667_{\pm 0.030}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}12.876})$ $199.760_{\pm 0.026}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}10.880})$ $1.331_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.374})$ $0.004_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.004_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.641_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.251})$ $1.000_{\pm 0.000}$ $0.919_{\pm 0.001}$
RA88.8%, TA78.6% 0.2 $0.838_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.051})$ $0.840_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.020})$ $0.799_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $199.413_{\pm 0.024}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}28.362})$ $199.620_{\pm 0.030}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}25.140})$ $1.022_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.162})$ $0.004_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.004_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.781_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.155})$ $1.000_{\pm 0.000}$ $0.825_{\pm 0.002}$
SSD 0.05 $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.950_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $198.769_{\pm 0.052}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.231})$ $197.320_{\pm 1.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.680})$ $1.866_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.081})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.509_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.023})$ $1.000_{\pm 0.000}$ $0.986_{\pm 0.006}$
0.1 $0.949_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.013})$ $0.900_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.060})$ $0.897_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $171.073_{\pm 0.209}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}21.809})$ $169.360_{\pm 2.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}23.980})$ $1.141_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.006_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.786_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $1.000_{\pm 0.000}$ $0.854_{\pm 0.006}$
UA100%, UAtf100%, 0.15 $0.913_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.880_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.080})$ $0.852_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $157.140_{\pm 1.209}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}29.651})$ $154.960_{\pm 0.907}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}33.920})$ $0.959_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.006_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.006_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.888_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $1.000_{\pm 0.000}$ $0.538_{\pm 0.007}$
RA98.4%, TA86.1% 0.2 $0.833_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.046})$ $0.800_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.060})$ $0.806_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $136.502_{\pm 3.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}34.549})$ $136.420_{\pm 2.422}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}38.060})$ $0.864_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $0.006_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.006_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.932_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $1.000_{\pm 0.000}$ $0.254_{\pm 0.005}$
NegGrad+ 0.05 $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.947_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $200.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $200.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.850_{\pm 0.036}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.065})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.512_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.020})$ $1.000_{\pm 0.000}$ $0.987_{\pm 0.001}$
0.1 $0.927_{\pm 0.104}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.950_{\pm 0.071}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.894_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $193.994_{\pm 8.493}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.112})$ $197.490_{\pm 3.550}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.150})$ $1.140_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.784_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $1.000_{\pm 0.000}$ $0.859_{\pm 0.003}$
UA100%, UAtf100%, 0.15 $0.862_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.042})$ $0.870_{\pm 0.042}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.090})$ $0.849_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $188.686_{\pm 0.954}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.894})$ $195.590_{\pm 0.863}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.710})$ $0.961_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.004_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.884_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $1.000_{\pm 0.000}$ $0.537_{\pm 0.003}$
RA99.0%, TA85.8% 0.2 $0.830_{\pm 0.027}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.043})$ $0.840_{\pm 0.085}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.020})$ $0.802_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $187.219_{\pm 0.064}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}16.168})$ $194.310_{\pm 0.948}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}19.830})$ $0.861_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.004_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.004_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.931_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $1.000_{\pm 0.000}$ $0.220_{\pm 0.002}$
Salun 0.05 $0.997_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.993_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.949_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $199.599_{\pm 0.207}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.401})$ $197.440_{\pm 1.244}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.560})$ $1.980_{\pm 0.050}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.196})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.479_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $1.000_{\pm 0.000}$ $0.989_{\pm 0.001}$
0.1 $0.975_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.039})$ $0.927_{\pm 0.023}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.033})$ $0.899_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $191.973_{\pm 1.616}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.910})$ $185.220_{\pm 0.918}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}8.120})$ $1.169_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.023})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.769_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $1.000_{\pm 0.000}$ $0.884_{\pm 0.001}$
UA100%, UAtf100%, 0.15 $0.961_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.057})$ $0.860_{\pm 0.040}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.100})$ $0.850_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $187.825_{\pm 3.461}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.034})$ $180.307_{\pm 2.908}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}8.573})$ $0.969_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.877_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $1.000_{\pm 0.000}$ $0.562_{\pm 0.003}$
RA98.4%, TA86.1% 0.2 $0.960_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.173})$ $0.840_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.020})$ $0.801_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $184.838_{\pm 3.478}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}13.787})$ $177.647_{\pm 2.627}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.167})$ $0.863_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.928_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $1.000_{\pm 0.000}$ $0.230_{\pm 0.009}$
Table 15: MIACR performance on CIFAR-10 with ResNet-18.
10% Forgetting 50% Forgetting
Methods $\alpha$ MIACR $\uparrow$ $\hat{q}$ MIACR $\uparrow$ $\hat{q}$
RT 0.05 $0.091_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.877_{\pm 0.004}$ $0.117_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.899_{\pm 0.007}$
0.1 $0.147_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.589_{\pm 0.008}$ $0.201_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.570_{\pm 0.001}$
MIA86.92% ($10\%$ Forgetting) 0.15 $0.203_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.485_{\pm 0.005}$ $0.272_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.472_{\pm 0.009}$
MIA82.79% ($50\%$ Forgetting) 0.2 $0.246_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.473_{\pm 0.001}$ $0.318_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.459_{\pm 0.003}$
FT 0.05 $0.039_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.745_{\pm 0.013}$ $0.036_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.081})$ $0.780_{\pm 0.011}$
0.1 $0.077_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.070})$ $0.627_{\pm 0.000}$ $0.103_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.098})$ $0.558_{\pm 0.012}$
MIA92.00% ($10\%$ Forgetting) 0.15 $0.128_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.075})$ $0.517_{\pm 0.008}$ $0.159_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.113})$ $0.494_{\pm 0.011}$
MIA92.92% ($50\%$ Forgetting) 0.2 $0.196_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.050})$ $0.483_{\pm 0.003}$ $0.244_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.074})$ $0.476_{\pm 0.004}$
RL 0.05 $0.083_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.627_{\pm 0.011}$ $0.050_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.067})$ $0.547_{\pm 0.000}$
0.1 $0.178_{\pm 0.027}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$ $0.572_{\pm 0.005}$ $0.137_{\pm 0.030}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.064})$ $0.547_{\pm 0.001}$
MIA74.21% ($10\%$ Forgetting) 0.15 $0.272_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.069})$ $0.492_{\pm 0.015}$ $0.194_{\pm 0.031}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.078})$ $0.547_{\pm 0.001}$
MIA61.15% ($50\%$ Forgetting) 0.2 $0.320_{\pm 0.025}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.074})$ $0.485_{\pm 0.011}$ $0.261_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.057})$ $0.546_{\pm 0.000}$
GA 0.05 $0.012_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.079})$ $0.862_{\pm 0.016}$ $0.012_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.105})$ $0.771_{\pm 0.008}$
0.1 $0.032_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.115})$ $0.502_{\pm 0.016}$ $0.055_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.146})$ $0.486_{\pm 0.005}$
MIA98.80% ($10\%$ Forgetting) 0.15 $0.076_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.127})$ $0.477_{\pm 0.007}$ $0.107_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.165})$ $0.474_{\pm 0.015}$
MIA98.86% ($50\%$ Forgetting) 0.2 $0.146_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.100})$ $0.476_{\pm 0.019}$ $0.164_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.154})$ $0.473_{\pm 0.011}$
Teacher 0.05 $0.013_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.078})$ $0.750_{\pm 0.014}$ $0.031_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$ $0.635_{\pm 0.018}$
0.1 $0.038_{\pm 0.023}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.109})$ $0.672_{\pm 0.028}$ $0.065_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.136})$ $0.582_{\pm 0.013}$
MIA87.24% ($10\%$ Forgetting) 0.15 $0.072_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.131})$ $0.625_{\pm 0.029}$ $0.110_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.162})$ $0.548_{\pm 0.007}$
MIA93.24% ($50\%$ Forgetting) 0.2 $0.113_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.133})$ $0.588_{\pm 0.019}$ $0.159_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.159})$ $0.532_{\pm 0.006}$
FF 0.05 $0.038_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.500_{\pm 0.003}$ $0.043_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.074})$ $0.508_{\pm 0.003}$
0.1 $0.051_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.096})$ $0.486_{\pm 0.018}$ $0.089_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.112})$ $0.509_{\pm 0.013}$
MIA71.52% ($10\%$ Forgetting) 0.15 $0.080_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.123})$ $0.474_{\pm 0.013}$ $0.130_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.142})$ $0.506_{\pm 0.007}$
MIA76.02% ($50\%$ Forgetting) 0.2 $0.109_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.137})$ $0.473_{\pm 0.002}$ $0.168_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.150})$ $0.499_{\pm 0.006}$
SSD 0.05 $0.011_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.080})$ $0.861_{\pm 0.012}$ $0.012_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.105})$ $0.748_{\pm 0.011}$
0.1 $0.031_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.116})$ $0.511_{\pm 0.011}$ $0.051_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.150})$ $0.488_{\pm 0.001}$
MIA98.78% ($10\%$ Forgetting) 0.15 $0.077_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.126})$ $0.480_{\pm 0.013}$ $0.104_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.168})$ $0.477_{\pm 0.015}$
MIA98.87% ($50\%$ Forgetting) 0.2 $0.139_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.107})$ $0.475_{\pm 0.013}$ $0.168_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.150})$ $0.477_{\pm 0.006}$
NegGrad+ 0.05 $0.076_{\pm 0.025}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.844_{\pm 0.024}$ $0.045_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.072})$ $0.863_{\pm 0.025}$
0.1 $0.128_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $0.481_{\pm 0.009}$ $0.109_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.092})$ $0.511_{\pm 0.008}$
MIA90.30% ($10\%$ Forgetting) 0.15 $0.174_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.480_{\pm 0.005}$ $0.167_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.105})$ $0.477_{\pm 0.010}$
MIA93.82% ($50\%$ Forgetting) 0.2 $0.213_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.033})$ $0.480_{\pm 0.004}$ $0.230_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.088})$ $0.472_{\pm 0.008}$
Salun 0.05 $0.055_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.691_{\pm 0.011}$ $0.044_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.073})$ $0.670_{\pm 0.008}$
0.1 $0.113_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.034})$ $0.681_{\pm 0.013}$ $0.115_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$ $0.630_{\pm 0.009}$
MIA57.58% ($10\%$ Forgetting) 0.15 $0.198_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.642_{\pm 0.015}$ $0.170_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.102})$ $0.610_{\pm 0.003}$
MIA59.12% ($50\%$ Forgetting) 0.2 $0.267_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $0.608_{\pm 0.011}$ $0.220_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.098})$ $0.586_{\pm 0.005}$
Table 16: Performance of our unlearning framework. We show the unlearning performance on CIFAR-10 with ResNet-18 and Tiny ImageNet with ViT in $10\%$ random data forgetting scenario.
$\lambda=0.2$ $\lambda=0.5$ $\lambda=1$
[…]
CIFAR-10 with ResNet-18
$0.05$ $0.788({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.076})$ $0.824({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.763({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.101})$ $0.825({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.054})$ $0.719({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.145})$ $0.82({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.059})$
RT $0.1$ $10.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.2})$ $98.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.4})$ $91.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.8})$ $0.914({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.924({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $14.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.4})$ $97.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.9})$ $90.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.4})$ $0.879({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.064})$ $0.912({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.033})$ $17.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}9.1})$ $96.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.9})$ $90.5\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.3})$ $0.838({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.105})$ $0.911({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.034})$
$0.15$ $0.956({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $0.959({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.936({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.039})$ $0.954({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.014})$ $0.906({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.069})$ $0.951({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.017})$
$0.2$ $0.977({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.976({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.963({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.025})$ $0.966({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.932({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.056})$ $0.965({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.016})$
$0.05$ $0.844({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.020})$ $0.829({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.050})$ $0.853({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.843({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.835({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.854({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.025})$
FT $0.1$ $6.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.8})$ $97.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.7})$ $90.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.0})$ $0.948({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.924({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $7.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.7})$ $96.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.8})$ $90.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $0.940({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.927({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.018})$ $9.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.6})$ $97.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.8})$ $91.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.6})$ $0.938({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.936({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$
$0.15$ $0.983({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.959({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.975({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.961({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.976({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.970({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$
$0.2$ $0.989({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.974({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.983({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.975({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.986({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.984({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$
$0.05$ $0.709({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.155})$ $0.736({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.143})$ $0.708({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.156})$ $0.731({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.148})$ $0.629({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.235})$ $0.669({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.210})$
RL $0.1$ $9.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.1})$ $96.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.1})$ $89.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.896({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$ $0.887({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.058})$ $9.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.3})$ $96.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.8})$ $89.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.1})$ $0.902({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.041})$ $0.896({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.049})$ $12.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.0})$ $95.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.4})$ $88.1\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.7})$ $0.845({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.098})$ $0.858({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.087})$
$0.15$ $0.946({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.931({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.037})$ $0.939({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.932({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.911({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.064})$ $0.913({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$
$0.2$ $0.964({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $0.949({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.959({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.950({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$ $0.936({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.938({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.043})$
Tiny ImageNet with ViT
$0.05$ $0.458({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.486})$ $0.516({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.433})$ $0.396({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.548})$ $0.489({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.460})$ $0.346({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.598})$ $0.481({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.468})$
RT $0.1$ $19.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.6})$ $98.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $86.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $0.729({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.163})$ $0.786({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.114})$ $26.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}11.7})$ $98.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.1})$ $85.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.2})$ $0.649({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.243})$ $0.765({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.135})$ $35.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}21.0})$ $98.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.2})$ $85.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.8})$ $0.549({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.343})$ $0.739({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.161})$
$0.15$ $0.841({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.889({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.039})$ $0.768({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.073})$ $0.88({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.030})$ $0.658({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.183})$ $0.861({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$
$0.2$ $0.898({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.108})$ $0.932({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.133})$ $0.839({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.049})$ $0.929({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.130})$ $0.743({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$ $0.918({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.119})$
$0.05$ $0.441({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.503})$ $0.399({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.550})$ $0.413({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.531})$ $0.401({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.548})$ $0.342({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.602})$ $0.363({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.586})$
FT $0.1$ $9.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.9})$ $97.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.4})$ $83.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.753({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.139})$ $0.683({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.217})$ $13.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $97.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.6})$ $83.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.718({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.174})$ $0.683({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.217})$ $20.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.3})$ $96.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $82.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.1})$ $0.627({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.265})$ $0.652({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.248})$
$0.15$ $0.884({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.043})$ $0.823({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$ $0.848({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.819({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$ $0.772({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.069})$ $0.802({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.048})$
$0.2$ $0.942({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.152})$ $0.893({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.094})$ $0.914({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.124})$ $0.890({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.091})$ $0.856({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.066})$ $0.877({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.078})$
$0.05$ $0.051({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.893})$ $0.111({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.838})$ $0.051({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.893})$ $0.121({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.828})$ $0.048({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.896})$ $0.119({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.830})$
RL $0.1$ $31.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}17.1})$ $95.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}17.9})$ $80.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.1})$ $0.278({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.614})$ $0.451({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.449})$ $36.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}21.5})$ $95.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.5})$ $80.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.6})$ $0.254({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.638})$ $0.449({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.451})$ $40.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}25.5})$ $94.5\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.3})$ $79.5\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.5})$ $0.236({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.656})$ $0.436({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.464})$
$0.15$ $0.579({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.262})$ $0.710({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.140})$ $0.541({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.300})$ $0.708({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.142})$ $0.480({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.361})$ $0.673({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.177})$
$0.2$ $0.752({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.038})$ $0.825({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.026})$ $0.718({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.072})$ $0.827({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.028})$ $0.642({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.148})$ $0.793({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$

### v2

@@
 Abstract
-Machine unlearning seeks to systematically remove specified data from a trained model, effectively achieving a state as though the data had never been encountered during training.
-While metrics such as Unlearning Accuracy (UA) and Membership Inference Attack (MIA) provide a baseline for assessing unlearning performance, they fall short of evaluating the completeness and reliability of forgetting.
-This is because the ground truth labels remain potential candidates within the scope of uncertainty quantification, leaving gaps in the evaluation of true forgetting.
-In this paper, we identify critical limitations in existing unlearning metrics and propose enhanced evaluation metrics inspired by conformal prediction.
-Our metrics can effectively capture the extent to which ground truth labels are excluded from the prediction set.
-Furthermore, we observe that many existing machine unlearning methods do not achieve satisfactory forgetting performance when evaluated with our new metrics.
-To address this, we propose an unlearning framework that integrates conformal prediction insights into Carlini & Wagner adversarial attack loss.
-Extensive experiments on the image classification task demonstrate that our enhanced metrics offer deeper insights into unlearning effectiveness, and that our unlearning framework significantly improves the forgetting quality of unlearning methods.
+Machine unlearning seeks to remove the influence of specified data from a trained model.
+While metrics such as unlearning accuracy (UA) and membership inference attack (MIA) provide baselines for assessing unlearning performance, they fall short of evaluating the forgetting reliability.
+In this paper, we find that the data misclassified across UA and MIA still have their ground truth labels included in the prediction set from the uncertainty quantification perspective, which raises a fake unlearning issue.
+To address this issue, we propose two novel metrics inspired by conformal prediction that more reliably evaluate forgetting quality.
+Building on these insights, we further propose a conformal prediction-based unlearning framework that integrates conformal prediction into Carlini & Wagner adversarial attack loss, which can significantly push the ground truth label out of the conformal prediction set.
+Through extensive experiments on image classification task, we demonstrate both the effectiveness of our proposed metrics and the superiority of our unlearning framework, which improves the UA of existing unlearning methods by an average of $6.6\%$ through the incorporation of a tailored loss term alone.
 1 Introduction
-Machine unlearning has become crucial in advancing data privacy, especially under legal requirements like the General Data Protection Regulation (GDPR) (Bourtoule et al., 2021).
-These regulations emphasize the right for individuals to have their data removed or forgotten, creating a demand for machine unlearning methods that can enable machine learning models to behave as if specific forget data were never used in training stage.
-The existing post hoc machine unlearning methods can be categorized into training-based (Graves et al., 2021; Warnecke et al., 2021; Thudi et al., 2022; Tarun et al., 2023) and training-free (Guo et al., 2019; Sekhari et al., 2021; Nguyen et al., 2020; Golatkar et al., 2020b, 2021; Foster et al., 2024) approaches, depending on whether they require any traditional model training steps during the unlearning process (Foster et al., 2024).
-Table 1: Grad-CAM maps of one original model in CIFAR-10 with ResNet18 and two corresponding unlearning models.
-The Prediction row indicates whether the model correctly predicts the image’s true label, while the In Set row represents whether the true label is included in the prediction set.
-Although the Finetune unlearning method, can misclassify the forget data, Grad-CAM can still highlight key features of the object under this model since the true label is included in the prediction set.
-In contrast, our unlearning method removes the true label from the set, with activation regions shifting significantly away from the object’s key features.
-This confirms that the forgetting quality is better if the true label can be excluded from the prediction set.
-Class Name Forget Data Original Model Finetune Method Our Method
-Wok [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Swimming Trunks [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Prediction
-In Set
-To measure the forgetting quality and predictive performance of the model after the unlearning process (i.e., unlearning model), several unlearning metrics have been proposed (Kashef, 2021; Shokri et al., 2017; Chen et al., 2021; Brophy & Lowd, 2021; Cao & Yang, 2015).
-However, existing unlearning metrics, such as unlearning accuracy (UA) (Brophy & Lowd, 2021; Foster et al., 2024) and membership inference attack (MIA) (Shokri et al., 2017; Chen et al., 2021), fall short in fully evaluating forgetting quality and reliability.
-These metrics primarily focus on whether models can predict forget data accurately without sufficiently addressing how well the model forgets.
+Machine unlearning has become crucial in advancing data privacy, especially under legal requirements like the General Data Protection Regulation (GDPR) Bourtoule et al.
+(2021).
+These regulations emphasize the right for individuals to have their data removed, creating a demand for unlearning methods that can enable models to behave as if specific data were never used in training stage.
+Beyond privacy considerations, machine unlearning has also emerged as a promising tool to remove harmful biases and stereotypes embodied in the model.
+The existing post hoc machine unlearning methods can be categorized into training-based Graves et al.
+(2021); Tarun et al.
+(2023); Thudi et al.
+(2022); Warnecke et al.
+(2021) and training-free Foster et al.
+(2024); Golatkar et al.
+(2021, 2020); Guo et al.
+(2019); Nguyen et al.
+(2020); Sekhari et al.
+(2021) approaches, depending on whether they require any model training steps during the unlearning process Foster et al.
+(2024).
+To measure the forgetting quality and predictive performance of an unlearning model, several unlearning metrics have been proposed Brophy and Lowd (2021); Cao and Yang (2015); Chen et al.
+(2021); Kashef (2021); Shokri et al.
+(2017).
+However, existing unlearning metrics, such as unlearning accuracy (UA) and membership inference attack (MIA), fall short in fully evaluating forgetting reliability.
+These metrics primarily focus on whether models can predict forget data accurately without sufficiently addressing how well the model forgets the data.
 In a nutshell, misclassifying forget data does not mean that the model has completely forgotten it to some extent.
-To verify this view, we apply conformal prediction (Papadopoulos et al., 2002; Lei & Wasserman, 2014) to the unlearning models.
-Through experiments, we find that although the model misclassifies some forget data, a significant portion of these misclassified instances still appear in the conformal prediction set.
-Pictures in Table 1, which visualize the important feature of models’ prediction of models by using Grad-CAM (Selvaraju et al., 2017), further explain this phenomenon.
-Despite the Finetune method incorrectly predicting the forget data, the Grad-CAM maps still focus heavily on the important features of the object itself.
-Based on the above findings, we design two novel metrics that capture the uncertainty and robustness of unlearning performance more effectively inspired by conformal prediction.
-Additionally, motivated by Carlini & Wagner (C&W) attack (Carlini & Wagner, 2017) and conformal prediction, we propose a general unlearning framework to improve training-based unlearning methods and promote reliable forgetting.
-Grad-CAM maps of our method in Table 1 reveal that once the true label no longer falls within the conformal prediction set, the activation regions shift significantly.
-To sum up, the contributions of our work are as follows:
-  • We identify pivotal limitations in current unlearning metrics, as they overlook misclassified data where ground truth labels remain potential candidates under uncertainty quantification.
+To verify this view, conformal prediction Lei and Wasserman (2014); Papadopoulos et al.
+(2002) as an uncertainty quantification technique, is applied in our work to recover the misclassified data in UA and MIA.
+Through extensive experiments,
+Figure 1: Grad-CAM maps of one original model and two corresponding unlearning models in CIFAR-10 with ResNet18.
+The Classification row indicates whether the model correctly predicts the image’s true label, while the In Set row represents whether the true label is included in the prediction set.
+Although the Finetune method, can misclassify the forget data, Grad-CAM can still highlight key features of the object since the true label is included in the prediction set.
+In contrast, when our unlearning method removes the true label from the set, activation regions shift significantly away from the object’s key features.
+This confirms that the forgetting quality improves if the true label can be excluded from the prediction set.
+we find that although the model misclassifies some forget data from the UA and MIA perspective, over $50\%$ of these misclassified data instances still appear in the conformal prediction set and can be easily recovered, which exposes a fake unlearning issue.
+As shown in Figure 1, the important features of models’ prediction by Grad-CAM Selvaraju et al.
+(2017) visualize this fake unlearning issue.
+Despite the Finetune method misclassifying the forget data, the Grad-CAM maps still focus heavily on the important features of the object itself.
+Based on the above insights, we design two novel metrics CR and MIACR that more effectively capture the uncertainty and robustness of unlearning performance inspired by conformal prediction to tackle the fake unlearning issue.
+Additionally, motivated by conformal prediction insights about fake unlearning and Carlini & Wagner (C&W) attack loss Carlini and Wagner (2017), we propose a general unlearning framework, which can improve existing training-based unlearning methods and promote reliable forgetting.
+Grad-CAM maps of our method in Figure 1 reveal that once the true label no longer falls within the conformal prediction set, the activation regions shift significantly.
+To sum up, our contributions are as follows:
+  • Our analysis reveals that conformal prediction can recover a substantial portion of data previously classified as forgotten by existing unlearning metrics.
+This fake unlearning issue underscores critical limitations in existing unlearning evaluation methodologies.
   • We design two novel metrics to address the limitations motivated by conformal prediction.
-  • We propose a general unlearning framework for training-based machine unlearning methods motivated by conformal prediction and C&W loss.
-  • Extensive experiments demonstrate the effectiveness of novel metrics and our unlearning framework.
+  • We propose an unlearning framework motivated by conformal prediction and C&W loss, enhancing existing training-based unlearning methods over both existing and our metrics.
 2 Enhancing Metrics for Machine Unlearning Based on Conformal Prediction
@@
 Machine Unlearning.
-Machine unlearning is the targeted removal of certain training data effects from a machine learning model.
-In our work, two different forgetting scenarios are considered: (i) random data forgetting focuses on randomly forgetting specific instances within the training data, and (ii) class-wise forgetting aims to remove all information associated with an entire class.
+In our work, we focus on the image classification task, which is widely used in prior unlearning literature Foster et al.
+(2024); Graves et al.
+(2021); Shen et al.
+(2024).
+Two forgetting scenarios are considered in this work: (i) random data forgetting focuses on randomly forgetting specific data instances within the training data, and (ii) class-wise forgetting aims to remove all data information associated with an entire class.
 Let $\mathcal{D}_{train}$ denote the original training data used to obtain an original model ${\bm{\theta}}_{o}$.
 We split the whole training data $\mathcal{D}_{train}$ into two subsets, forget data $\mathcal{D}_{f}$ and retain data $\mathcal{D}_{r}=\mathcal{D}_{train}\setminus\mathcal{D}_{f}$.
-In random data forgetting, $\mathcal{D}_{test}$ represents test data.
-In class-wise forgetting, $\mathcal{D}_{tf}$ corresponds to the test-forget data, exclusively containing the forget class, while $\mathcal{D}_{tr}$ represents the test-retain data within the test data $\mathcal{D}_{test}$.
-Let $\theta_{u}$ denote the model after the unlearning process, where the influence of forget data $\mathcal{D}_{f}$ has been removed.
-The purpose of machine unlearning is to enable $\theta_{u}$ to forget $\mathcal{D}_{f}$ (thus no longer recognizing $\mathcal{D}_{tf}$ in the class-wise forgetting scenario), while maintaining accuracy on the retain data $\mathcal{D}_{r}$ and avoiding a significant drop in accuracy on $\mathcal{D}_{test}$ in the random data forgetting scenario or $\mathcal{D}_{tr}$ in the class-wise setting.
+Let $\mathcal{D}_{test}$ represent test data.
+$\theta_{u}$ denotes the model after the unlearning process, where the information of forget data $\mathcal{D}_{f}$ has been removed.
+The purpose of machine unlearning is to enable $\theta_{u}$ to forget $\mathcal{D}_{f}$, while maintaining accuracy on $\mathcal{D}_{r}$ and $\mathcal{D}_{test}$.
 Conformal Prediction.
-Conformal prediction is proposed to quantify uncertainty in machine learning models, providing prediction sets that contain the true label with a guaranteed probability (Angelopoulos & Bates, 2021).
-Among the various types of conformal prediction, this work specifically focuses on split conformal prediction (SCP) [^1] Note that while the goal is to remove the influence of the forget data so that it behaves similarly to the calibration data, the exchangeability property may not always hold in machine unlearning settings.
-Here, we are directly leveraging the concept of conformal prediction to evaluate machine unlearning performance..
-To construct the conformal prediction set, it involves four steps:
+Conformal prediction is proposed to quantify uncertainty in machine learning models, providing prediction sets that contain the ground truth label with a theoretically guaranteed probability Angelopoulos and Bates (2021).
+Among the various types of conformal prediction, this work specifically focuses on split conformal prediction (SCP).
+To construct the conformal prediction set, it involves four steps on the unlearning model:
   1. 1.
-    Calibration Data. SCP first chooses unseen data as calibration data. The number of calibration data points should be enough to evaluate the model’s uncertainty.
+    Calibration Data. SCP first chooses unseen data as calibration data, which must be held out from both the training and test sets to ensure independence.
   2. 2.
@@
     $S({\bm{x}},y_{i})=1-p_{i}({\bm{x}}),$ (1)
-    where $p_{i}({\bm{x}})$ represents the probability output (through a softmax function) for the class $y_{i}$.
+    where $p_{i}({\bm{x}})$ represents the probability of different class $y_{i}$.
   3. 3.
     Quantile Computation.
-Given a target miscoverage rate $\alpha\in[0,1]$, SCP obtains threshold $\hat{q}$ by taking the $1-\alpha$ quantile of the non-conformity score on the calibration data $\mathcal{D}_{c}$.
+Given a target miscoverage rate $\alpha\in[0,1]$, SCP obtains threshold $\hat{q}$ by taking the $1-\alpha$ quantile of the non-conformity score of the ground truth labels $y_{t}$ on the calibration data $({\bm{x}},y_{t})\in\mathcal{D}_{c}$,
+    $\hat{q}=\text{Quantile}_{1-\alpha}\big{(}S({\bm{x}},y_{t})\big{)}.$ (2)
   4. 4.
@@
     ${\mathbb{C}}({\bm{x}})=\{y_{i}:S({\bm{x}},y_{i})\leq\hat{q}\},$ (3)
-    where $y_{i}$ represents different classes.
-2.2 Identifying Limitations in Current Unlearning Metrics
-Evaluating the effectiveness of machine unlearning involves a trade-off between how well the model forgets unwanted data (forgetting quality) and how much it retains useful knowledge (predictive performance).
-The first key question we pose is as follows:
-(Q1) Are existing commonly used metrics sufficient to evaluate unlearning performance?
+2.2 Identifying Fake Unlearning in Existing Unlearning Metrics
+Evaluating the effectiveness of machine unlearning involves a trade-off between how well the model forgets data and retains useful knowledge. The first key question we pose is as follows:
+(Q1) Can we recover the data that is identified as forgotten by the metrics UA and MIA?
 […]
@@
 RL 7.55 97.41 90.60 74.21 10.48 93.91 85.78 61.15
+We propose a conformal prediction-based recovery method.
+If the ground truth of forget data falls within the conformal prediction set, we consider the recovery successful.
+Thus, fake unlearning is defined as the scenario where the data identified as forgotten by existing metrics can be recovered by conformal prediction.
+To substantiate our claim, we first apply metrics, UA (i.e., $1-$ the accuracy on forget data), retain accuracy (RA, i.e., accuracy on retain data), test accuracy (TA, i.e., accuracy on test data) and MIA, to evaluate 3 classic unlearning methods, Retrain (RT), Finetune (FT) Warnecke et al.
+(2021), and Random Label (RL) Graves et al.
+(2021).
+The results are trained on CIFAR-10 with ResNet-18 in a random data forgetting scenario.
+In Table 1, the UA and MIA results prove that the models indeed fail to correctly classify some forget data and identify membership.
+However, higher UA and lower MIA cannot ensure that these forget data points can not appear in any form within the model’s predictions.
 Table 2: Mis-label (mis-classification) count and in-set ratio of UA and MIA metrics for RT, FT and RL on CIFAT-10 with ResNet-18 under $10\%$ and $50\%$ random data forgetting scenarios.
@@
 RL 1,289 1,011 78.4% 9,713 8,295 85.4%
-We apply the most commonly used metrics, unlearning accuracy (UA), retain accuracy (RA), test accuracy (TA) and MIA, to evaluate three classic machine unlearning methods: Retrain (RT), Finetune (FT) (Warnecke et al., 2021), and Random Label (RL) (Graves et al., 2021).
-These methods are implemented on a ResNet-18 model trained with the CIFAR-10 dataset in a random data forgetting scenario.
-In Table 2, compared with RA on $\mathcal{D}_{r}$, the UA of the three methods proves that the model indeed forgets some forget data.
-The MIA column also indicates that many data points seem to be forgotten.
-These metrics appear to provide valuable insights into the extent of forgetting achieved.
-However, higher unlearning accuracy cannot ensure that these forget data points are no longer represented in any form within the model’s predictions.
-To validate our hypothesis, we employ the process of conformal prediction to examine whether the true labels of the forget data appear within the prediction set.
-We set the confidence level to $95\%$, i.e., $\alpha=0.05$, and set the calibration set size to $2,000$.
-In Table 3, we count both UA metric and MIA metric’s misclassified data points (mis-label) and, after applying conformal prediction, determined how many of these points fall within the conformal prediction set (in-set).
-Table 3 demonstrates that even though the model misclassifies part of the forget data, on average 54.6% of these misclassified instances are still present in the conformal prediction set.
-This indicates that a high UA does not mean the model has truly forgotten the data.
-Relying solely on UA to evaluate the forgetting quality is insufficient.
-In MIA, the prediction of ‘0’ indicates a data point is predicted as “truly forget”, while ‘1’ signifies “still retain”.
-And “mis-label” refers to the number of the model predicts the forget data points as ‘0’.
-In conformal prediction of MIA, there are three potential prediction set results, i.e., $\{0\},\{1\},\{0,1\}$.
-The “in-set” here refers to the number of mislabeled data points whose conformal prediction set includes $1$.
-Thus, the in-set ratio indicates that, although the MIA fails to identify an average of 18.33% of the forget data points, conformal prediction can still recognize more than half of these forget data points within its prediction set.
-Overall, high in-set ratios in conformal prediction highlight that a considerable portion of the forget data remains identifiable through conformal prediction.
-This implies these unlearning methods may not be fully effective in eliminating traces of the forget data from another perspective.
+We employ conformal prediction to investigate whether we can recover forget data’s ground truth, specifically, whether the ground truth labels still appear within the conformal prediction sets.
+We set the confidence level to $95\%$, i.e., $\alpha=0.05$, and set the calibration set size to $2000$.
+In Table 2, we count the number of data points that are identified as truly forgotten by UA and MIA (marked as mis-label) and count how many of these points can still be recovered (marked as in-set).
+The results of UA reveal that even though the model misclassifies part of the forget data, on average $54.6\%$ of these misclassified data instances are still recovered by conformal prediction.
+This finding demonstrates that a high UA does not mean the model has truly forgotten the data, and thus relying solely on UA to evaluate the forgetting quality is fragile.
+A similar phenomenon occurs on results of MIA.
+In MIA, the prediction of ‘0’ indicates a data point is truly forgotten, while ‘1’ signifies it is still identified as the training membership.
+The mis-label column of MIA refers to the number of data points that are predicted as ‘0’.
+The in-set here refers to the number of mis-label data points whose conformal prediction set still includes ‘1’.
+Thus, the recover ratio indicates that, although the MIA fails to identify an average of $18.33\%$ of the forget data as training membership, conformal prediction can still recover $54.7\%$ of these forget data within prediction sets.
+For more results of other unlearning methods, see Table 6 in AppendixB.1.
+Overall, the high recover ratio observed across both UA and MIA in Tables 2 and 6 highlights that these unlearning methods are not fully effective at eliminating traces of the forget data from the uncertainty quantification perspective and can still be recovered by conformal prediction.
+This encloses that the fake unlearning issue arises when the ground truth label of misclassified data falls within the conformal prediction set.
 2.3 Designing Metrics Motivated by Conformal Prediction
-Based on the limitations of UA and MIA metrics shown in Section 2.2, it raises a question as follows:
-(Q2) Can we develop metrics to address the shortcomings of UA and MIA?
+Based on the limitation of UA and MIA metrics shown in Section 2.2, it raises a question as follows:
+(Q2) Can we develop metrics to address the fake unlearning issue of UA and MIA?
 Thus, we propose enhanced UA and MIA metrics that draw intuition from conformal prediction.
@@
 Conformal Ratio (CR).
-To overcome the limitations of UA, we introduce a novel metric, Conformal Ratio (CR), which incorporates both coverage and set size in conformal prediction to provide a more comprehensive evaluation.
+To overcome the fake unlearning inherent in UA, we introduce a novel metric, CR, which incorporates both coverage and set size in conformal prediction to provide a more comprehensive evaluation.
 Before defining CR, we introduce Coverage and Set Size.
-The definition of Coverage, given a dataset $\mathcal{D}$, is as follows:
+Given a dataset $\mathcal{D}$, the definition of Coverage is as follows:
 $\text{Coverage}:=\frac{1}{|\mathcal{D}|}\sum_{({\bm{x}},y_{t})\in\mathcal{D}}\mathbb{I}(y_{t}\in{\mathbb{C}}({\bm{x}})),$ (4)
 where $y_{t}$ is the true label of data point ${\bm{x}}$.
+Indicator function $\mathbb{I}(\cdot)$ returns $1$ if the enclosed condition is true and 00 otherwise.
 Coverage reflects the probability that the true label falls within the prediction set ${\mathbb{C}}({\bm{x}})$.
-For $\mathcal{D}=\mathcal{D}_{f}$, high coverage indicates that the model retains significant information about forget data, suggesting incomplete unlearning.
-Conversely, higher coverage is better for test data $\mathcal{D}_{test}$.
-Given a dataset $\mathcal{D}$, Set Size can be defined as follows:
+For $\mathcal{D}=\mathcal{D}_{f}$, high coverage indicates that the model retains significant information about forget data, suggesting fake unlearning.
+Given a dataset $\mathcal{D}$, Set Size is defined as follows:
 $\text{Set Size}:=\frac{1}{|\mathcal{D}|}\sum_{({\bm{x}},y_{t})\in\mathcal{D}}|{\mathbb{C}}({\bm{x}})|,$ (5)
 where $|{\mathbb{C}}({\bm{x}})|$ represents the prediction set size of a specific instance.
-When $\mathcal{D}=\mathcal{D}_{f}$, a smaller set size implies that the model cannot forget specific information about $\mathcal{D}_{f}$ well.
-In contrast, when $\mathcal{D}=\mathcal{D}_{test}$, a small set size may indicate that the model maintains a strong performance after unlearning.
+When $y_{t}\in{\mathbb{C}}({\bm{x}})$, a small set size indicates that fewer non-ground truth classes are included in the prediction set, reflecting stronger fake unlearning.
 Based on Coverage and Set Size, we introduce the definition of CR for a dataset $\mathcal{D}$ as follows:
@@
 CR balances the information captured by Coverage and Set Size.
-A lower CR value implies successful unlearning on forget data $\mathcal{D}_{f}$, but a significant drop in the model’s utility for $\mathcal{D}_{test}$.
-CR is inspired by conformal prediction which is proposed to assess the model’s behavior on new, unseen data (e.g., validation or test data), not on the training data.
+A lower CR value implies stronger unlearning on a given dataset.
+CR is inspired by conformal prediction which is proposed to assess the model’s behavior on new and unseen data, not on the training data.
 Thus, we emphasize that CR only measures forget data $\mathcal{D}_{f}$ and test data $\mathcal{D}_{test}$.
 MIA Conformal Ratio (MIACR).
-To address the limitations of the current MIA metric, we propose a new metric called MIACR.
-For three potential prediction sets {0}, {1} and {0, 1}, set {0} is the only ideal case for MIA prediction since ‘1’ represents the data point is part of the training data.
-Therefore, we introduce a new metric to measure the probability of the set {0} occurring, referred to as MIACR:
+MIACR is proposed to address the limitation of the existing MIA metric.
+Among three potential conformal prediction sets {0}, {1} and {0, 1}, only set {0} is ideal case for MIA, because the presence of ‘1’ represents the data point can still be recognised as a training member.
+Therefore, we introduce a new metric MIACR as:
 $\text{MIACR}:=\frac{1}{|\mathcal{D}_{f}|}\sum_{({\bm{x}},y_{t})\in\mathcal{D}_{f}}\mathbb{I}({\mathbb{C}}({\bm{x}})=\{0\}),$ (7)
 where ${\mathbb{C}}({\bm{x}})=\{0\}$ denotes prediction set is $\{0\}$.
-MIACR only calculates the average of ${\mathbb{C}}({\bm{x}})=\{0\}$, and a higher MIACR value represents better forgetting for $\mathcal{D}_{f}$.
+A higher MIACR score indicates a stronger forgetting.
+While MIA consider a data point forgotten once the logit for label ‘0’ exceeds that for label ‘1’.
+MIACR applies a stricter rule by also requiring that label ‘1’ be absent from the conformal prediction set, giving a more rigorous check of membership status and forgetting quality.
 Superiority of Our Metrics.
-Unlike existing metrics, both the CR and MIACR metrics incorporate the fact that the true labels of some forgotten data points may remain within the prediction set, even when the predictions deviate from the true labels.
-This allows for a more robust assessment of the model’s forgetting quality and predictive performance.
-Evaluation Criteria.
-Similar to existing unlearning metrics, we consider two different criteria to measure unlearning performance with our metrics.
-➊ Gap to RT Criterion: A lower gap to RT method is better for both CR and MIACR metrics.
-The performance gap relative to the RT method is represented in ().
-➋ Limit-Based Criterion: For the CR metric, a lower CR value of forget data $\mathcal{D}_{f}$ indicates better unlearning performance while a higher CR value of $\mathcal{D}_{test}$ is desirable.
-For the MIACR metric, a higher MIACR value for forget data ($\mathcal{D}_{f}$) reflects improved unlearning effectiveness.
-2.3.2 Confidence Level and Calibration Data
-Figure 1: The stability of $\hat{q}$ in different calibration set sizes (results of RT method on CIFAR-10 with ResNet-18).
-When the calibration set size is greater than $2,000$, the fluctuation of $\hat{q}$ can be kept within a certain range.
-In conformal prediction, miscoverage rate $\alpha$ and calibration set size are two crucial parameters for determining prediction set coverage and reliability.
-We next discuss the suitable settings for these two parameters and the rationale behind our chosen configuration.
-In terms of $\alpha$, a higher confidence level (which is $1-\alpha$) provides more reliable coverage but typically results in larger prediction sets.
-Common settings of confidence level are $0.95$ and $0.9$, corresponding to $\alpha$ values of $0.05$ and $0.1$, respectively.
-To make our work more broadly applicable and provide a comprehensive reference, we expand our experiments by setting $\alpha$ to a wider range of values: $0.05$, $0.1$, $0.15$, and $0.2$.
-Notably, our default and recommended value for $\bm{\alpha}$ is $\bm{0.1}$ and subsequent experimental analysis is based on $\bm{\alpha=0.1}$.
-Higher values of $\alpha$ would result in smaller, potentially more precise prediction sets, but this reduction comes at the cost of decreased confidence, which may not be acceptable in many applications.
-Table 4: Unlearning performance of 9 unlearning methods on CIFAR-10 with ResNet-18 in $10\%$ random data forgetting scenario.
-The sign $\uparrow$ represents the greater is better, while $\downarrow$ denotes ideally small.
-The results are average values from $3$ independent trials and the standard deviation values are reported in Appendix B.
-The performance gap relative to the RT method is represented in () in results tables.
-It shows the unlearning methods that excel under traditional metrics (UA, RA and TA) do not necessarily perform well under the CR metric.
-Coverage Set Size CR
-Methods $\alpha$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{test}\uparrow$ $\mathcal{D}_{f}\uparrow$ $\mathcal{D}_{test}\downarrow$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{test}\uparrow$ $\hat{q}$
-RT 0.1 $0.881({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.895({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.934({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.947({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.943({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.945({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.192$
-[…]
-FT 0.1 $0.968({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.087})$ $0.899({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $0.969({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.035})$ $0.924({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.023})$ $0.998({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.972({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$ $0.079$
-[…]
-RL 0.1 $0.913({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.897({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.975({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.041})$ $0.980({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.033})$ $0.936({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.916({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.572$
-[…]
-GA 0.1 $0.990({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.109})$ $0.905({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.990({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.056})$ $0.928({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $0.998({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.973({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.028})$ $0.062$
-[…]
-Teacher 0.1 $0.967({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$ $0.898({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.963({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.929({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.018})$ $0.998({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.969({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $0.591$
-[…]
-FF 0.1 $0.933({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.899({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $7.129({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.195})$ $6.566({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.619})$ $0.131({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.812})$ $0.137({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.808})$ $0.998$
-[…]
-SSD 0.1 $0.987({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.106})$ $0.902({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.990({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.056})$ $0.926({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $0.998({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.973({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.028})$ $0.063$
-[…]
-NegGrad+ 0.1 $0.895({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.014})$ $0.898({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.964({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.030})$ $0.950({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.928({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.946({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.044$
-[…]
-Salun 0.1 $0.936({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.896({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.956({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.022})$ $0.954({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.979({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.939({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.489$
-[…]
-As for calibration data, it should be independent of train and test data.
-One requirement for the choice of the calibration set size is that the distribution of the calibration set should match the distribution of the test data as closely as possible.
-That means a small sample size may lead to distributional dissimilarity and unstable coverage estimation.
-Therefore, a sufficient sample size of calibration data must be ensured to obtain stable estimates.
+The existing accuracy-based metrics UA and MIA suffer from a fake unlearning issue, as true labels of misclassified data points may still remain within the prediction set.
+In contrast, our metrics CR and MIACR eliminate this issue by examining the entire conformal prediction set, yielding a firmer evaluation of forgetting quality and predictive reliability.
+Besides evidence in Tables 2 and  6, Figures 710 in Appendix also support this claim.
+The results show the probability of true labels for data recognized as forgotten by CR is consistently closer to 0 when compared to UA, confirming the stricter criterion.
+Evaluation Criteria of Our Metrics Similar to existing unlearning metrics, we consider two different criteria [^1] The appropriate evaluation criteria vary across unlearning application scenarios Kurmanji et al.
+(2023): criterion ➊ is particularly relevant for user privacy scenario, while criterion ➋ focuses on bias removal scenario.to measure unlearning performance with our metrics, Gap to RT Criterion: A lower gap to the RT method is better for both CR and MIACR metrics.
+The gap relative to RT is represented in blue text (•) in our result tables.
+Limit-Based Criterion: For the CR, a lower CR value of forget data $\mathcal{D}_{f}$ indicates stronger forgetting performance while a higher CR value of $\mathcal{D}_{test}$ represents higher preserved model utility.
+For the MIACR, a higher MIACR value for $\mathcal{D}_{f}$ reflects better unlearning effectiveness.
+2.3.2 Discussion of Confidence Level and Calibration Set Size
+In conformal prediction, the confidence level $1-\alpha$ (i.e., miscoverage rate $\alpha$) and calibration set size are two crucial factors for prediction set coverage and reliability.
+We next discuss the suitable settings for the confidence level and calibration set size, and the rationale behind them.
+Figure 2: The stability of $\hat{q}$ in different calibration set sizes. When the calibration set size is greater than $2000$, the fluctuations of $\hat{q}$ remain within a stable range.
+Confidence Level $\bm{1-\alpha}$.
+A smaller miscoverage rate $\alpha$, i.e., a higher confidence level $1-\alpha$, guarantees more reliable coverage, while a higher value of $\alpha$ results in a lower confidence level.
+In the conformal prediction related works Angelopoulos and Bates (2021); Papadopoulos et al.
+(2002); Romano et al.
+(2020); Tailor et al.
+, $\alpha=0.05$ has become the standard in most cases.
+This value can be rooted in its widespread adoption in statistical hypothesis testing, where it balances the trade-off between false positives and practical usability.
+Following previous work, we set $\alpha$ to $0.05$, but we still report results for a higher range of $\alpha$ values ($0.10$, $0.15$, and $0.20$) in Appendix B.3 and B.4 to account for scenarios where a more relaxed confidence level may be needed.
+Unless otherwise stated, all subsequent analyzes use the recommended default of $\alpha=0.05$.
+Calibration Set Size.
+As for calibration data, a portion of the validation data is set aside as calibration data, ensuring it remains independent from both the training and test data.
+A key requirement in selecting the calibration set size is to avoid producing an abnormal threshold $\hat{q}$, as a small sample size may introduce outliers and lead to unstable coverage estimates.
+Therefore, it is essential to ensure a sufficiently calibration set to obtain reliable and stable estimates.
 Figure 2 illustrates the stability of $\hat{q}$ across varying calibration set sizes.
+The results are smoothed using B-spline.
 We implement them on CIFAR-10 with ResNet-18 in $10\%$ and $50\%$ random data forgetting scenarios.
-The results shown in Figure 1 are smoothed using B-spline.
-It reveals that $2,000$ is a sufficient calibration set size for different settings on CIFAR-10 with ResNet-18 to get a stable threshold $\hat{q}$.
-We also analyze the calibration set size choice of the class-wise forgetting scenario and find that $50$ data points per class as a calibration set is enough.
-The reason is that, the model’s prediction distribution becomes more complex in a dataset with more classes, requiring a larger calibration set to stabilize the estimation.
-2.4 Evaluating Unlearning Methods with New Metrics
-Having established the new metrics CR and MIACR, a natural question to ask is:
-(Q3) Do existing machine unlearning methods truly achieve forgetting, and how do they perform under our new evaluation metrics?
-In this section, we assess the performance of various unlearning methods using the newly introduced metrics, specifically CR, together with coverage and set size.
-We employ 9 different unlearning methods, RT, FT (Warnecke et al., 2021), RL (Graves et al., 2021), Gradient Ascent (GA) (Thudi et al., 2022), Teacher (Tarun et al., 2023), FisherForgetting (FF) (Golatkar et al., 2020a), SSD (Foster et al., 2024), NegGrad+ (Kurmanji et al., 2024) and Salun (Fan et al., 2023).
-The experimental results are presented in Table 4, which summarizes the unlearning performance under $10\%$ random data forgetting scenario on CIFAR-10 with ResNet-18.
-For more experimental results on CR and MIACR, please check Section 4 and Appendix B.
-The table 4 includes coverage percentages, set size, CR for both $\mathcal{D}_{f}$ and $\mathcal{D}_{test}$, as well as $\hat{q}$ in different settings of $\alpha$.
-As observed, it shows a decreasing coverage and set size as the alpha increases in all methods since the threshold $\hat{q}$ is relaxed due to the reduced confidence level.
-Based on evaluation criterion ➊, while NegGrad+ outperforms RL under the traditional UA metric, RL exhibits superior performance under the CR metric.
-This observation suggests that methods excelling in traditional UA metric may not perform well under the CR metric.
-The underlying rationale behind this is that the CR metric takes into account the possibility that the true labels of some misclassified forget data points may still remain within the prediction set.
-This observation aligns with the insights we discussed in Section 2.2 regarding the limitation of UA.
-Based on evaluation criterion ➋, NegGrad+ and RL exhibit the best trade-off between forgetting quality and model performance.
-FF achieves the lowest CR, indicating the highest forgetting quality, but it exhibits the worst prediction performance on $\mathcal{D}_{test}$.
-On the other hand, GA, Teacher, SSD, FT and Salun maintain relatively high CR of the forget data $\mathcal{D}_{f}$ across different alpha settings, reflecting their poor forgetting quality.
+The results show that for different settings using ResNet-18 on CIFAR-10, after the calibration set size is larger than $1000$, abnormal $\hat{q}$ values do not occur anymore, and a stable threshold $\hat{q}$ can be obtained.
+Similarly, we analyze the calibration set size of the class-wise forgetting scenario and find that fewer calibration data points are required compared to random data forgetting.
+This is because the targeted removal of specific classes reduces the complexity of the distribution, unlike the broader variability introduced by random data removal.
 3 Enhancing Machine Unlearning via Conformal Prediction
-Our analysis indicates that machine unlearning methods generally do not exhibit strong performance under our newly developed evaluation metrics.
-Even methods with relatively low CR values demonstrate significant compromises in other metrics.
-This observation raises an important question:
-(Q4) Can we explore advanced learning techniques via conformal prediction to optimize unlearning model’s performance under our evaluation metrics?
-A training-based machine unlearning method is typically optimized for prediction loss but cannot efficiently support the improvement of forgetting quality in our new metrics.
-Therefore, we propose a novel and general unlearning framework for training-based machine unlearning methods to enhance machine unlearning.
-The key innovation of this framework lies in leveraging conformal prediction principles combined with a loss function inspired by the C&W attack (Carlini & Wagner, 2017) to improve forgetting quality.
-The original C&W loss function is designed to find adversarial examples that cause misclassification.
-Here we extend it to an unlearning loss.
-For forget data $\mathcal{D}_{f}$, the objective of the unlearning loss function is to decrease the model’s confidence in the true labels of $\mathcal{D}_{f}$.
-Based on this, the C&W-inspired unlearning loss can be defined as:
-${\mathcal{L}}_{\text{cw}}({\bm{x}},y_{t})=\max\left(p_{t}({\bm{x}})-\max_{i\neq t}\{p_{i}({\bm{x}})\},\ -\Delta\right),$ (7)
-where $({\bm{x}},y_{t})\in\mathcal{D}_{f}$.
-$p_{i}({\bm{x}})$ is the possibility (after softmax function) of class $y_{i}$, while $z_{t}({\bm{x}})$ is the true label $y_{t}$’s possibility output.
-We denote $\max_{i\neq t}\{p_{i}({\bm{x}})\}$ as the highest possibility value of the incorrect class.
-$\Delta$ is the margin parameter (also known as the confidence parameter) that controls the attack strength.
-This loss is minimized when $p_{t}({\bm{x}})-\max_{i\neq t}\{p_{i}({\bm{x}})\}\leq-\Delta$ to make $\max_{i\neq t}\{p_{i}({\bm{x}})\}-p_{t}({\bm{x}})\geq\Delta$.
-It means that the loss maximizes the difference between the highest possibility value for class $y_{i}$ ($i\neq t$) and the possibility value for the true label $y_{t}$, which makes it more difficult for the model to include the true label in the prediction set during conformal prediction process.
-Additionally, increasing the value of $\Delta$ further reduces the confidence of the true label $y_{t}$ being included in the prediction set.
-In our work, we set $\Delta=0.01$.
-Based on the insight discussed in Section 2, we further improve the C&W-inspired unlearning loss function by combining conformal prediction.
+Based on the findings in Table 2 in Section 2.1 and Table 6 in Appendix B.1, we observe that existing training-based unlearning methods are typically optimized with respect to loss functions that do not directly support the improvement of forgetting quality from our fake unlearning perspective.
+Specifically, the optimization objectives of existing methods fail to ensure that the ground truth labels are sufficiently pushed out of the conformal prediction set, which is key to overcoming fake unlearning.
+This raises a critical question:
+(Q4) Can we explore advanced unlearning techniques via conformal prediction to optimize existing unlearning model’s forgetting quality?
+Therefore, we propose a novel and general conformal prediction-based unlearning framework (CPU) tailored for training-based unlearning methods, aimed at enhancing their forgetting quality.
+A key insight driving our framework is to overcome the issue exposed by fake unlearning.
+This emphasizes that the non-conformity scores of ground truth labels should be pushed as far beyond the conformal prediction threshold $\hat{q}$ as possible.
+Interestingly, this goal aligns naturally with the design of the C&W attack loss Carlini and Wagner (2017), which motivates our creative adaptation to the unlearning scenario.
+Let us first apply the original C&W loss directly to the unlearning scenario, without yet incorporating conformal prediction.
+For forget data $\mathcal{D}_{f}$, the goal of the unlearning loss is to decrease the model’s confidence in the true labels of $\mathcal{D}_{f}$.
+Based on this, the C&W-inspired unlearning loss is defined as:
+${\mathcal{L}}_{\text{cw}}({\bm{x}},y_{t})=\max\{p_{t}({\bm{x}})-\max_{i\neq t}\{p_{i}({\bm{x}})\},\ -\Delta\},$ (8)
+where $({\bm{x}},y_{t})\in\mathcal{D}_{f}$ and ${}\max\{\cdot\}$ is a maximum operator that selects the largest value from the set.
+$p_{i}({\bm{x}})$ is the probability of class $y_{i}$, and $p_{t}({\bm{x}})$ refers specifically to the probability assigned to the true label $y_{t}$.
+We denote $\max_{i\neq t}\{p_{i}({\bm{x}})\}$ as the highest probability value of the non-ground truth classes.
+This loss ${\mathcal{L}}_{\text{cw}}$ maximizes the difference between the highest probability value for class $y_{i}\ (i\neq t)$ and the probability value for the true class $y_{t}$.
+It tries to decrease the probability of true class $y_{t}$ and further increase that of the class $y_{i}$ with the highest probability.
+The margin parameter $\Delta$ controls the enforced margin between the true class and the strongest competing class.
+When the $\max_{i\neq t}\{p_{i}({\bm{x}})\}-p_{t}({\bm{x}})<\Delta$, this loss encourages the model to decrease the true label’s probability $p_{t}({\bm{x}})$.
+Increasing the value of $\Delta$ further increase the margin between $\max_{i\neq t}\{p_{i}({\bm{x}})\}$ and $p_{t}({\bm{x}})$.
+With this C&W loss, we can indeed reduce the probability assigned to the true label $y_{t}$, thereby compelling the model to misclassify the data point into another class $y_{i}$.
+However, this loss still fails to guarantee that the true label $y_{t}$ can be excluded from the conformal prediction set.
+If we let the threshold in conformal prediction play the role of $\max_{i\neq t}\{p_{i}({\bm{x}})\}$ in Eq.
+8, and push non-conformity score of $y_{t}$ further away from this threshold, the above issue can be effectively resolved.
+Therefore, we further improve the C&W-inspired unlearning loss function by combining conformal prediction.
 In conformal prediction, calibration data helps in estimating non-conformity scores and determining a threshold to ensure valid statistical guarantees about the model’s uncertainty estimates.
-Therefore, we reserve a portion of calibration data $\mathcal{D}_{c}^{\prime}$ for the unlearning phase, which is kept separate from the calibration data $\mathcal{D}_{c}$ used in the evaluation phase.
-According to Eq.
-1, calibration data $\mathcal{D}_{c}^{\prime}$ is used to calculate non-conformity scores and generate a threshold $\bar{q}$ based on an $\alpha$.
-Then, we calculate the non-conformity scores of $\mathcal{D}_{f}$ to obtain the corresponding prediction set.
-Then, by revising C&W-inspired unlearning loss with this calibration step, a general unlearning loss function is defined as follows:
-${\mathcal{L}}_{\text{unlearn}}({\bm{x}},y_{t})=\max\left(S({\bm{x}},y_{t})-\bar{q},\ -\Delta\right),$ (8)
-where $({\bm{x}},y_{t})\in\mathcal{D}_{f}$.
-We replace $p_{t}({\bm{x}})$ in Eq.
-7 with the non-conformity score $S({\bm{x}},y_{t})$ of true label $y_{t}$, and replace $\max_{i\neq t}\{p_{i}({\bm{x}})\}$ with $\bar{q}$ which will be updated before each epoch.
-It ensures that ${\mathcal{L}}_{\text{unlearn}}$ adheres to the principle of ${\mathcal{L}}_{\text{cw}}$, which encourages $\bar{q}-S({\bm{x}},y_{t})\geq\Delta$.
-It is a general framework applicable across various training-based machine unlearning methods.
-To preserve the effectiveness of specific machine unlearning methods, we retain their original loss ${\mathcal{L}}_{\text{original}}$.
-Consequently, we sum these terms to form the final objective loss function as follows:
+A portion of calibration data $\mathcal{D}_{c}^{\prime}$ can be reserved for the unlearning phase, which is kept separate from the calibration data $\mathcal{D}_{c}$ used in the evaluation phase.
+With calibration data $\mathcal{D}_{c}^{\prime}$, the threshold $\bar{q}$ for the unlearning phase is easily calculated given an $\alpha$.
+Given $\bar{q}$, by revising C&W-inspired unlearning loss with a calibration step, a general unlearning loss function is defined as follows:
+${\mathcal{L}}_{\text{unlearn}}({\bm{x}},y_{t})=\max\{\bar{q}-S({\bm{x}},y_{t}),\ -\Delta\}.$ (9)
+We replace probability $p_{t}({\bm{x}})$ and $\max_{i\neq t}\{p_{i}({\bm{x}})\}$ in Eq.
+8 with the threshold $\bar{q}$ and non-conformity score $S({\bm{x}},y_{t})$ respectively.
+The threshold $\bar{q}$ is updated in each training epoch to obtain an accurate value.
+Since $\bar{q}$ is computed merely as a quantile, this process incurs negligible computational overhead.
+The loss ${\mathcal{L}}_{\text{unlearn}}$ adheres to the same principle of ${\mathcal{L}}_{\text{cw}}$, which encourages $S({\bm{x}},y_{t})-\bar{q}\geq\Delta$.
+It helps to increase non-conformity score $S({\bm{x}},y_{t})$ of true label $y_{t}$ to surpass the threshold $\bar{q}$.
+As an improvement over the loss ${\mathcal{L}}_{\text{cw}}$, the loss ${\mathcal{L}}_{\text{unlearn}}$ makes it more difficult for the model to include the true label in conformal prediction set.
+In this loss, even a small value of $\Delta$ is sufficient to achieve the desired effect, because the true label $y_{t}$ is excluded from the conformal prediction set once its non-conformity score $S({\bm{x}},y_{t})$ exceeds the threshold $\bar{q}$.
+Therefore, in our work, we set $\Delta=0.01$.
+As a general framework, to preserve the efficacy of specific unlearning methods themselves, we reserve their original loss ${\mathcal{L}}_{\text{original}}$ in our framework.
+Consequently, we combine these terms to form the final objective loss function as:
 ${\mathcal{L}}_{\text{total}}={\mathcal{L}}_{\text{original}}+\lambda\cdot{\mathcal{L}}_{\text{unlearn}},$ (10)
@@
 Datasets and Models.
-We report experiments on CIFAR-10 (Krizhevsky, 2009) with ResNet-18 (He et al., 2016) and Tiny ImageNet (Le & Yang, 2015) with ViT (Dosovitskiy et al., 2021).
-Table 5: Performance of 8 unlearning methods on Tiny ImageNet with ViT in $10\%$ random data forgetting scenario.
-Coverage Set Size CR
-Methods $\alpha$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{test}\uparrow$ $\mathcal{D}_{f}\uparrow$ $\mathcal{D}_{test}\downarrow$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{test}\uparrow$ $\hat{q}$
-RT 0.1 $0.892({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.900({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.151({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.144({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.775({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.786({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.853$
+For both random data forgetting and class-wise forgetting, we report experiments on CIFAR-10 Krizhevsky (2009) with ResNet-18 He et al.
+(2016) and Tiny ImageNet Le and Yang (2015) with ViT Dosovitskiy et al.
+(2021).
+See Appendix B for additional experimental results of class-wise forgetting scenario.
+Baselines and Metrics.
+We employ 9 different unlearning methods, including RT, FT Warnecke et al.
+(2021), RL Graves et al.
+(2021), Gradient Ascent (GA) Thudi et al.
+(2022), Bad Teacher (Teacher) Tarun et al.
+(2023), SSD Foster et al.
+(2024), NegGrad+ Kurmanji et al.
+(2023), Salun Fan et al.
+(2024) and SFRon Huang et al.
+(2025).
+We evaluate the performance of various unlearning methods using the existing metrics, including UA, RA, TA, MIA, as well as our proposed metrics CR and MIACR.
+Implementation Details.
+For the hyperparameter in our work, miscoverage rate $\alpha\in[0.05,0.1,0.15,0.2]$, margin parameter $\Delta=0.01$, unlearning loss weight $\lambda\in[0,0.2,0.5,1]$.
+See Appendix A-C for more training details and experimental results under settings of $\alpha\in[0.1,0.15,0.2]$, $\lambda=1$ and other forgetting scenarios.
+Table 3: Unlearning performance of 9 unlearning methods on CIFAR-10 with ResNet-18 and Tiny ImageNet with ViT in $10\%$ random data forgetting.
+The results are average values from $3$ independent trials and the standard deviation values are reported in Appendix B.
+For evaluation criterion ➊, performance differences compared to the RT method are highlighted with ().
+For clarity in observing criterion ➋, the sign $\uparrow$ represents greater is better, while $\downarrow$ denotes ideally small.
+It shows the unlearning methods that excel under the existing metric UA do not necessarily perform well under our CR metric due to the fake unlearning issue.
+Existing Metrics Coverage Set Size CR
+Methods UA $\uparrow$ RA $\uparrow$ TA $\uparrow$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{test}\uparrow$ $\mathcal{D}_{f}\uparrow$ $\mathcal{D}_{test}\downarrow$ $\mathcal{D}_{f}\downarrow$ $\mathcal{D}_{test}\uparrow$
+CIFAR-10 with ResNet-18
+RT $8.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $99.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $91.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $0.941({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.944({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.089({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.074({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.864({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.879({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$
+FT $3.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.8})$ $98.1\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.6})$ $91.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.2})$ $0.994({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.951({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $1.008({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.081})$ $1.026({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.048})$ $0.986({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.122})$ $0.927({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.048})$
+RL $7.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.0})$ $97.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.3})$ $90.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.2})$ $0.970({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.949({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $1.242({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.153})$ $1.197({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.123})$ $0.788({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.076})$ $0.796({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.083})$
+GA $0.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}8.0})$ $99.5\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.2})$ $94.1\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.3})$ $0.994({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.945({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $1.002({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.087})$ $1.009({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.065})$ $0.994({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.130})$ $0.936({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.057})$
+Teacher $0.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}7.8})$ $99.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.3})$ $93.5\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.7})$ $0.991({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.050})$ $0.941({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $1.003({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$ $1.021({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.993({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.129})$ $0.922({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.043})$
+SSD $0.5\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}8.1})$ $99.5\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.2})$ $94.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.996({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.945({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.999({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.090})$ $1.008({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.066})$ $0.994({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.130})$ $0.936({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.057})$
+NegGrad+ $8.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.1})$ $98.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $92.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.4})$ $0.934({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.948({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $1.068({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $1.086({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.875({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.873({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$
+Salun $3.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.9})$ $98.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.8})$ $91.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$) $0.987({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.046})$ $0.950({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $1.132({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.043})$ $1.143({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.069})$ $0.872({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.832({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$
+SFRon $4.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.8})$ $97.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.3})$ $91.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.4})$ $0.977({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.953({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $1.100({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $1.143({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.069})$ $0.889({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.025})$ $0.834({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.045})$
+Tiny ImageNet with ViT
+RT $14.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $98.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $86.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $0.944({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.949({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.876({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.840({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.503({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.516({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$
+FT $6.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}7.8})$ $97.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $84.1\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.9})$ $0.994({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.050})$ $0.950({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $2.133({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.257})$ $2.440({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.600})$ $0.466({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.037})$ $0.389({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.127})$
+RL $26.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}12.2})$ $96.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.8})$ $81.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.6})$ $0.969({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.025})$ $0.952({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $17.890({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}16.014})$ $8.572({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.732})$ $0.054({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.449})$ $0.111({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.405})$
+GA $3.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}11.5})$ $97.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.4})$ $84.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.1})$ $0.996({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.947({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $1.539({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.337})$ $2.018({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.178})$ $0.647({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.144})$ $0.469({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$
+Teacher $17.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.6})$ $86.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}12.1})$ $79.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}7.0})$ $0.977({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.033})$ $0.956({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $5.473({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.597})$ $5.080({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.240})$ $0.179({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.324})$ $0.188({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.328})$
+SSD $1.5\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}13.2})$ $98.5\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.3})$ $86.1\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.1})$ $0.998({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.054})$ $0.950({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $1.354({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.522})$ $1.827({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.013})$ $0.737({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.234})$ $0.520({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$
+NegGrad+ $19.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.7})$ $98.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.5})$ $84.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.0})$ $0.999({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.890({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.059})$ $0.949({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.927})$ $1.614({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.227})$ $2.184({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.681})$ $2.499({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.984})$
+Salun $9.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.5})$ $97.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.1})$ $83.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.995({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.051})$ $0.964({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $2.803({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.927})$ $2.726({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.886})$ $1.311({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.808})$ $1.157({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.641})$
+SFRon $9.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.4})$ $97.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.8})$ $83.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.1})$ $0.989({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.045})$ $0.948({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $2.000({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.124})$ $2.208({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.368})$ $0.495({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.429({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$
+4.2 Measure Unlearning Methods via New Metrics
+In this section, we explore how existing unlearning methods perform with the consideration of the fake unlearning perspective.
+We evaluate the performance of 9 various unlearning methods using the proposed metrics CR and MIACR, together with Coverage and Set Size.
+The experimental results are presented in Table 3, which summarizes the unlearning performance under $10\%$ random data forgetting scenario on CIFAR-10 and Tiny ImageNet, respectively.
+See Tables 7 - 13 in Appendix B for additional experimental results on other forgetting scenarios.
+CR Metric.
+We take the results on CIFAR-10 as an example for analysis of CR on forget data $\mathcal{D}_{f}$.
+According to evaluation criterion ➊, the top 4 methods under the UA metric are NegGrad+, RL, SFRon, and Salun, as their unlearning accuracy is closest to the RT method.
+However, this ranking shifts slightly under the CR metric, where the top 4 become Salun, NegGrad+, SFRon, and RL.
+CR metric identifies that Salun performs better in forgetting quality and can deal with the fake unlearning issue well, while RL faces a fake unlearning situation and performs poorly on our metric CR.
+This observation suggests that methods excelling in traditional UA metric may not perform well under the CR metric.
+The underlying rationale behind this is that the CR metric takes into account the possibility that the true labels of some misclassified forget data points may still remain within the prediction set.
+This observation aligns with the insights we discussed in Section 2.2 regarding the fake unlearning issue of UA metric.
+Regarding evaluation criterion ➋, a similar pattern is observed as with criterion ➊.
+Under the UA metric, the top 4 methods in terms of forgetting quality are NegGrad+, RT, RL and SFRon.
+However, under the CR metric, the top 4 shift to RL, RT, Salun and NegGrad.
+This indicates that some unlearning methods, such as NegGrad+ show weak forgetting quality when viewed from the fake unlearning perspective.
+This also highlights that the CR captures critical scenarios overlooked by UA, specifically the potential retention of true labels within prediction sets for the forget data points.
+CR ensures a more robust and reliable evaluation for unlearning quality.
+Table 4: MIACR results on CIFAR-10 with ResNet-18 in both 10% and 50% random data forgetting.
+$10\%$ Forgetting $50\%$ Forgetting
 […]
-FT 0.1 $0.978({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$ $0.903({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $1.234({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.083})$ $1.317({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.173})$ $0.792({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.017})$ $0.685({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.101})$ $0.935$
+RT $86.92({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.091({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $82.79({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.117({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$
+FT $92.00({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.08})$ $0.039({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $92.92({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}10.13})$ $0.036({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.081})$
+RL $74.21({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}12.71})$ $0.083({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $61.15({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}21.64})$ $0.050({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.067})$
+GA $98.80({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}11.88})$ $0.012({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.079})$ $98.86({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}16.07})$ $0.012({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.105})$
+Teacher $87.24({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.32})$ $0.013({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.078})$ $93.24({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}10.45})$ $0.031({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$
+SSD $98.78({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}11.86})$ $0.011({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.080})$ $98.87({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}16.08})$ $0.012({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.105})$
+NegGrad+ $90.30({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.38})$ $0.076({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $93.82({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}11.03})$ $0.045({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.072})$
+Salun $57.58({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}29.34})$ $0.055({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $59.12({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}23.67})$ $0.044({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.073})$
+SFRon $91.55({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.63})$ $0.017({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.074})$ $92.52({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}9.73})$ $0.017({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.100})$
+MIACR Metric.
+In Table 4, we show the MIACR results on CIFAR-10 under both $10\%$ and $50\%$ random data forgetting.
+Under evaluation criterion ➊, most methods show superior MIA and MIACR performance in the $10\%$ forgetting scenario compared to $50\%$ forgetting, because larger forget sets pose greater challenges for unlearning methods.
+This demonstrates that the general trend of membership leakage risk remains broadly consistent across MIA and MIACR.
+However, the unlearning methods considered superior under MIA metric, such as Teacher and SFRon, fail to maintain their exceptional performance when evaluated using the MIACR criterion.
+Similarly, under evaluation criterion ➋, Salun, which is considered optimal according to MIA metrics, does not demonstrate the best performance under MIACR evaluation.
+While MIA fails to accurately predict approximately $50\%$ of the forget data in Salun, MIACR can still identify membership within the prediction set.
+Overall, the results show that, compared to MIACR, the existing MIA metric still leaves privacy concerns.
+Although MIA may fail to predict some forget data points as training members, these points can still appear in the conformal prediction set with high confidence.
+In contrast, MIACR more strictly controls potential membership leakage risk by measuring the probability that only non-member predictions (i.e., label ‘0’) appear in the prediction set.
+4.3 Performance of Our Unlearning Framework
+In this experiment, we apply RT, FT, and RL methods to our framework CPU, i.e., CPU-RT, CPU-FT, CPU-RL.
+Table 5 presents the results for CIFAR-10 with ResNet-18 and Tiny ImageNet with ViT in $10\%$ random data forgetting.
+We vary $\lambda$ in the range $[0,0.2,0.5,1]$, where $\lambda=0$ represents the baseline without our framework applied.
+See Table 14 in Appendix C for the results of $\lambda=1$.
 […]
-RL 0.1 $0.892({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.902({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $2.639({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.488})$ $1.843({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.699})$ $0.338({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.437})$ $0.489({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.297})$ $0.971$
-[…]
-GA 0.1 $0.986({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.094})$ $0.900({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.104({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$ $1.224({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.080})$ $0.894({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.119})$ $0.736({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.050})$ $0.899$
-[…]
-Teacher 0.1 $0.930({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.038})$ $0.902({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $1.991({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.840})$ $1.959({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.815})$ $0.467({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.308})$ $0.460({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.326})$ $0.971$
-[…]
-SSD 0.1 $0.993({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.101})$ $0.897({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $1.039({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.112})$ $1.134({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.956({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.181})$ $0.791({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.852$
-[…]
-NegGrad+ 0.1 $0.995({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.103})$ $0.848({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.898({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.253})$ $1.093({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.051})$ $1.225({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.450})$ $1.287({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.501})$ $0.933$
-[…]
-Salun 0.1 $0.977({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.085})$ $0.924({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $1.229({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.078})$ $1.281({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.137})$ $0.918({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.143})$ $0.884({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.097})$ $0.939$
-[…]
-Table 6: MIACR performance on CIFAR-10 with ResNet-18. We show the MIA value in $10\%$ random data forgetting scenario, while the MIA results in $50\%$ scenario are presented in Appendix B.
-$10\%$ Forgetting $50\%$ Forgetting
-Methods $\alpha$ MIACR $\uparrow$ $\hat{q}$ MIACR $\uparrow$ $\hat{q}$
-RT 0.1 $0.147({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.589$ $0.201({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.570$
-[…]
-FT 0.1 $0.077({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.070})$ $0.627$ $0.103({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.098})$ $0.558$
-[…]
-RL 0.1 $0.178({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$ $0.572$ $0.137({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.064})$ $0.547$
-[…]
-GA 0.1 $0.032({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.115})$ $0.502$ $0.055({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.146})$ $0.486$
-[…]
-Teacher 0.1 $0.038({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.109})$ $0.672$ $0.065({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.136})$ $0.582$
-[…]
-FF 0.1 $0.051({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.096})$ $0.486$ $0.089({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.112})$ $0.509$
-[…]
-SSD 0.1 $0.031({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.116})$ $0.511$ $0.051({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.150})$ $0.488$
-[…]
-NegGrad+ 0.1 $0.128({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $0.481$ $0.109({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.092})$ $0.511$
-[…]
-Salun 0.1 $0.113({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.034})$ $0.681$ $0.115({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$ $0.630$
-[…]
-Implementation Details.
-For CIFAR-10/Tiny ImageNet, we randomly select $200$/$50$ data points per class ($2,000$/$10,000$ data points in total) as calibration data $\mathcal{D}_{c}$ and $\mathcal{D}_{c}^{\prime}$, respectively.
-The calibration data $\mathcal{D}_{c}$ does not participate in the model training or unlearning processes and is only used for calibrating the threshold $\hat{q}$, while $\mathcal{D}_{c}^{\prime}$ is used in the process of our unlearning framework to generate $\bar{q}$.
-For the hyperparameter in our work, miscoverage rate $\alpha\in[0.05,0.1,0.15,0.2]$, margin parameter $\Delta=0.01$, unlearning loss weight $\lambda\in[0,0.2,0.5,1]$.
-Due to space limitations, more training details and experimental results under settings of $\alpha\in[0.05,0.15]$, $\lambda=0.2$ and class-wise forgetting scenario can be found in Appendix A-C.
-4.2 Measure Machine Unlearning Method via New Metrics
-CR Metric.
-For the CR metric, we already show the results of CIFAR-10 with ResNet-18 in Tabel 4.
-Here, we further evaluate the unlearning performance in Tiny ImageNet with ViT under $10\%$ random data forgetting scenario by CR metric.
-The results of Tiny ImageNet with ViT and CIFAR-10 with ReNet18 in $50\%$ random data forgetting and class-wise forgetting scenarios can be found in Appendix B.
-We remark that it is computationally intractable for the FF method in Tiny ImageNet with ViT, so we omit its results.
-The results of Tiny ImageNet with ViT in $10\%$ random data forgetting scenario are summarized in Table 5.
-Our findings reveal the following insights:
-For all machine unlearning methods, as $\alpha$ level increases, it results in reduced Coverage and smaller Set Size.
-This happens because a higher $\alpha$ loosens the conformal threshold $\hat{q}$, allowing fewer predictions to be included within the prediction set for each data point.
-On the contrary, the CR tends to increase with increasing $\alpha$.
-Although both Coverage and Set Size may decrease, Set Size often decreases more significantly.
-Consequently, the CR value generally becomes larger as $\alpha$ increases.
-When $\alpha$ is set to $0.2$, most methods show a value of Set Size less than $1$ in both Table 4 and 5.
-The intuition behind it is that conformal prediction, as a static predictor, is intrinsically tied to the model’s base prediction performance and accuracy.
-When the model’s accuracy is significantly higher than the confidence level, conformal prediction can achieve the required coverage with ease.
-In fact, it can generate empty prediction sets for some data points while still meeting the target coverage.
-Thus, the choice of $\alpha$ is crucial.
-Overly high $\alpha$ values may skew evaluation results by failing to let CR accurately reflect model performance.
-Notably, when $\alpha=0.2$, RL’s coverage on the forget data $\mathcal{D}_{f}$ decreases significantly compared to other methods.
-This results from the use of label corruption during RL’s unlearning process, which enlarges the distribution difference between calibration data and forget data.
-We can mitigate this gap if we know the details of specific unlearning methods, and a potential solution to this issue is discussed in Appendix B.
-Regarding the forgetting quality of specific machine unlearning methods, the Teacher method achieves the strongest performance on the traditional UA metric under both evaluation criteria ➊ and ➋.
-However, its forgetting quality is significantly weaker when measured by the CR metric.
-This highlights that the CR captures critical scenarios overlooked by UA, specifically the potential retention of true labels within prediction sets for the forget data points.
-By explicitly addressing this phenomenon, CR ensures a more robust and reliable evaluation for unlearning quality.
-Table 7: Performance of our unlearning framework.
-We show the performance on CIFAR-10 with ResNet-18 and Tiny ImageNet with ViT in $10\%$ random data forgetting scenario.
+Table 5: Performance of our unlearning framework CPU.
+We show the performance on CIFAR-10 with ResNet-18 and Tiny ImageNet with ViT in $10\%$ random data forgetting.
 $\lambda=0$ represents the baseline without our framework applied.
-It shows our framework significantly improves the forgetting quality, while preserving stable predictive performance.
+It shows our framework significantly improves the forgetting quality, not only across our metric but also existing metric UA, while preserving stable predictive performance.
 $\lambda=0$ $\lambda=0.2$ $\lambda=0.5$
@@
 CIFAR-10 with ResNet-18
-RT $0.1$ $8.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $99.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $91.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $0.943({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.945({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $10.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.2})$ $98.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.4})$ $91.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.8})$ $0.914({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.924({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $14.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.4})$ $97.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.9})$ $90.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.4})$ $0.879({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.064})$ $0.912({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.033})$
-$0.2$ $0.988({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.981({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.977({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.976({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.963({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.025})$ $0.966({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$
-FT $0.1$ $3.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.8})$ $98.1\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.6})$ $91.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.2})$ $0.998({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.972({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$ $6.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.8})$ $97.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.7})$ $90.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.0})$ $0.948({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.924({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $7.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.7})$ $96.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.8})$ $90.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $0.940({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.927({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.018})$
-$0.2$ $1.000({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.993({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.989({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.974({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.983({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.975({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$
-RL $0.1$ $7.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.0})$ $97.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.3})$ $90.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.2})$ $0.936({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.916({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $9.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.1})$ $96.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.1})$ $89.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.896({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$ $0.887({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.058})$ $9.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.3})$ $96.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.8})$ $89.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.1})$ $0.902({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.041})$ $0.896({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.049})$
-$0.2$ $0.976({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.959({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.022})$ $0.964({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $0.949({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.959({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.950({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$
+CPU-RT $8.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $99.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $91.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $0.864({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.879({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $10.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.2})$ $98.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.4})$ $91.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.8})$ $0.788({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.076})$ $0.824({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $14.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.4})$ $97.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.9})$ $90.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.4})$ $0.763({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.101})$ $0.825({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.054})$
+CPU-FT $3.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.8})$ $98.1\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.6})$ $91.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.2})$ $0.986({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.122})$ $0.927({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.048})$ $6.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.8})$ $97.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.7})$ $90.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.0})$ $0.844({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.020})$ $0.829({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.050})$ $7.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.7})$ $96.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.8})$ $90.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $0.853({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.843({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$
+CPU-RL $7.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.0})$ $97.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.3})$ $90.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.2})$ $0.788({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.076})$ $0.796({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.083})$ $9.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.1})$ $96.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.1})$ $89.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.709({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.155})$ $0.736({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.143})$ $9.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.3})$ $96.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.8})$ $89.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.1})$ $0.708({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.156})$ $0.731({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.148})$
 Tiny ImageNet with ViT
-RT $0.1$ $14.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $98.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $86.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $0.775({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.786({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $19.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.6})$ $98.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $86.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $0.729({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.163})$ $0.786({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.114})$ $26.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}11.7})$ $98.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.1})$ $85.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.2})$ $0.649({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.243})$ $0.765({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.135})$
-$0.2$ $0.934({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.935({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.898({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.108})$ $0.932({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.133})$ $0.839({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.049})$ $0.929({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.130})$
-FT $0.1$ $6.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}7.8})$ $97.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $84.1\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.9})$ $0.791({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.087})$ $0.685({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.201})$ $9.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.9})$ $97.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.4})$ $83.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.753({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.139})$ $0.683({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.217})$ $13.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $97.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.6})$ $83.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.718({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.174})$ $0.683({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.217})$
-$0.2$ $0.969({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.905({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.942({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.152})$ $0.893({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.094})$ $0.914({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.124})$ $0.890({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.091})$
-RL $0.1$ $26.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}12.2})$ $96.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.8})$ $81.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.6})$ $0.338({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.540})$ $0.488({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.398})$ $31.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}17.1})$ $95.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}17.9})$ $80.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.1})$ $0.278({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.614})$ $0.451({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.449})$ $36.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}21.5})$ $95.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.5})$ $80.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.6})$ $0.254({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.638})$ $0.449({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.451})$
-$0.2$ $0.821({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.143})$ $0.848({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.110})$ $0.752({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.038})$ $0.825({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.026})$ $0.718({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.072})$ $0.827({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.028})$
-MIACR Metric.
-In Table 6, we show the MIACR results on CIFAR-10 under both $10\%$ and $50\%$ random data forgetting scenarios.
-Compared to MIACR, traditional MIA remains a privacy concern.
-Albeit MIA cannot predict some forget data points as training data, it still has high confidence in appearing in the prediction set.
-For example, under evaluation criterion ➊, the teacher’s MIA is the method closest to the RT method, but it performs poorly in the CR metric.
-Similarly, FF and Salun are the top two unlearning methods in the traditional MIA metric (ideally small) under evaluation criterion ➋, but FF has low MIACR (ideally greater) values in both 10% and 50% forgetting scenarios.
-4.3 Performance of Our Unlearning Framework
-In this experiment, we apply the RT, FT, and RL methods to our framework. Table 7 presents the results for CIFAR-10 with ResNet-18 and Tiny ImageNet with ViT in $10\%$ random data forgetting scenario.
+CPU-RT $14.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $98.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $86.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $0.503({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.516({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $19.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.6})$ $98.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $86.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $0.458({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.045})$ $0.516({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $26.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}11.7})$ $98.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.1})$ $85.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.2})$ $0.396({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.107})$ $0.489({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$
+CPU-FT $6.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}7.8})$ $97.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $84.1\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.9})$ $0.466({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.037})$ $0.389({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.127})$ $9.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.9})$ $97.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.4})$ $83.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.441({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.062})$ $0.399({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.117})$ $13.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $97.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.6})$ $83.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.413({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.090})$ $0.401({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.115})$
+CPU-RL $26.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}12.2})$ $96.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.8})$ $81.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.6})$ $0.054({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.449})$ $0.111({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.405})$ $31.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}17.1})$ $95.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}17.9})$ $80.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.1})$ $0.051({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.452})$ $0.111({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.405})$ $36.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}21.5})$ $95.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.5})$ $80.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.6})$ $0.051({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.452})$ $0.121({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.395})$
+(a) $\mathcal{D}_{f}$
+(b) $\mathcal{D}_{r}$
+(c) $\mathcal{D}_{test}$
+(d) $\mathcal{D}_{f}$
+(e) $\mathcal{D}_{r}$
+(f) $\mathcal{D}_{test}$
+Figure 3: CPU-FT accuracy of $\mathcal{D}_{f}$, $\mathcal{D}_{r}$ and $\mathcal{D}_{test}$ under different $\lambda$ values across each epoch on CIFAR-10 (a-c) and Tiny ImageNet (d-f).
+As $\lambda$ increases, accuracy on $\mathcal{D}_{f}$ drops significantly, while retain and test accuracy remain stable.
 […]
-From the perspective of evaluation criterion ➊, the FT method achieves an average gap of $1.47$ across UA, RA, and TA on ResNet-18 when $\lambda=0.5$, compared to an average gap of $2.2$ when $\lambda=0$.
-Similarly, on the ViT model, FT reduces the average gap from $3.53$ to $1.63$ when $\lambda=0.5$.
-If UA of an unlearning method is already lower than that of RT, e.g., RL, our framework cannot further improve their performance under evaluation criterion ➊.
+Moreover, in Figure 3, we further present the CPU-FT accuracy on forget data $\mathcal{D}_{f}$, retain data $\mathcal{D}_{r}$ and test data $\mathcal{D}_{test}$ under different $\lambda$ values across each epoch on Tiny ImageNet with ViT for $10\%$ random data forgetting.
+As $\lambda$ increases, the accuracy on $\mathcal{D}_{f}$ drops quickly, showing stronger unlearning effectiveness, while the accuracy on $\mathcal{D}_{r}$ and $\mathcal{D}_{test}$ remains stable.
+In summary, the experimental results demonstrate that our framework notably enhances the forgetting quality while maintaining stable predictive performance.
 […]
@@
 Machine unlearning has emerged as a vital research topic due to several privacy, regulatory, and ethical concerns associated with machine learning models.
-It refers to the process of selectively removing specific data points from a trained machine-learning model.
-Generally, post-hoc machine unlearning can be divided into training-based (Graves et al., 2021; Warnecke et al., 2021; Thudi et al., 2022; Tarun et al., 2023) and training-free approaches (Guo et al., 2019; Sekhari et al., 2021; Nguyen et al., 2020; Golatkar et al., 2020b, 2021; Foster et al., 2024).
-To evaluate these methods, several unlearning metrics have been proposed, including UA (Brophy & Lowd, 2021; Foster et al., 2024) and MIA (Shokri et al., 2017; Chen et al., 2021).
+It refers to the process of selectively removing specific data points from a trained machine learning model.
+Generally, post-hoc machine unlearning can be divided into training-based Graves et al.
+(2021); Tarun et al.
+(2023); Thudi et al.
+(2022); Warnecke et al.
+(2021) and training-free approaches Foster et al.
+(2024); Golatkar et al.
+(2021, 2020); Guo et al.
+(2019); Nguyen et al.
+(2020); Sekhari et al.
+(2021).
+To evaluate these methods, several unlearning metrics have been proposed, including UA Brophy and Lowd (2021); Foster et al.
+(2024) and MIA Shokri et al.
+(2017); Chen et al.
+(2021).
 However, these metrics often fail to account for the confidence of the forgetting quality.
-To address this limitation, we improve it in our work motivated by conformal prediction (Angelopoulos & Bates, 2021), which stands out among uncertainty quantification techniques for its ability to provide well-calibrated, reliable confidence measures.
+To address this limitation, we improve it in our work motivated by conformal prediction Angelopoulos and Bates (2021), which stands out among uncertainty quantification techniques for its ability to provide well-calibrated, reliable confidence measures.
 As a generic methodology, conformal prediction can transform the outputs of any black box prediction algorithm into a prediction set.
-Due to its versatility, many works have specifically designed numerous conformal prediction methods tailored to particular prediction problems (Papadopoulos et al., 2002; Lei & Wasserman, 2014; Romano et al., 2020; Lei et al., 2018).
-One work (Becker & Liebig, 2022) has primarily focused on parameter-level uncertainty without fully addressing the broader implications of unlearning on prediction confidence.
+Due to its versatility, many works have specifically designed numerous conformal prediction methods tailored to particular prediction problems Lei et al.
+(2018); Lei and Wasserman (2014); Papadopoulos et al.
+(2002); Romano et al.
+(2020).
+One work Becker and Liebig (2022) has primarily focused on parameter-level uncertainty without fully addressing the broader implications of unlearning on prediction confidence.
 It assesses the sensitivity of model parameters to the target data through Fisher Information Matrix, but they often rely on computationally intensive operations and may struggle to scale to large models or datasets.
@@
 Together, we provide a more rigorous foundation for privacy-preserving machine learning.
-Impact Statement
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Appendix
 Appendix A Setting Details
-For CIFAR-10 with a ResNet-18 architecture, we train the original model from scratch for 200 epochs using SGD with a Cosine Annealing learning rate schedule, starting from an initial learning rate of 0.1.
-We set momentum to 0.9 and a batch size of 64.
+For CIFAR-10 with ResNet-18 architecture, we train the original model from scratch for 200 epochs using SGD with a Cosine Annealing learning rate schedule, starting from an initial learning rate of 0.1.
+We set the momentum to 0.9 and a batch size of 64.
+The RT model adopts the same training configuration.
+Other models are trained for the following durations: FT for 20 epochs, RL for 10 epochs, SalUn for 10 epochs, GA for 1 epoch (to avoid over-forgetting and significant RA degradation), NegGrad+ for 10 epochs (reduced to 2 epochs in class-wise scenarios), and SFRon for 10 epochs.
+All other hyperparameters match those of the original model.
 For the ViT architecture, we initialize the original model by training a pretrained ViT model for 15 epochs on Tiny ImageNet.
 We start with a learning rate of 0.001, while other training parameters match those used for ResNet-18.
-In class-wise forgetting, in each dataset, we designate the same class as the forget data $\mathcal{D}_{f}$ across all unlearning methods, with the remaining classes used as the retain data $\mathcal{D}_{r}$.
+We use SGD and set the momentum to 0.9 and a batch size of 64.
+The RT model follows the same training procedure as the original model.
+Other models are trained for the following durations: FT for 5 epochs, RL for 5 epochs, SalUn for 5 epochs, GA for 1 epoch, NegGrad+ for 5 epochs, and SFRon for 5 epochs.
+All other hyperparameters are consistent with the original model’s training.
+For CIFAR-10/Tiny ImageNet, we randomly select $200$/$50$ data points per class ($2000$/$10000$ data points in total) as calibration data $\mathcal{D}_{c}$ and $\mathcal{D}_{c}^{\prime}$, respectively.
+The calibration data $\mathcal{D}_{c}$ does not participate in the model training or unlearning processes and is only used for calibrating the threshold $\hat{q}$, while $\mathcal{D}_{c}^{\prime}$ is used in the process of our unlearning framework to generate $\bar{q}$.
+All experiments are conducted on 1 Tesla V100-SXM2 GPU card with 32GB memory in a single node.
 Appendix B Evaluating MU methods
@@
 Teacher 40 4 10% 400 37 9.3%
-FF 2,995 2,751 91.9% 15,083 14,061 93.2%
 SSD 25 2 8.0% 116 9 7.8%
 NegGrad+ 435 115 26.4% 711 249 35.5%
-Salun 185 117 63.2% 1065 695 65.3%
+Salun 185 117 63.2% 1,065 695 65.3%
+SFRon 240 125 52.1% 1,000 610 61.0%
 Mis-label and In-set Ratio of MIA
@@
 Teacher 638 586 91.8% 1,689 895 53.0%
-FF 1,424 1,424 100% 5,996 4,850 80.9%
 SSD 61 11 18.0% 282 24 8.5%
@@
 Salun 2,121 1,848 87.1% 10,221 9,121 89.2%
+SFRon 423 121 28.6% 1,871 433 23.1%
 Conformal prediction is applied to UA and MIA predictions to determine the number of misclassified data points (mis-label) and the number of these points that fall within the conformal prediction set (in-set) across 9 unlearning methods.
 We evaluate both the UA and MIA metrics by counting the misclassified data points and calculating how many of them are included in the conformal prediction set.
-The detailed results are presented in Table 8.
-B.2 CR Metric
-Tables 9 and 10 show the unlearning performance of seven unlearning methods on CIFAR-10 with ResNet-18 in $10\%$ and $50\%$ random data forgetting scenarios, while Table 11 is the results in class-wise forgetting scenario.
-Tables 12 and 13 present the unlearning performance on Tiny ImageNet with ResNet-18 in random data forgetting scenario, while Table 11 details the unlearning performance in class-wise forgetting scenario.
-For all machine unlearning methods, increasing the $\alpha$ level leads to lower Coverage and smaller Set Size.
-Conversely, the CR tends to grow as $\alpha$ increases.
-Most methods yield a Set Size of less than $1$ when $\alpha$ is set too low, since conformal prediction is fundamentally linked to the model’s baseline prediction performance (accuracy).
-If the model’s accuracy significantly exceeds the confidence level, conformal prediction can easily achieve the required coverage.
-In fact, it may produce empty prediction sets for certain data points while still satisfying the coverage target.
-Excessively high $\alpha$ values can distort evaluation results, preventing the CR metric from accurately reflecting the model’s performance.
-Therefore, the choice of $\alpha$ should be carefully considered in relation to the model’s inherent performance.
-The more experimental results for random data forgetting scenarios have already been analyzed in detail in Section 4 and will not be reiterated here.
+The detailed results are presented in Table 6, which is the extended results of Table 2.
+B.2 Distribution Comparison of Forgotten Data on UA and CR
+As shown in Figures 7-10, we further analyze the probability and loss distributions of ground truth labels for data identified as truly forgotten by CR (i.e., out-set) and UA (i.e., mis-label) respectively.
+The distribution curves are fitted using KDE for clearer visualization.
+The softmax outputs for ‘out-set’ are consistently near 00 compared to ‘mis-label’, which strongly suggests that ‘out-set’ more rigorously captures real forgotten data.
+In cross entropy loss distribution, forgotten data identified by CR consistently show higher cross entropy loss than UA.
+Higher loss indicates better forgetting quality, which further validates that CR better remove fake unlearning data.
+B.3 CR Metric
+Tables 7 and 8 show the unlearning performance of 9 unlearning methods on CIFAR-10 with ResNet-18 in $10\%$ and $50\%$ random data forgetting scenarios, while Table 9 is the results in class-wise forgetting scenario.
+Tables 10 and 11 present the unlearning performance on Tiny ImageNet with ResNet-18 in random data forgetting scenario, while Table 12 details the unlearning performance in class-wise forgetting scenario.
+For class-wise forgetting scenario, we note $\mathcal{D}_{test}=\mathcal{D}_{tf}\cup\mathcal{D}_{tr}$.
+$\mathcal{D}_{tf}$ corresponds to the test-forget data exclusively containing the forget class, while $\mathcal{D}_{tr}$ represents the test-retain data within the test data $\mathcal{D}_{test}$.
+For all unlearning methods, as $\alpha$ level increases, it results in reduced Coverage and smaller Set Size.
+This happens because a higher $\alpha$ loosens the conformal threshold $\hat{q}$, allowing fewer predictions to be included within the prediction set for each data point.
+On the contrary, the CR tends to increase with increasing $\alpha$.
+Although both Coverage and Set Size may decrease, Set Size often decreases more significantly.
+Consequently, the CR value of $\mathcal{D}_{f}$ generally becomes larger as $\alpha$ increases.
+It is natural that the adjustment of $\alpha$ affects both Coverage and Set Size.
+However, the final CR value really depends on the model’s performance itself.
+For a strict evaluation, we encourage setting $\alpha$ to $0.5$.
+When $\alpha$ is set to $0.2$, most methods show a value of Set Size less than $1$ in both Table 7, 8, 10, 11.
+The intuition behind it is that conformal prediction, as a static predictor, is intrinsically tied to the model’s base prediction performance and accuracy.
+When the model’s accuracy is significantly higher than the confidence level, conformal prediction can achieve the required coverage with ease.
+In fact, it can generate partial empty prediction sets for some data points while still meeting the target coverage.
+Thus, the choice of $\alpha$ is crucial.
+Overly high $\alpha$ values may skew evaluation results by failing to let CR accurately reflect model performance.
+Therefore, we emphasize that a small $\alpha$ is generally appropriate for most unlearning scenarios.
 Notably, the insights gained from the random data forgetting scenario can also be extended to the class-wise forgetting scenario.
 Additionally, in the class-wise scenario, some unlearning methods like RT and RL with UA = $100$ and CR approaching 00 indicate they are truly effective at forgetting the specified class.
-B.3 MIACR Metric
-Table 15 presents the performance of seven machine unlearning methods on CIFAR-10 in ResNet-18, evaluated with the MIACR metric.
+B.4 MIACR Metric
+Table 13 presents the performance of 9 machine unlearning methods on CIFAR-10 in ResNet-18, evaluated with the MIACR metric.
 In addition to the settings discussed in Section 4, we include results for $\alpha\in[0.05,0.15]$ in Table 13.
@@
 Our unlearning framework shifts the distribution of the forget data to the right, demonstrating improved forgetting quality.
-B.4 How can we better measure forgetting under distribution shifts
-Recall that the coverage value of RL has a relatively large deviation in Table 5 since it employs label corruption in its unlearning strategy which can cause distribution shifts.
+B.5 Measuring Forgetting under Distribution Shifts
+RL and Salun are unlearning methods employ label corruption in their unlearning strategy which can cause distribution shifts.
 Here, we introduce how to better measure forgetting under these circumstances.
@@
 Furthermore, we conduct an ablation study and analyze the impact of using our unlearning framework.
-As illustrated in Figures 3 and 4, we compare the density distributions of non-conformity scores for calibration data $\mathcal{D}_{c}$ and forget data $\mathcal{D}_{f}$ under the RT, FT, and RL unlearning methods.
+As illustrated in Figures 5 and 6, we compare the density distributions of non-conformity scores for calibration data $\mathcal{D}_{c}$ and forget data $\mathcal{D}_{f}$ under the RT, FT, and RL unlearning methods.
 We set $\lambda$ to 1.
 Clearly, a higher non-conformity score for $\mathcal{D}_{f}$ indicates that it is less likely to be included in the conformal prediction set, reflecting more effective forgetting.
-Comparing Figures 3 and 4, it is evident that after applying our unlearning framework, the distribution of non-conformity scores for forget data shifts noticeably to the right.
-This demonstrates the effectiveness of our framework in enhancing forgetting quality.
-Table 9: Unlearning performance of 9 unlearning methods on CIFAR-10 with ResNet-18 in $10\%$ random data forgetting scenario.
+Comparing Figures 5 and 6, after applying our unlearning framework, we observe a significant rightward shift in the non-conformity score distribution of forget data, which is a promising signal according to evaluation criterion ➋.
+Furthermore, the FT distribution in Figure 6 exhibits substantial overlap with the calibration data, nearly matching the distribution observed in RT.
+Based on evaluation criterion ➊, since calibration data represents unseen examples, the similarity between forget data and calibration data distributions provides strong evidence of effective forgetting.
+Overall, the results evaluated on both evaluation criteria ➊ and ➋ consistently confirm the efficacy of our framework in enhancing forgetting quality.
+Appendix D Limitations
+While our work provides a new perspective on evaluating and improving machine unlearning via uncertainty quantification, there are a few limitations to consider.
+First, our current study focuses exclusively on image classification tasks.
+While this setting is representative and widely used in prior unlearning literature Foster et al.
+(2024); Graves et al.
+(2021); Shen et al.
+(2024), extending our framework to other domains such as language modeling or generative models remains an important direction for future work.
+Second, our evaluation method requires a separate validation set for uncertainty calibration.
+But this requirement is generally not restrictive in real-world unlearning settings, as model controllers (e.g., service providers) typically have full access to the training and validation data.
+Moreover, in the scarce training set scenario, where access to data is inherently limited, evaluation is often unnecessary because retraining from scratch is computationally feasible and serves as a natural baseline.
+Thus, the need for calibration data does not pose a significant barrier to applying our approach in most practical contexts.
+Appendix E Impact Statement
+Our work on machine unlearning through uncertainty quantification perspective offers several positive contributions to society.
+Enhanced Privacy Protection.
+Effective machine unlearning is crucial for upholding individual privacy rights, particularly in contexts where regulations like GDPR mandate the "right to be forgotten." Our method provides a mechanism to verify that unlearning is genuine, ensuring that user data is truly removed from models when requested.
+Support for Trustworthy AI Development.
+Our approach provides technical foundations for AI systems capable of "forgetting" harmful, biased, or outdated information, facilitating continuous improvement and the trustworthiness of AI systems.
+(a) FT
+(b) RL
+(c) Salun
+(d) SFRon
+[…]
+(a) FT
+(b) RL
+(c) Salun
+(d) SFRon
+[…]
+(a) FT
+(b) RL
+(c) Salun
+(d) SFRon
+[…]
+(a) FT
+(b) RL
+(c) Salun
+(d) SFRon
+[…]
+Table 7: Unlearning performance of 10 unlearning methods on CIFAR-10 with ResNet-18 in $10\%$ random data forgetting scenario.
 The results are reported in the format a±$\pm$±b, where a is the mean and b is the standard deviation from $3$ independent trials.
@@
 0.2 $0.865_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.085})$ $0.806_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.866_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.077})$ $0.816_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.998_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.988_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.426_{\pm 0.007}$
-FF 0.05 $0.973_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.949_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $7.966_{\pm 0.212}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.877})$ $7.408_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.334})$ $0.122_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.742})$ $0.128_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.751})$ $0.999_{\pm 0.000}$
-0.1 $0.933_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.899_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $7.129_{\pm 0.148}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.195})$ $6.566_{\pm 0.166}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.619})$ $0.131_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.812})$ $0.137_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.808})$ $0.998_{\pm 0.001}$
-UA59.9%, RA40.1%, TA41.1% 0.15 $0.888_{\pm 0.029}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.068})$ $0.852_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.013})$ $6.431_{\pm 0.078}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.590})$ $5.903_{\pm 0.262}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.036})$ $0.138_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.837})$ $0.144_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.824})$ $0.996_{\pm 0.002}$
-0.2 $0.835_{\pm 0.048}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.794_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.014})$ $5.750_{\pm 0.034}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.961})$ $5.219_{\pm 0.368}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.395})$ $0.145_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.843})$ $0.153_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.828})$ $0.993_{\pm 0.003}$
 SSD 0.05 $0.996_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.945_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.999_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.090})$ $1.008_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.066})$ $0.994_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.130})$ $0.936_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.057})$ $0.622_{\pm 0.019}$
@@
 0.2 $0.788_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.794_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.014})$ $0.794_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.821_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.992_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $0.966_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.221_{\pm 0.005}$
-Table 10: Unlearning performance of 9 unlearning methods on CIFAR-10 with ResNet18 in $50\%$ random data forgetting scenario.
+SFRon 0.05 $0.977_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.953_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $1.100_{\pm 0.023}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $1.143_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.069})$ $0.889_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.025})$ $0.834_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.045})$ $0.926_{\pm 0.018}$
+0.1 $0.945_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.064})$ $0.905_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.986_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.052})$ $0.977_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.030})$ $0.958_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.927_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.018})$ $0.435_{\pm 0.043}$
+UA4.8%, RA97.4%, TA91.4% 0.15 $0.895_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.075})$ $0.847_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.912_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.071})$ $0.879_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.982_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.963_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.082_{\pm 0.007}$
+0.2 $0.857_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.077})$ $0.808_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.868_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.079})$ $0.826_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.988_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.978_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.025_{\pm 0.005}$
+Table 8: Unlearning performance of 10 unlearning methods on CIFAR-10 with ResNet18 in $50\%$ random data forgetting scenario.
 Coverage Set Size CR
@@
 $0.2$ $0.818_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.798_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.016})$ $0.823_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.826_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $0.997_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.033})$ $0.971_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.013})$ $0.793_{\pm 0.012}$
-FF $0.05$ $0.972_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.017})$ $0.954_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $8.023_{\pm 0.189}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.736})$ $7.461_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.247})$ $0.121_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.621})$ $0.128_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.652})$ $0.999_{\pm 0.000}$
-$0.1$ $0.930_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.897_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $7.091_{\pm 0.151}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.068})$ $6.521_{\pm 0.171}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.501})$ $0.131_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.747})$ $0.138_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.748})$ $0.998_{\pm 0.001}$
-UA60.0%, RA40.1%, TA40.6% $0.15$ $0.887_{\pm 0.024}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.054})$ $0.852_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $6.402_{\pm 0.034}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.519})$ $5.837_{\pm 0.438}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.931})$ $0.139_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.804})$ $0.146_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.788})$ $0.996_{\pm 0.001}$
-$0.2$ $0.840_{\pm 0.046}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.058})$ $0.803_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $5.805_{\pm 0.042}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.994})$ $5.253_{\pm 0.376}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.403})$ $0.145_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.819})$ $0.153_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.804})$ $0.993_{\pm 0.003}$
 SSD $0.05$ $0.993_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.038})$ $0.944_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.999_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.288})$ $1.001_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.213})$ $0.995_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.253})$ $0.941_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.161})$ $0.585_{\pm 0.014}$
@@
 $0.2$ $0.856_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.074})$ $0.796_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $0.872_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.060})$ $0.844_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.982_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $0.943_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.669_{\pm 0.008}$
-Table 11: Unlearning performance of 9 unlearning methods on CIFAR-10 with ResNet18 in class-wise forgetting scenario.
+SFRon 0.05 $0.977_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.022})$ $0.953_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $1.100_{\pm 0.023}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.188})$ $1.143_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.071})$ $0.889_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.147})$ $0.834_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.054})$ $0.926_{\pm 0.018}$
+0.1 $0.945_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$ $0.905_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.986_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.037})$ $0.977_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.044})$ $0.958_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.081})$ $0.927_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.042})$ $0.435_{\pm 0.043}$
+UA4.0%, RA97.3%, TA91.6% 0.15 $0.895_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.062})$ $0.847_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.912_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.879_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$ $0.982_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.039})$ $0.963_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.082_{\pm 0.007}$
+0.2 $0.857_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.075})$ $0.808_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.868_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.056})$ $0.826_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $0.988_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $0.978_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.020})$ $0.025_{\pm 0.005}$
+Table 9: Unlearning performance of 10 unlearning methods on CIFAR-10 with ResNet18 in class-wise forgetting scenario.
 Coverage Set Size CR
@@
 RA99.5%, TA94.0% 0.2 $0.809_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.191})$ $0.841_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.159})$ $0.816_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $8.141_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.859})$ $7.525_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.475})$ $0.824_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.099_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.112_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.990_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.916_{\pm 0.005}$ $0.010_{\pm 0.003}$
-FF 0.05 $0.984_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.016})$ $0.962_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.038})$ $0.950_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $7.760_{\pm 1.386}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.240})$ $7.479_{\pm 1.223}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.521})$ $7.627_{\pm 0.058}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.479})$ $0.129_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.130_{\pm 0.021}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.030})$ $0.125_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.715})$ $0.999_{\pm 0.000}$ $0.997_{\pm 0.003}$
-0.1 $0.964_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.917_{\pm 0.037}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.083})$ $0.909_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$ $7.360_{\pm 0.999}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.640})$ $6.988_{\pm 0.749}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.012})$ $6.853_{\pm 0.390}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.930})$ $0.132_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.132_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.133_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.823})$ $0.998_{\pm 0.000}$ $0.996_{\pm 0.003}$
-UA100%, UAtf100%, 0.15 $0.922_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.078})$ $0.852_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.148})$ $0.863_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $6.941_{\pm 1.088}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.059})$ $6.562_{\pm 0.856}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.438})$ $6.278_{\pm 0.232}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.395})$ $0.134_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.034})$ $0.131_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$ $0.138_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.832})$ $0.997_{\pm 0.001}$ $0.994_{\pm 0.006}$
-RA34.3%, TA36.1% 0.2 $0.897_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.103})$ $0.818_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.182})$ $0.789_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.025})$ $6.757_{\pm 0.978}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.243})$ $6.352_{\pm 0.757}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.648})$ $5.307_{\pm 0.527}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.477})$ $0.134_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.034})$ $0.130_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.030})$ $0.150_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.832})$ $0.993_{\pm 0.002}$ $0.993_{\pm 0.006}$
 SSD 0.05 $0.995_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.935_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.065})$ $0.940_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.024})$ $1.030_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}8.970})$ $1.067_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}8.933})$ $0.991_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.157})$ $0.966_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.866})$ $0.876_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.776})$ $0.949_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.109})$ $0.804_{\pm 0.015}$ $0.447_{\pm 0.007}$
@@
 RA99.5%, TA94.3% 0.2 $0.895_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.105})$ $0.816_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.184})$ $0.823_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.895_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}9.105})$ $0.850_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}9.150})$ $0.831_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.999_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.899})$ $0.960_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.860})$ $0.991_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $0.078_{\pm 0.003}$ $0.002_{\pm 0.009}$
+NegGrad+ 0.05 $0.989_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.961_{\pm 0.056}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.039})$ $0.945_{\pm 0.027}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $9.432_{\pm 0.803}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.568})$ $9.038_{\pm 1.360}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.962})$ $1.053_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.096})$ $0.105_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.107_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.897_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.058})$ $1.000_{\pm 0.000}$ $0.835_{\pm 0.085}$
+0.1 $0.980_{\pm 0.029}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.020})$ $0.954_{\pm 0.065}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.046})$ $0.881_{\pm 0.028}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $9.250_{\pm 1.061}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.750})$ $8.836_{\pm 1.647}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.164})$ $0.913_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.106_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.109_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.965_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $1.000_{\pm 0.000}$ $0.057_{\pm 0.021}$
+UA96.2%, UAtf95.2%, 0.15 $0.952_{\pm 0.068}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.048})$ $0.908_{\pm 0.130}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.092})$ $0.849_{\pm 0.026}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $8.600_{\pm 1.980}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.400})$ $8.077_{\pm 2.719}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.923})$ $0.868_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.014})$ $0.113_{\pm 0.018}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.013})$ $0.116_{\pm 0.023}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.016})$ $0.977_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $1.000_{\pm 0.000}$ $0.012_{\pm 0.003}$
+RA97.6%, TA92.8% 0.2 $0.958_{\pm 0.060}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.042})$ $0.921_{\pm 0.111}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.079})$ $0.814_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $8.673_{\pm 1.876}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.327})$ $8.219_{\pm 2.519}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.781})$ $0.828_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.112_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.115_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.983_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $1.000_{\pm 0.000}$ $0.004_{\pm 0.003}$
 Salun 0.05 $0.996_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $0.941_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.059})$ $0.952_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $9.996_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $9.892_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.108})$ $1.028_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.121})$ $0.100_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.095_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $0.926_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.087})$ $1.000_{\pm 0.000}$ $0.785_{\pm 0.049}$
@@
 RA99.6%, TA94.3% 0.2 $0.915_{\pm 0.019}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.085})$ $0.807_{\pm 0.038}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.193})$ $0.820_{\pm 0.035}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.005})$ $9.893_{\pm 0.024}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.107})$ $9.511_{\pm 0.192}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.489})$ $0.828_{\pm 0.039}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $0.092_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.085_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.015})$ $0.990_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $1.000_{\pm 0.000}$ $0.001_{\pm 0.001}$
-Table 12: Unlearning performance of 8 unlearning methods on Tiny ImageNet with ViT in $10\%$ random data forgetting scenario.
+SFRon 0.05 $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.952_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.013})$ $10.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $10.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.022_{\pm 0.030}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.127})$ $0.100_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.100_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.932_{\pm 0.024}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.092})$ $1.000_{\pm 0.000}$ $0.677_{\pm 0.206}$
+0.1 $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.908_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.026})$ $10.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $10.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.937_{\pm 0.028}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.014})$ $0.100_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.100_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.970_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.014})$ $1.000_{\pm 0.000}$ $0.089_{\pm 0.092}$
+UA100%, UAtf100%, 0.15 $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.840_{\pm 0.026}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.016})$ $10.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $10.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.849_{\pm 0.026}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.033})$ $0.100_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.100_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.989_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.019})$ $1.000_{\pm 0.000}$ $0.002_{\pm 0.001}$
+RA99.3%, TA94.4% 0.2 $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.807_{\pm 0.024}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $10.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $10.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.813_{\pm 0.025}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.017})$ $0.100_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.100_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.992_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.010})$ $1.000_{\pm 0.000}$ $0.001_{\pm 0.001}$
+Table 10: Unlearning performance of 9 unlearning methods on Tiny ImageNet with ViT in $10\%$ random data forgetting scenario.
 Coverage Set Size CR
@@
 0.2 $0.870_{\pm 0.081}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.080})$ $0.810_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.845_{\pm 0.036}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.925_{\pm 0.046}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.071})$ $0.924_{\pm 0.047}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.894_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.041})$ $0.630_{\pm 0.003}$
-Table 13: Unlearning performance of 8 unlearning methods on Tiny ImageNet with ViT in $50\%$ random data forgetting scenario.
+SFRon 0.05 $0.989_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.045})$ $0.948_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $2.000_{\pm 0.059}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.124})$ $2.208_{\pm 0.037}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.368})$ $0.495_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.429_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.086})$ $0.986_{\pm 0.000}$
+0.1 $0.960_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.068})$ $0.899_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $1.227_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.076})$ $1.268_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.123})$ $0.783_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.709_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.077})$ $0.902_{\pm 0.003}$
+UA9.3%, RA97.0%, TA83.9% 0.15 $0.917_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.076})$ $0.849_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $1.024_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.068})$ $1.015_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.059})$ $0.896_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.016})$ $0.837_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.689_{\pm 0.012}$
+0.2 $0.866_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.076})$ $0.802_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.916_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.070})$ $0.892_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.037})$ $0.946_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.012})$ $0.899_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.426_{\pm 0.018}$
+Table 11: Unlearning performance of 9 unlearning methods on Tiny ImageNet with ViT in $50\%$ random data forgetting scenario.
 Coverage Set Size CR
@@
 0.2 $0.900_{\pm 0.044}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.114})$ $0.825_{\pm 0.049}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.910_{\pm 0.097}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.054})$ $0.969_{\pm 0.037}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.105})$ $0.928_{\pm 0.040}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$ $0.876_{\pm 0.063}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.045})$ $0.893_{\pm 0.002}$
-Table 14: Unlearning performance of 8 unlearning methods on Tiny ImageNet with ViT in class-wise forgetting scenario.
+SFRon 0.05 $0.994_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.048})$ $0.947_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $2.010_{\pm 0.188}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.136})$ $2.327_{\pm 0.087}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.222})$ $0.497_{\pm 0.045}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.057})$ $0.407_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.043})$ $0.983_{\pm 0.002}$
+0.1 $0.980_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.087})$ $0.900_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $1.245_{\pm 0.060}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.023})$ $1.338_{\pm 0.039}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.126})$ $0.788_{\pm 0.041}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.058})$ $0.673_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.069})$ $0.909_{\pm 0.003}$
+UA6.3%, RA96.8%, TA82.9% 0.15 $0.951_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.113})$ $0.849_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $1.041_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.065})$ $1.044_{\pm 0.023}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.067})$ $0.913_{\pm 0.028}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.813_{\pm 0.016}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.738_{\pm 0.029}$
+0.2 $0.910_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.125})$ $0.803_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.947_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.091})$ $0.910_{\pm 0.022}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.046})$ $0.961_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.044})$ $0.884_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.038})$ $0.523_{\pm 0.068}$
+Table 12: Unlearning performance of 9 unlearning methods on Tiny ImageNet with ViT in class-wise forgetting scenario.
 Coverage Set Size CR
@@
 RA98.4%, TA86.1% 0.2 $0.960_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.173})$ $0.840_{\pm 0.020}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.020})$ $0.801_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $184.838_{\pm 3.478}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}13.787})$ $177.647_{\pm 2.627}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.167})$ $0.863_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.001})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.928_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $1.000_{\pm 0.000}$ $0.230_{\pm 0.009}$
+SFRon 0.05 $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.948_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.002})$ $200.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $200.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $2.264_{\pm 0.254}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.479})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.423_{\pm 0.050}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.110})$ $1.000_{\pm 0.000}$ $0.990_{\pm 0.003}$
+0.1 $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.064})$ $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.040})$ $0.900_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $200.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}7.118})$ $200.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.660})$ $1.266_{\pm 0.044}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.120})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.711_{\pm 0.026}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.077})$ $1.000_{\pm 0.000}$ $0.912_{\pm 0.017}$
+UA100%, UAtf100%, 0.15 $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.096})$ $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.040})$ $0.850_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $200.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}13.209})$ $200.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}11.120})$ $1.009_{\pm 0.012}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.051})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.843_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.049})$ $1.000_{\pm 0.000}$ $0.668_{\pm 0.029}$
+RA96.1%, TA84.3% 0.2 $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.213})$ $1.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.140})$ $0.802_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $200.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}28.949})$ $200.000_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}25.520})$ $0.886_{\pm 0.006}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.026})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.005_{\pm 0.000}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.905_{\pm 0.007}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$ $1.000_{\pm 0.000}$ $0.358_{\pm 0.017}$
 Table 13: MIACR performance on CIFAR-10 with ResNet-18.
@@
 MIA93.24% ($50\%$ Forgetting) 0.2 $0.113_{\pm 0.008}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.133})$ $0.588_{\pm 0.019}$ $0.159_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.159})$ $0.532_{\pm 0.006}$
-FF 0.05 $0.038_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.053})$ $0.500_{\pm 0.003}$ $0.043_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.074})$ $0.508_{\pm 0.003}$
-0.1 $0.051_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.096})$ $0.486_{\pm 0.018}$ $0.089_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.112})$ $0.509_{\pm 0.013}$
-MIA71.52% ($10\%$ Forgetting) 0.15 $0.080_{\pm 0.015}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.123})$ $0.474_{\pm 0.013}$ $0.130_{\pm 0.017}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.142})$ $0.506_{\pm 0.007}$
-MIA76.02% ($50\%$ Forgetting) 0.2 $0.109_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.137})$ $0.473_{\pm 0.002}$ $0.168_{\pm 0.010}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.150})$ $0.499_{\pm 0.006}$
 SSD 0.05 $0.011_{\pm 0.011}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.080})$ $0.861_{\pm 0.012}$ $0.012_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.105})$ $0.748_{\pm 0.011}$
@@
 MIA59.12% ($50\%$ Forgetting) 0.2 $0.267_{\pm 0.009}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $0.608_{\pm 0.011}$ $0.220_{\pm 0.005}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.098})$ $0.586_{\pm 0.005}$
+SFRon 0.05 $0.017_{\pm 0.001}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.074})$ $0.711_{\pm 0.009}$ $0.017_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.100})$ $0.715_{\pm 0.008}$
+0.1 $0.040_{\pm 0.004}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.107})$ $0.626_{\pm 0.025}$ $0.046_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.155})$ $0.562_{\pm 0.013}$
+MIA91.55% ($10\%$ Forgetting) 0.15 $0.113_{\pm 0.003}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.090})$ $0.517_{\pm 0.003}$ $0.134_{\pm 0.013}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.138})$ $0.498_{\pm 0.003}$
+MIA92.52% ($50\%$ Forgetting) 0.2 $0.184_{\pm 0.002}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.062})$ $0.487_{\pm 0.002}$ $0.206_{\pm 0.014}({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.112})$ $0.483_{\pm 0.002}$
 Table 14: Performance of our unlearning framework. We show the unlearning performance on CIFAR-10 with ResNet-18 and Tiny ImageNet with ViT in $10\%$ random data forgetting scenario.
@@
 CIFAR-10 with ResNet-18
-$0.05$ $0.788({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.076})$ $0.824({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.763({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.101})$ $0.825({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.054})$ $0.719({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.145})$ $0.82({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.059})$
+$0.05$ $0.788({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.076})$ $0.824({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.055})$ $0.763({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.101})$ $0.825({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.054})$ $0.719({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.145})$ $0.820({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.059})$
 RT $0.1$ $10.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.2})$ $98.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.4})$ $91.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.8})$ $0.914({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.029})$ $0.924({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $14.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.4})$ $97.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.9})$ $90.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.4})$ $0.879({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.064})$ $0.912({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.033})$ $17.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}9.1})$ $96.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.9})$ $90.5\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.3})$ $0.838({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.105})$ $0.911({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.034})$
@@
 Tiny ImageNet with ViT
-$0.05$ $0.458({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.486})$ $0.516({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.433})$ $0.396({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.548})$ $0.489({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.460})$ $0.346({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.598})$ $0.481({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.468})$
-RT $0.1$ $19.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.6})$ $98.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $86.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $0.729({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.163})$ $0.786({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.114})$ $26.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}11.7})$ $98.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.1})$ $85.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.2})$ $0.649({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.243})$ $0.765({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.135})$ $35.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}21.0})$ $98.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.2})$ $85.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.8})$ $0.549({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.343})$ $0.739({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.161})$
-$0.15$ $0.841({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.889({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.039})$ $0.768({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.073})$ $0.88({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.030})$ $0.658({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.183})$ $0.861({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.011})$
-$0.2$ $0.898({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.108})$ $0.932({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.133})$ $0.839({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.049})$ $0.929({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.130})$ $0.743({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$ $0.918({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.119})$
-$0.05$ $0.441({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.503})$ $0.399({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.550})$ $0.413({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.531})$ $0.401({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.548})$ $0.342({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.602})$ $0.363({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.586})$
-FT $0.1$ $9.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.9})$ $97.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.4})$ $83.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.753({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.139})$ $0.683({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.217})$ $13.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $97.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.6})$ $83.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.718({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.174})$ $0.683({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.217})$ $20.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.3})$ $96.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $82.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.1})$ $0.627({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.265})$ $0.652({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.248})$
-$0.15$ $0.884({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.043})$ $0.823({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$ $0.848({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.007})$ $0.819({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.031})$ $0.772({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.069})$ $0.802({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.048})$
-$0.2$ $0.942({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.152})$ $0.893({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.094})$ $0.914({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.124})$ $0.890({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.091})$ $0.856({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.066})$ $0.877({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.078})$
-$0.05$ $0.051({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.893})$ $0.111({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.838})$ $0.051({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.893})$ $0.121({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.828})$ $0.048({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.896})$ $0.119({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.830})$
-RL $0.1$ $31.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}17.1})$ $95.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}17.9})$ $80.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.1})$ $0.278({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.614})$ $0.451({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.449})$ $36.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}21.5})$ $95.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.5})$ $80.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.6})$ $0.254({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.638})$ $0.449({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.451})$ $40.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}25.5})$ $94.5\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.3})$ $79.5\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.5})$ $0.236({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.656})$ $0.436({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.464})$
-$0.15$ $0.579({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.262})$ $0.710({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.140})$ $0.541({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.300})$ $0.708({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.142})$ $0.480({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.361})$ $0.673({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.177})$
-$0.2$ $0.752({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.038})$ $0.825({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.026})$ $0.718({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.072})$ $0.827({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.028})$ $0.642({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.148})$ $0.793({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$
+$0.05$ $0.458({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.045})$ $0.516({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.396({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.107})$ $0.489({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.027})$ $0.346({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.157})$ $0.481({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.035})$
+RT $0.1$ $19.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.6})$ $98.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $86.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.0})$ $0.729({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.046})$ $0.786({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $26.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}11.7})$ $98.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.1})$ $85.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.2})$ $0.649({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.126})$ $0.765({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.021})$ $35.7\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}21.0})$ $98.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.2})$ $85.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.8})$ $0.549({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.226})$ $0.739({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.047})$
+$0.15$ $0.841({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.039})$ $0.889({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.000})$ $0.768({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.112})$ $0.880({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.009})$ $0.658({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.222})$ $0.861({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.028})$
+$0.2$ $0.898({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.036})$ $0.932({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.003})$ $0.839({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.095})$ $0.929({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.006})$ $0.743({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.191})$ $0.918({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.017})$
+$0.05$ $0.441({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.062})$ $0.399({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.117})$ $0.413({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.090})$ $0.401({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.115})$ $0.342({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.161})$ $0.363({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.153})$
+FT $0.1$ $9.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.9})$ $97.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.4})$ $83.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.753({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.022})$ $0.683({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.103})$ $13.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.9})$ $97.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}1.6})$ $83.6\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $0.718({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.057})$ $0.683({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.103})$ $20.0\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.3})$ $96.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.4})$ $82.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.1})$ $0.627({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.148})$ $0.652({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.134})$
+$0.15$ $0.884({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.004})$ $0.823({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.066})$ $0.848({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.032})$ $0.819({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.070})$ $0.772({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.108})$ $0.802({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.087})$
+$0.2$ $0.942({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.008})$ $0.893({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.042})$ $0.914({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.020})$ $0.890({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.045})$ $0.856({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.078})$ $0.877({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.058})$
+$0.05$ $0.051({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.452})$ $0.111({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.405})$ $0.051({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.452})$ $0.121({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.395})$ $0.048({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.455})$ $0.119({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.397})$
+RL $0.1$ $31.8\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}17.1})$ $95.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}17.9})$ $80.9\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.1})$ $0.278({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.497})$ $0.451({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.335})$ $36.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}21.5})$ $95.3\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}3.5})$ $80.4\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}5.6})$ $0.254({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.521})$ $0.449({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.337})$ $40.2\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}25.5})$ $94.5\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.3})$ $79.5\%({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}6.5})$ $0.236({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.539})$ $0.436({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.350})$
+$0.15$ $0.579({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.301})$ $0.710({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.179})$ $0.541({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.339})$ $0.708({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.181})$ $0.480({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.400})$ $0.673({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.216})$
+$0.2$ $0.752({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.182})$ $0.825({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.110})$ $0.718({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.216})$ $0.827({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.108})$ $0.642({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.292})$ $0.793({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.142})$